### v5

# Models of Set Theory: Extensions and Dead-ends

Abstract.
This paper is a contribution to the study of extensions of arbitrary models of $\mathrm{ZF}$ (Zermelo-Fraenkel set theory), with no regard to countability or well-foundedness of the models involved. Our main results include the theorems below; in Theorems A and B, $\mathcal{N}$ is said to be a conservative elementary extension of $\mathcal{M}$ if $\mathcal{N}$ elementarily extends $\mathcal{M}$, and the intersection of every $\mathcal{N}$-definable set with $M$ is $\mathcal{M}$-definable (parameters allowed). In Theorem B, ZFC is the result of augmenting ZF with the axiom of choice.
Theorem A. Every model $\mathcal{M}$ of $\mathrm{ZF+}\exists p\left(\mathrm{V}=\mathrm{HOD}(p)\right)$ has a conservative elementary extension $\mathcal{N}$ that contains an ordinal above all of the ordinals of $\mathcal{M}$.
Theorem B. If $\mathcal{N}$ is a conservative elementary extension of a model $\mathcal{M}$ of $\mathrm{ZFC}$, and $\mathcal{N}$ has the same natural numbers as $\mathcal{M}$, then $\mathcal{M}$ is cofinal in $\mathcal{N}$.
Theorem C. Every consistent extension of $\mathrm{ZFC}$ has a model $\mathcal{M}$ of power $\aleph_{1}$ such that $\mathcal{M}$ has no proper end extension to a model of $\mathrm{ZF}$.
1. INTRODUCTION
This paper concerns the model theory of Zermelo-Fraenkel ($\mathrm{ZF}$) set theory. More specifically, it concerns the types of extensions that are possible for arbitrary models of $\mathrm{ZF}$ (as opposed to countable or well-founded ones), a topic that I have worked on intermittently for over four decades. Indeed, Theorem C of the abstract answers the following question that was initially posed in my 1984 doctoral dissertation [En-1, Question 1.1.12]:
[…]
Note that assuming that strongly inaccessible cardinals are cofinal in the class of ordinals, every well-founded model of $\mathrm{ZF}$ can be properly end extended to a model of $\mathrm{ZF}$. The motivation for the above question emerges from the comparative study of the model theory of $\mathrm{ZF}$ and $\mathrm{PA}$ (Peano Arithmetic). Prima facie $\mathrm{ZF}$ and $\mathrm{PA}$ are unrelated; after all $\mathrm{PA}$ axiomatizes basic intuitions concerning the familiar arithmetical structure of natural numbers, whereas $\mathrm{ZF}$ codifies the laws governing Cantor’s mysterious universe of sets, a vastly more complex structure. However, thanks to a clever coding idea, introduced first by Ackermann [A], $\mathrm{ZF}$ and $\mathrm{PA}$ turn out to be intimately connected at a formal level: $\mathrm{PA}$ is bi-interpretable [^2] The role of TC was elucidated in [ESV] by showing that TC cannot be dropped in this bi-interpretability result. The details of this bi-interpretability result are worked out by Kaye and Wong in [KW]. Their work shows that indeed the two theories are definitionally equivalent (aka synonymous), i.e., they have a common definitional extension. with the theory $\mathrm{ZF}_{\mathrm{fin}}+\mathrm{TC}$, where $\mathrm{ZF}_{\mathrm{fin}}$ is obtained from $\mathrm{ZF}$ by replacing the axiom of infinity by its negation, and $\mathrm{TC}$ expresses “the transitive closure of every set exists”. In this light it is natural to compare and contrast the model theoretic behavior of models of $\mathrm{ZF}$ and $\mathrm{PA}$ in order to elucidate the role of the axiom of infinity in Zermelo-Fraenkel set theory. A central result in the model theory of PA is the MacDowell-Specker Theorem [MS], and its refinement (independently) by Phillips [Ph] and Gaifman [G-2], that states that every model of PA (of any cardinality) has a proper conservative elementary end extension. Using the aforementioned bi-interpretation between $\mathrm{PA}$ and $\mathrm{ZF}$, the refinement of the MacDowell-Specker Theorem is equivalent to:
[…]
In the above, $\mathcal{N}$ is said to be a conservative elementary extension of $\mathcal{M}$ if $\mathcal{N}$ elementarily extends $\mathcal{M}$, and the intersection of every $\mathcal{N}$-definable set with $M$ is $\mathcal{M}$-definable (parameters allowed). It is known that Theorem $\triangledown$ becomes false if $\mathrm{ZF}_{\mathrm{fin}}+\mathrm{TC}$ is replaced by $\mathrm{ZF}$, but the answer to Question $\heartsuit$ (Theorem 5.18) remained open until now. [^3] By a remarkable theorem of Keisler and Morley [KS] every model of $\mathrm{ZF}$ of countable cofinality has arbitrarily large elementary end extensions, however, as shown by Keisler and Silver [KS] not every model of $\mathrm{ZF}$ has an elementary end extension. For an exposition of the general framework for the analogues of the MacDowell-Specker Theorem for set theory, see [En-5].
The foremost motivation in writing this paper was to present the rather intricate details of the solution to the above 40-year old question to my younger self. Another motivation was to publicize what can be described as a long overdue ‘Additions and Corrections’ to my 1987 paper [En-3], namely Theorems A and B of the abstract (Theorem 3.1 and Corollary 4.2 of the paper, respectively).
Section 2 is devoted to preliminaries needed for Sections 3, 4, and 5 in which the main results are presented. Section 3 contains the proof of Theorem A of the abstract (and closely related material). Section 4 contains results on faithful extensions (a generalization of conservative extensions), including Theorem 4.1 that yields Theorem B of the abstract (as Corollary 4.2). A key result in Section 4 is Theorem 4.4 which is used in Section 5 in conjunction with other machinery to establish Theorem C of the abstract. Section 6 is an Addendum and Corrigendum to my aforementioned 1987 paper [En-3]. The Appendix presents the proof of a theorem due to Rubin, Shelah, and Schmerl that plays a key role in the proof of Theorem C of the abstract.
The main results of the paper were presented online in February 2025 at the CUNY set theory seminar. I am grateful to Vika Gitman and Gunter Fuchs and other participants of the seminar for helpful feedback. I am also grateful to Corey Switzer for illuminating discussions about the forcing argument used in the proof of Step 2 of Stage 2 of Theorem A.1 (in the Appendix).
2. PRELIMINARIES
Here we collect the basic definitions, notations, conventions, and results that are relevant for the subsequent sections.
2.1. Definitions and Basic Facts (Languages and theories of sets).
(a) $\mathcal{L}_{\mathrm{set}}=\{=,\in\}$ is the usual language of $\mathrm{ZF}$-set theory. We treat $\mathrm{ZF}$ as being axiomatized as usual, except that instead of including the scheme of replacement among the axioms of $ZF$, we include the schemes of separation and collection, as in [CK, Appendix A]. Thus, in our setup the axioms of Zermelo set theory $\mathrm{Z}$ are obtained by removing the scheme of collection from the axioms of $\mathrm{ZF}$. More generally, for $\mathcal{L}\supseteq\mathcal{L}_{\mathrm{set}}$ we construe $\mathrm{ZF}(\mathcal{L})$ to be the natural extension of $\mathrm{ZF}$ in which the schemes of separation and collection are extended to $\mathcal{L}$-formulae; similarly we will use $\mathrm{Z}(\mathcal{L})$ to denote Zermelo set theory over $\mathcal{L}$, i.e., as the result of extending $\mathrm{Z}$ with the $\mathcal{L}$-separation scheme $\mathrm{Sep}(\mathcal{L})$. Officially speaking, $\mathrm{Sep}(\mathcal{L})$ consists of the universal closures of $\mathcal{L}$-formulae of the form:
$\forall v\exists w\forall x(x\in w\longleftrightarrow x\in v\wedge\varphi(x,\overline{y})).$
Thus $\mathrm{ZF}(\mathcal{L})$ is the result of augmenting $\mathrm{Z}(\mathcal{L})$ with the $\mathcal{L}$-collection scheme $\mathrm{Coll}(\mathcal{L})$, which consists of the universal closures of $\mathcal{L}$-formulae of the form:
$\left(\forall x\in v\text{ }\exists y\text{\ }\varphi(x,y,\overline{z})\right)\rightarrow\left(\exists w\text{ }\forall x\in v\text{ }\exists y\in w\text{ }\varphi(x,y,\overline{z})\right).$
When X is a new predicate, we write $\mathcal{L}_{\mathrm{set}}(\mathrm{X})$ instead of $\mathcal{L}_{\mathrm{set}}\cup\{\mathrm{X\}}$, and similarly we write $\mathrm{ZF}(\mathrm{X})$, $\mathrm{Sep}(X)$, $\mathrm{Coll}(X)$, etc. instead of $\mathrm{ZF}(\mathcal{L})$, $\mathrm{Sep}(\mathcal{L})$, $\mathrm{Coll}(\mathcal{L})$, etc. (respectively) when $\mathcal{L}=\mathcal{L}_{\mathrm{set}}(\mathrm{X})$.
(b) For $n\in\omega$, we employ the common notation ($\Sigma_{n},$ $\Pi_{n},$ $\Delta_{n}$) for the Levy hierarchy of $\mathcal{L}_{\mathrm{set}}$-formulae, as in the standard references in modern set theory by Kunen [Ku], Jech [J] and Kanamori [Kan]. In particular, $\Delta_{0}=\Sigma_{0}=\Pi_{0}$ corresponds to the collections of $\mathcal{L}_{\mathrm{set}}$-formulae all of whose quantifiers are bounded. In addition, we will also consider the Takahashi hierarchy of formulae ($\Sigma_{n}^{\mathcal{P}},$ $\Pi_{n}^{\mathcal{P}},$ $\Delta_{n}^{\mathcal{P}}$) (introduced in [Takahashi], and further studied by Mathias’ [M]), where $\Delta_{0}^{\mathcal{P}}$ is the smallest class of $\mathcal{L}_{\mathrm{set}}$-formulae that contains all atomic formulae, and is closed both under Boolean connectives and under quantification in the form $Qx\in y$ and $Qx\subseteq y$ where $x$ and $y$ are distinct variables, and $Q$ is $\exists$ or $\forall$. The classes $\Sigma_{1}^{\mathcal{P}},\Pi_{1}^{\mathcal{P}}$, etc. are defined inductively from the class $\Delta_{0}^{\mathcal{P}}$ in the same way that the formula-classes $\Sigma_{1},\Pi_{1}$, etc. are defined from $\Delta_{0}$-formulae.
(c) For $n\in\omega$. and $\mathcal{L}\supseteq\mathcal{L}_{\mathrm{set}},$ the Levy hierarchy can be naturally extended to $\mathcal{L}$-formulae ($\Sigma_{n}(\mathcal{L}),$ $\Pi_{n}(\mathcal{L}),$ $\Delta_{n}(\mathcal{L})$), where $\Delta_{0}(\mathcal{L})$ is the smallest family of $\mathcal{L}$-formulae that contains all atomic $\mathcal{L}$-formulae and is closed under Boolean operations and bounded quantification.
(d) $\mathrm{KPR}$ (Kripke-Platek with ranks) is the subtheory of $\mathrm{ZF}$ whose axioms consist of KP plus an axiom that asserts that for all ordinals $\alpha,$$\{x:\mathrm{rank}(x)<\alpha\}$ exists (where $\mathrm{rank}(x)$ is the usual rank function on sets). Following recent practice (initiated by Mathias [M]), the foundation scheme of $\mathrm{KP}$ is only limited to $\Pi_{1}$-formulae; thus in this formulation $\mathrm{KP}$ can prove the scheme of $\in$-induction for $\Sigma_{1}$-formulae. In contrast to Barwise’s KP in [Bar], which includes the full scheme of foundation, our version of KP is finitely axiomatizable. Thus $\mathrm{KPR}$ is finitely axiomatizable.
(e) For $\mathcal{L}\supseteq\mathcal{L}_{\mathrm{set}}$, the dependent choice scheme, denoted $\Pi_{\infty}^{1}$-$\mathrm{DC(}\mathcal{L}\mathrm{)}$, consists of $\mathcal{L}$-sentences of the following form:
$\forall z\ \left[\forall x\exists y\ \varphi(x,y,z)\longrightarrow\forall x\exists f\ \left(f:\omega\rightarrow\mathrm{V},\ f(0)=x,\mathrm{and}\ \forall n\in\omega\ \varphi(f(n),f(n+1),z)\right)\right].$
In the presence of $\mathrm{ZF(\mathcal{L}),}$ thanks to the Reflection Theorem, $\Pi_{\infty}^{1}$-$\mathrm{DC(\mathcal{L})}$ is equivalent to the single sentence $\mathrm{DC}$ below:
[…]
Note that $\mathrm{DC}$ is provable in $\mathrm{ZFC}$.
2.2. Definitions and basic facts. (Model theoretic concepts) In what follows we make the blanket assumption that $\mathcal{M}$, $\mathcal{N}$, etc. are $\mathcal{L}$-structures, where $\mathcal{L}\supseteq\mathcal{L}_{\mathrm{set}}$.
(a) We follow the convention of using $M$, $M^{\ast},$ $M_{0}$, etc. to denote (respectively) the universes of structures $\mathcal{M}$, $\mathcal{M}^{\ast},$ $\mathcal{M}_{0},$ etc. We denote the membership relation of $\mathcal{M}$ by $\in^{\mathcal{M}}$; thus an $\mathcal{L}_{\mathrm{set}}$-structure $\mathcal{M}$ is of the form $(M,\in^{\mathcal{M}})$.
(b) $\mathrm{Ord}^{\mathcal{M}}$ is the class of “ordinals” of $\mathcal{M}$, i.e.,
$\mathrm{Ord}^{\mathcal{M}}:=\left\{m\in M:\mathcal{M}\models\mathrm{Ord}(m)\right\},$
where $\mathrm{Ord}(x)$ expresses “$x$ is transitive and is well-ordered by $\in$”. More generally, for a formula $\varphi(\overline{x})$ with parameters from $M$, where $\overline{x}=\left(x_{1},\cdot\cdot\cdot,x_{k}\right)$, we write $\varphi^{\mathcal{M}}$ for $\left\{\overline{m}\in M^{k}:\mathcal{M\models\varphi}\left(m_{1},\cdot\cdot\cdot,m_{k}\right)\right\}.$ A subset $D$ of $M^{k}$ is $\mathcal{M}$-definable if it is of the form $\varphi^{\mathcal{M}}$ for some choice of $\varphi.$ We write $\mathbb{\omega}^{\mathcal{M}}$ for the set of finite ordinals (i.e., natural numbers) of $\mathcal{M}$, and $\mathbb{\omega}$ for the set of finite ordinals in the real world, whose members we refer to as metatheoretic natural numbers. $\mathcal{M}$ is said to be $\omega$-standard if $\left(\mathbb{\omega},\mathbb{\in}\right)^{\mathcal{M}}\cong\left(\mathbb{\omega},\mathbb{\in}\right).$ For $\alpha\in\mathrm{Ord}^{\mathcal{M}}$ we often use $\mathcal{M}_{\alpha}$ to denote the substructure of $\mathcal{M}$ whose universe is $\mathrm{V}_{\alpha}^{\mathcal{M}},$ here $\mathrm{V}_{\alpha}$ is defined as usual within $\mathrm{ZF}$ as $\{x:\mathrm{rank}(x)<\alpha\}$, where $\mathrm{rank}(x)=\sup\{\mathrm{rank}(y)+1:y\in x\}.$
(c) For $c\in M$, $\mathrm{Ext}_{\mathcal{M}}(c)$ is the $\mathcal{M}$-extension of $c$, i.e., $\mathrm{Ext}_{\mathcal{M}}(c):=\{m\in M:m\in^{\mathcal{M}}c\}.$
(d) We say that $X\subseteq M$ is coded (in $\mathcal{M)}$ if there is some $c\in M$ such that $\mathrm{Ext}_{\mathcal{M}}(c)=X.$ $X$ is piecewise coded in $\mathcal{M}$ if $X\cap\mathrm{Ext}_{\mathcal{M}}(m)$ is coded for each $m\in M.$
(e) Suppose $\mathcal{M}\models\mathrm{ZF}$. We say that $A\subseteq M$ is $\mathcal{M}$-amenable if $(\mathcal{M},A)\models\mathrm{ZF}(\mathrm{X})$, here $A$ is the interpretation of $X$.
2.3. Definitions. (“geometric shapes” of extensions). Suppose $\mathcal{L}\supseteq\mathcal{L}_{\mathrm{set}}$, $\mathcal{M}$ and $\mathcal{N}$ are $\mathcal{L}$-structures, and $\mathcal{M}$ a submodel [^4] The notion of a submodel here is as in Model Theory; i.e., $\mathcal{M}\subseteq\mathcal{N}$ means $M\subseteq N$ and the interpretation of each nonlogical symbol of $\mathcal{L}$ in $\mathcal{M}$ is the restriction to $M$ of the corresponding of interpretation in $\mathcal{N}$ (in particular, $\mathrm{Ext}_{\mathcal{M}}(a)\subseteq\mathrm{Ext}_{\mathcal{N}}(a)$ for all $a\in M).$ of $\mathcal{N}$ (written $\mathcal{M}\subseteq\mathcal{N)}.$
(a) $\mathcal{M}^{\ast}$ is the convex hull of $\mathcal{M}$ in $\mathcal{N}$ if $M^{\ast}=\bigcup\limits_{a\in M}\mathrm{Ext}_{\mathcal{N}}(a).$
(b) Suppose $a\in M.$ $\mathcal{N}$ fixes $a$ if $\mathrm{Ext}_{\mathcal{M}}(a)=\mathrm{Ext}_{\mathcal{N}}(a)$, and $\mathcal{N}$ enlarges $a$ if $\mathrm{Ext}_{\mathcal{M}}(a)\subsetneq\mathrm{Ext}_{\mathcal{N}}(a).$
(c) $\mathcal{N}$ end extends $\mathcal{M}$ (written $\mathcal{M}\subseteq_{\mathrm{end}}\mathcal{N}$), if $\mathcal{N}$ fixes every $a\in M$.  End extensions are also referred to in the literature as transitive extensions, and in the old days as outer extensions.
(d) $\mathcal{N}$ is a powerset-preserving end extension of $\mathcal{M}$ (written $\mathcal{M}\subseteq_{\mathrm{end}}^{\mathcal{P}}\mathcal{N}$) if $(i)$ $\mathcal{M}\subseteq_{\mathrm{end}}\mathcal{N}$, and $(ii)$ for if $a\in M$, $b\in N$, and $\mathcal{N}\models(b\subseteq a)$, then $b\in M$.
(e) $\mathcal{N}$ is a rank extension of $\mathcal{N}$ (written $\mathcal{M}\subseteq_{\mathrm{rank}}\mathcal{N}$) if $\mathcal{N}$ is an end extension of $\mathcal{M}$, and for all $a\in M$ and all $b\in N\backslash M$, $\mathcal{N}\models\mathrm{rank}(a)\in\mathrm{rank}(b)$. Here we assume that $\mathcal{M}$ and $\mathcal{N}$ are models of a sufficient fragment of $\mathrm{ZF}$ (such as $\mathrm{KP}$) in which the rank function is well-defined. $\mathcal{N}$ is a topped rank extension of $\mathcal{M}$ if $\mathrm{Ord}^{\mathcal{N}}\backslash\mathrm{Ord}^{\mathcal{M}}$ has a least element. [^5] Recently rank extensions are also referred to as top extensions, we will not use this terminology as it leads to the expression “a topped top extension”.
(f) $\mathcal{N}$ is a cofinal extension of $\mathcal{M}$ (written $\mathcal{M}\subseteq_{\mathrm{cof}}\mathcal{N)}$ if for every $b\in N$ there is some $a\in M$ such that $b\in\mathrm{Ext}_{\mathcal{N}}(a)$, i.e., $\mathcal{N}$ is the convex hull of $\mathcal{M}$ in $\mathcal{N}$.
(g) $\mathcal{N}$ is taller than $\mathcal{M}$ if there is some $b\in M$ such that $M\subseteq\mathrm{Ext}_{\mathcal{N}}(b).$ For models of $\mathrm{ZF}$ (or even $\mathrm{KPR}$) this is equivalent to the existence of an “ordinal” in $\mathrm{Ord}^{\mathcal{N}}$ that exceeds all the “ordinals” in $\mathrm{Ord}^{\mathcal{M}}.$
2.4. Definitions. (“logical shapes” of extensions). Suppose $\mathcal{L}\supseteq\mathcal{L}_{\mathrm{set}}$, and $\mathcal{M}$ and $\mathcal{N}$ are $\mathcal{L}$-structures such that $\mathcal{M}$ is a submodel of $\mathcal{N}$.
[…]
(b) $\mathcal{N}$ is a self-extension of $\mathcal{M}$ (written $\mathcal{M}\subseteq_{\mathrm{self}}\mathcal{N}$) if $(i)$ and $(ii)$ below hold:
$(i)$ $\mathcal{N}$ is interpretable in $\mathcal{M}$, i.e., if $\mathcal{N}$ is isomorphic to a structure $\mathcal{N}^{\ast}$ whose universe $N^{\ast}$, together with the interpretation $R^{\mathcal{N}^{\ast}}$ of each $R\in\mathcal{L}$, are all $\mathcal{M}$-definable.
$(ii)$ There is an $\mathcal{M}$-definable embedding of $\mathcal{M}$ into $\mathcal{N}$.
$\mathcal{N}$ is an elementary self-extension of $\mathcal{M}$ if in $(ii)$ above, the map $j$ is an elementary embedding.
(c) $\mathcal{N}$ is a conservative extension of $\mathcal{M}$ (written $\mathcal{M}\subseteq_{\mathrm{cons}}\mathcal{N}$) if for every $\mathcal{N}$-definable $D$, $M\cap D$ is $\mathcal{M}$-definable.
(d) $\mathcal{N}$ is a faithful extension of $\mathcal{M}$ (written $\mathcal{M}\subseteq_{\mathrm{faith}}\mathcal{N)}$ if for every $\mathcal{N}$-definable $D$, $M\cap D$ is $\mathcal{M}$-amenable (as in Definition 2.2(e)). [^6] The notion of a faithful extension was first introduced and studied in the context of elementary extensions in [En-3].
2.5. Remark. Suppose $\mathcal{M}$ and $\mathcal{N}$ are $\mathcal{L}_{\mathrm{set}}$-structures with $\mathcal{M}\subseteq\mathcal{N}\mathrm{.}$ It is clear that we have the following chain of implications:
$\left(\mathcal{M}\subseteq_{\mathrm{self}}\mathcal{N}\right)\Rightarrow\left(\mathcal{M}\subseteq_{\mathrm{cons}}\mathcal{N}\right)\Rightarrow\left(\mathcal{M}\subseteq_{\mathrm{faith}}\mathcal{N}\right).$
[…]
It is also noteworthy that the most common examples of definable extensions in the context of models of set theory are elementary extensions. However, the Boolean valued approach to forcing provides a wealth of definable extensions that are not elementary. More specifically, if $\mathcal{M}$ is a model of $\mathrm{ZFC}$, $\mathbb{B}$ is a complete Boolean algebra from the point of view of $\mathcal{M}$, and $\mathcal{U}$ is an ultrafilter on $\mathbb{B}$ that is in $\mathcal{M}$, then there is an $\mathcal{M}$-definable embedding $j$ of $\mathcal{M}$ into the model $\mathcal{M}^{\mathbb{B}}/\mathcal{U}$ (where $\mathcal{M}^{\mathbb{B}}$ is the $\mathbb{B}$-valued forcing extension of the universe, as calculated in $\mathcal{M}$). However, in typical cases $\mathcal{M}$ and $\mathcal{M}^{\mathbb{B}}/\mathcal{U}$ are not elementarily equivalent, e.g., we can always arrange $\mathbb{B}$ so that the truth-value of the continuum hypothesis changes in the transition between the $\mathcal{M}$ and $\mathcal{M}^{\mathbb{B}}/\mathcal{U}$. Furthermore, even though $\mathcal{M}$ and $\mathcal{M}^{\mathbb{B}}/\mathcal{U}$ might be elementarily equivalent in some cases, it can be verified that in general, the embedding $j$ is not elementary if $\mathbb{B}$ is a nontrivial notion of forcing.
2.6. Remark. Suppose $\mathcal{M}$ and $\mathcal{N}$ are $\mathcal{L}_{\mathrm{set}}$-structures with $\mathcal{M}\subseteq\mathcal{N}\mathrm{.}$ The following are readily verifiable from the definitions involved:
(a) For each $n\in\omega,$ $\mathcal{M}\preceq_{\Sigma_{n}}\mathcal{N}$ iff $\mathcal{M}\preceq_{\Pi_{n}}\mathcal{N}.$
(b) If $\mathcal{M}\subseteq_{\mathrm{end}}\mathcal{N}$, then $\mathcal{M}\preceq_{\Delta_{0}}\mathcal{N}$; thus $\Delta_{1}$-notions are absolute between $\mathcal{M}$ and $\mathcal{N}$, and $\Sigma_{1}$-formulae with parameters from $M$ are preserved in the passage from $\mathcal{M}$ to $\mathcal{N}$.
(c) Suppose $\mathcal{M}\preceq_{\Delta_{0}}\mathcal{N}$, $\mathcal{N}$ fixes $a\in M$, and there is bijection $f$ in $\mathcal{M}$ between $a$ and some $b\in M$. Then $\mathcal{N}$ fixes $b$ as well (by $\Delta_{0}$-elementarity, $f$ remains a bijection between $a$ and $b$ as viewed from $\mathcal{N}$).
(d) Suppose $\mathcal{M}$ and $\mathcal{N}$ are models of $\mathrm{KPR}$, $\mathcal{M}\preceq_{\Delta_{0}^{\mathcal{P}}}\mathcal{N}$, and $\alpha\in\mathrm{Ord}^{M}$. Then $\mathrm{V}_{\alpha}^{\mathcal{M}}=\mathrm{V}_{\alpha}^{\mathcal{N}}$; this follows from the fact that the formula expressing that $x$ is an ordinal and $y=\mathrm{V}_{x}$ as a $\Sigma_{1}^{\mathcal{P}}$-formula in $\mathrm{KPR}$ (using the usual recursive definition of the $\mathrm{V}_{\alpha}$-hierarchy).
(e) If $\mathcal{M}\subseteq_{\mathrm{end}}^{\mathcal{P}}\mathcal{N}$, and $\mathcal{M}$ and $\mathcal{N}$ both satisfy the axioms of extensionality and powerset, then $\mathcal{M}\preceq_{\Delta_{0}}\mathcal{N}$ (as noted by Forster and Kaye [FS]). Since the formula expressing that $x$ is an ordinal and $y=\mathrm{V}_{x}$ can be written as a $\Sigma_{1}^{P}$-formula, it is absolute between models $\mathcal{M}$ and $\mathcal{N}$ such that $\mathcal{M}\subseteq_{\mathrm{end}}^{\mathcal{P}}\mathcal{N}$. Thus for models $\mathcal{M}$ and $\mathcal{N}$ of $\mathrm{KPR}$, if $\mathcal{M}\subseteq_{\mathrm{end}}^{\mathcal{P}}\mathcal{N}$, then $\mathcal{M}\subseteq_{\mathrm{rank}}\mathcal{N}$.
(f) If $\mathcal{M}\preceq_{\Delta_{0}}\mathcal{N}$, and $\mathcal{M}^{\ast}$ is the convex hull of $\mathcal{M}$ in $\mathcal{N}$, then $\mathcal{M}\subseteq_{\mathrm{cof}}\mathcal{M}^{\ast}\subseteq_{\mathrm{end}}\mathcal{N}$. If $\mathcal{M}$ and $\mathcal{N}$ are furthermore assumed to be models of $\mathrm{KPR}$, and $\mathcal{M}\preceq_{\Delta_{0}^{\mathcal{P}}}\mathcal{N}$, then (e) above implies that $\mathcal{M}\subseteq_{\mathrm{cof}}\mathcal{M}^{\ast}\subseteq_{\mathrm{rank}}\mathcal{N}$.
(g) If both $\mathcal{M}$ and $\mathcal{N}$ are models of $\mathrm{KPR}$, and $\mathcal{M\preceq_{\Delta_{0}}N}$, then the statement “$\mathcal{N}$ is taller than $\mathcal{M}$” is equivalent to “there is some $\gamma\in\mathrm{Ord}^{\mathcal{N}}$ that exceeds each $\alpha\in\mathrm{Ord}^{\mathcal{M}}$”.
(h) A conservative extension is a faithful extension. The converse fails: Consider strongly inaccessible cardinals $\kappa$ and $\lambda$ with $\kappa<\lambda$, then $\left(\mathrm{V}_{\kappa},\in\right)\subsetneq_{\mathrm{faith}}\left(\mathrm{V}_{\lambda},\in\right)$ but by Tarski’s undefinability of truth $\left(\mathrm{V}_{\lambda},\in\right)$ is not a conservative extension of $\left(\mathrm{V}_{\kappa},\in\right)$ since the satisfaction predicate of $\left(\mathrm{V}_{\kappa},\in\right)$ is definable in $\left(\mathrm{V}_{\lambda},\in\right).$
(i) Generic extensions of models of ZF are end extensions but not rank extensions since they have the same ordinals. However, by (f) above, powerset preserving end extensions of models of KPR are rank extensions, thus for models $\mathcal{M}$ and $\mathcal{N}$ of $\mathrm{KPR}$ we have:
[…]
To see that the converse of the above implication fails, consider strongly inaccessible cardinals $\kappa$ and $\lambda$ with $\kappa<\lambda$ and $\kappa=$ the first strongly inaccessible cardinals. Then $\left(\mathrm{V}_{\kappa},\in\right)$ satisfies the $\Pi_{1}$-statement expressing “there is no strongly inaccessible cardinal”, but $\left(\mathrm{V}_{\lambda},\in\right)$ satisfies the negation of the same statement.
(j) As we shall see in Lemma 4.3 faithful end extensions of models of $\mathrm{ZF}$ are rank extensions.
(k) Thanks to the provability of the induction scheme (over natural numbers) in $\mathrm{ZF}$, if $\mathcal{N}$ is a faithful extension of a model $\mathcal{N}$ (where $\mathcal{M}$ is a model of $\mathrm{ZF}$), then $\mathcal{N}$ fixes each element of $\omega^{\mathcal{M}}$ (but $\mathcal{N}$ need not fix $\omega^{\mathcal{M}}$, e.g., consider the case when $\mathcal{N}$ is an internal ultrapower of $\mathcal{M}$ modulo a nonprincipal ultrafilter over $\omega^{\mathcal{M}}$).
2.7. Definition. Reasoning within $ZF$, for each object $a$ in the universe of sets, let $c_{a\text{ }}$ be a constant symbol denoting $a$ (where the map $a\mapsto c_{a}$ is $\Delta_{1}).$ For each finite extension $\mathcal{L}$ of $\mathcal{L}_{\mathrm{Set}}$, let $\mathrm{Sent}_{\mathcal{L}^{+}}(x)$ be the $\mathcal{L}_{\mathrm{Set}}$-formula that defines the class $\mathrm{Sent}_{\mathcal{L}^{+}}$ of sentences in the language $\mathcal{L}^{+}=$$\mathcal{L}\cup\{c_{a}:a\in\mathrm{V}\}$, and let $\mathrm{Sent}_{\mathcal{L}^{+}}(i,x)$ be the $\mathcal{L}_{\mathrm{set}}$-formula that expresses “$i\in\omega,$ $x\in\mathrm{Sent}_{\mathcal{L}^{+}},$ and $x$ is a $\Sigma_{i}(\mathcal{L}^{+})$-sentence”. Suppose $\mathcal{L}$ is a finite extension of $\mathcal{L}_{\mathrm{Set}}$, $\mathcal{M}\models\mathrm{ZF}\mathcal{(L)}$, $S\subseteq M$, and $k\in\omega^{\mathcal{M}}$.
[…]
$(i)\ \ \left[\left(S\left(c_{x}=c_{y}\right)\leftrightarrow x=y\right)\wedge\left(S\left(c_{x}\in c_{y}\right)\leftrightarrow x\in y\right)\right].$
$(ii)\ \ \left[\mathrm{Sent}_{\mathcal{L}^{+}}(k,\varphi)\wedge\left(\varphi=\lnot\psi\right)\right]\rightarrow\left[S(\varphi)\leftrightarrow\lnot S\mathsf{(}\psi\mathsf{)}\right]\mathsf{.}$
$(iii)$ $\ \left[\mathrm{Sent}_{\mathcal{L}^{+}}(k,\varphi)\wedge\left(\varphi=\psi_{1}\vee\psi_{2}\right)\right]\rightarrow\left[S(\varphi)\leftrightarrow\left(S(\psi_{1})\vee S(\psi_{2})\right)\right]\mathsf{.}$
$(iv)$ $\ \left[\mathrm{Sent}_{\mathcal{L}^{+}}(k,\varphi)\wedge\left(\varphi=\exists v\ \psi(v)\right)\right]\rightarrow\left[S(\varphi)\leftrightarrow\exists x\ S\mathsf{(}\psi\mathsf{(}c_{x}\mathsf{))}\right].$
(b) $S$ is a $\Sigma_{\omega}$-satisfaction class for $\mathcal{M}$ if for each $k\in\omega,$ $S$ is a $\Sigma_{k}$-satisfaction class over $\mathcal{M}$.
2.8. Tarski’s Definability and Undefinability of Truth TheoremsSuppose $\mathcal{M}$ is an $\mathcal{L}_{\mathrm{set}}$-structure.
(a) (Definability of Truth) If $\mathcal{M}$ is coded as an element of a model $\mathcal{N}$ of a sufficiently strong [^8] The two canonical fragments of ZF that are capable of defining Tarskian truth predicates for their internal set structures are KP (see [Bar, III.1]; the proof makes it clear that only $\Pi_{1}$-Foundation is invoked) and the fragment of Zermelo set theory in which the separation scheme is limited to $\Delta_{0}$-formulae(see [M, Proposition 3.10]). fragment of $\mathrm{ZF}$, then for some $\mathcal{N}$-coded $S$, $S$ is a $\Sigma_{\omega}$-satisfaction class for $\mathcal{M}$ (we refer to $S$ as the Tarskian satisfaction class for $\mathcal{M}$, as defined in $\mathcal{N}$).
(b) (Undefinability of Truth) If $S$ is a $\Sigma_{\omega}$-satisfaction class for $\mathcal{M}$, then $S$ is not $\mathcal{M}$-definable
It is a well-known result of Levy that if $\mathcal{M}\models\mathrm{ZF,}$ then there is a $\Delta_{0}$-satisfaction class for $\mathcal{M}$ that is definable in $\mathcal{M}$ both by a $\Sigma_{1}$-formula and a $\Pi_{1}$-formula (see [J, p. 186] for a  proof). This makes it clear that for each $n\geq 1,$ there is a $\Sigma_{n}$-satisfaction class for $\mathcal{M}$ that is definable in $\mathcal{M}$ by a $\Sigma_{n}$-formula. Levy’s result extends to models of $\mathrm{ZF(}\mathcal{L}\mathrm{)}$ if $\mathcal{L}$ is finite as  follows:
[…]
The Reflection Theorem is often formulated as a theorem scheme of $\mathrm{ZF}$ (e.g., as in [J]), the proof strategy of the Reflection Theorem applies equally well to $\mathrm{ZF}(\mathcal{L})$ for finite extensions $\mathcal{L}$ of $\mathcal{L}_{\mathrm{set}}$.
2.10. Montague-Vaught-Levy Reflection Theorem. Let $\mathcal{L}$ be a finite extension of $\mathcal{L}_{\mathrm{set}}.$ For each $n\in\omega$, $\mathrm{ZF(}\mathcal{L}\mathrm{)}$ proves that there are arbitrarily large ordinals $\gamma$ such that the submodel of the universe determined by $V$γ is a $\Sigma_{n}(\mathcal{L})$-elementary submodel of the universe. [^9] This is expressible with the help of $\mathrm{Sat}_{\Sigma_{n}(\mathcal{L})}.$
2.11. $\mathbf{\Sigma}_{\mathbf{n}}$-Elementary Chains Theorem. Suppose $n\in\omega,$ $\mathcal{L}\supseteq\mathcal{L}_{\mathrm{Set}}$, $(I,<_{I})$ is a linear order, and $\left\{\mathcal{M}_{i}:i\in I\right\}$ is a collection of $\mathcal{L}$-structures such that $\mathcal{M}_{i}\ \mathcal{\preceq}_{\Sigma_{n}(\mathcal{L})}\ \mathcal{M}_{i^{\prime}}$ whenever $i<_{I}i^{\prime}.$ Then we have:
$\mathcal{M}_{i}\ \mathcal{\preceq}_{\Sigma_{n}(\mathcal{L})}\ \bigcup\limits_{i\in I}\mathcal{M}_{i}.$
The following is a special case of a result of Gaifman [G-1, Theorem 1, p.54].
2.12. Gaifman Splitting Theorem. Let $\mathcal{L}\supseteq\mathcal{L}_{\mathrm{Set}}.$ Suppose $\mathcal{M}\models\mathrm{ZF}(\mathcal{L})$, $\mathcal{N}$ is an $\mathcal{L}$-structure with $\mathcal{M}\preceq_{\Delta_{0}(\mathcal{L})}\mathcal{N}$, and $\mathcal{M}^{\ast}$ is the convex hull of $\mathcal{M}$ in $\mathcal{N}$. Then the following hold:
(a) $\mathcal{M}\preceq_{\mathrm{cof}}\mathcal{M}^{\ast}\preceq_{{}_{\Sigma_{0}(\mathcal{L}),\mathrm{end}}}\mathcal{N}$.
(b) If $\mathcal{M\preceq N}$, then $\mathcal{M}\preceq_{\mathrm{cof}}\mathcal{M}^{\ast}\preceq_{{}_{\mathrm{end}}}\mathcal{N}$.
Proof. To simplify the notation, we present the proof for $\mathcal{L}=\mathcal{L}_{\mathrm{set}}$ since the same reasoning handles the general case. To prove (a) suppose $\mathcal{M}\prec_{\Delta_{0},\mathrm{cof}}\mathcal{N},$ where $\mathcal{M}\models\mathrm{ZF}\mathsf{.}$ Note that we are only assuming that $\mathcal{N}$ is an $\mathcal{L}_{\mathrm{set}}$-structure such that $\mathcal{M}$ is a submodel of $\mathcal{N}$, and we are not assuming that $\mathcal{N}\models\mathrm{ZF}\mathsf{.}$ The definition of $\mathcal{M}^{\ast}$ makes it clear that $\mathcal{M}$ is cofinal in $\mathcal{N}$. Our goal is to show that $\mathcal{M}\prec_{\Sigma_{n}}\mathcal{M}^{\ast}$ for each $n\in\omega,$ so let us fix $n\in\omega.$ Thanks to the Reflection Theorem there is an unbounded subset $U$ of $\mathrm{Ord}^{\mathcal{M}}$ such that:
(1)   If $\alpha$ and $\beta$ are in $U$ with $\alpha<\beta$, then $\mathcal{M}\models\mathrm{V}_{\alpha}\prec_{\Sigma_{n}}\mathrm{V}_{\beta}.$
Let $\mathrm{Sat}(s,y)$ be the $\mathcal{L}_{\mathrm{set}}$-formula expressing “$s$ is the Tarskian satisfaction predicate for $(y,\in)$”. $\mathrm{Sat}(s,y)$ is $\Delta_{0}.$ Since in $\mathrm{ZF}$ for each $y$ there is a (unique) $s$ such that $\mathrm{Sat}(s,y)$ holds, this makes it clear that the statement expressing $(x,\in)\prec_{\Sigma_{n}}(y,\in)$ can be written as a $\Delta_{0}$-statement in parameters $x$, $y$, and $s$. Thus,
(2)   If $m_{1}$ and $m_{2}$ are elements of $M$ such that $\mathcal{M}\models(m_{1},\in)\prec_{\Sigma_{n}}(m_{2},\in)$, and $\mathcal{M}\prec_{\Delta_{0}}\mathcal{M}^{\ast}$, then $\mathcal{M}^{\ast}\models(m_{1},\in)\prec_{\Sigma_{n}}(m_{2},\in).$
Let $\alpha\in\mathrm{Ord}^{\mathcal{M}},$ let $a_{\alpha}\in M$ such that $\mathcal{M}\models a_{\alpha}=\mathrm{V}_{\alpha}$,  using (1) and (2) we can conclude:
(3)   If $\alpha$ and $\beta$ are in $U$ with $\alpha<\beta$, then $\left(a_{\alpha},\in\right)^{\mathcal{M}^{\ast}}\prec_{\Sigma_{n}}\left(a_{\beta},\in\right)^{\mathcal{M}^{\ast}}.$
On the other hand, by putting $\mathcal{M}\prec_{\mathrm{cof}}\mathcal{M}^{\ast}$ together with $\mathcal{M}=\bigcup\limits_{\alpha\in U}(a_{\alpha},\in)^{\mathcal{M}}$ we have:
(4) $\ \ \mathcal{M}^{\ast}=\bigcup\limits_{\alpha\in U}(a_{\alpha},\in)^{\mathcal{M}^{\ast}}$.
Thanks to (3), (4), and the Elementary Chains Theorem we conclude:
(5)   If $\alpha$ is in $U$, then $(a_{\alpha},\in)^{\mathcal{M}^{\ast}}\prec_{\Sigma_{n}}\mathcal{M}^{\ast}.$
Thus, given a $\Sigma_{n}$-formula $\varphi(\overline{x})$ such that $\mathcal{M}\models\varphi(\overline{p})$ for some sequence $\overline{p}$ of parameters from $M$, there is a large enough $\alpha\in U$ such that:
$\mathcal{M}\models\overline{p}\in a_{\alpha}\wedge\varphi^{a_{\alpha}}(\overline{p}).$
Since the statement $\varphi^{a_{\alpha}}(\overline{p})$ is $\Delta_{0}$, $\mathcal{M}^{\ast}\models\varphi^{a_{\alpha}}(\overline{p})$, i.e., $(a_{\alpha},\in)^{\mathcal{M}^{\ast}}\models\varphi(\overline{p})$, which together with (5) implies that $\mathcal{M}^{\ast}\models\varphi(\overline{a}).$ This concludes the proof of (a).
(b) follows easily from (a).
$\square$
2.13. Remark. In general, if $\mathcal{M}\subseteq\mathcal{N}$, where $\mathcal{M}$ and $\mathcal{N}$ are both models of $\mathrm{ZF}$, the condition “$\mathrm{Ord}^{\mathcal{M}}$ is cofinal in $\mathrm{Ord}^{\mathcal{N}}$” does not imply that $\mathcal{M}$ is cofinal in $\mathcal{N}$ (for example consider $\mathcal{M}$ and $\mathcal{N}$ where $\mathcal{N}$ is a set-generic extension of $\mathcal{M}$). However, the conjunction of $\Sigma_{1}$-elementarity and “$\mathrm{Ord}^{\mathcal{M}}$ is cofinal in $\mathrm{Ord}^{\mathcal{N}}$”, where $\mathcal{M}\models\mathrm{ZF}$ (or even $\mathrm{KPR}$) and $\mathcal{N}$ is an $\mathcal{L}_{\mathrm{set}}$-set structure implies that $\mathcal{M}$ is cofinal in $\mathcal{N}$, since the statement $(y\in\mathrm{Ord}\wedge x=\mathrm{V}_{y})$ is $\Pi_{1}$, and therefore it is absolute between $\mathcal{M}$ and $\mathcal{N}$ if $\mathcal{M}\prec_{\Sigma_{1}}\mathcal{N}.$ [^10] It is well-known that for models $\mathcal{M}$ and $\mathcal{N}$ of $\mathrm{ZF}$, $\mathcal{M}\prec_{\Sigma_{1}}\mathcal{N}$ iff $\mathcal{M}\prec_{\Delta_{0}^{\mathcal{P}}}\mathcal{N}$. For a refinement of the nontrivial direction of this result, see [M, Proposition 6.17]. This is how the existence of elementary embedding between inner models of $\mathrm{ZF}$ is usually formalized, as in Kanamori’s monograph [Kan, Proposition 5.1].
Recall that for models of $\mathrm{ZF}$, the $\mathcal{L}_{\mathrm{set}}$-sentence $\exists p\left(\mathrm{V}=\mathrm{HOD}(p)\right)$ expresses: “there is some $p$ such that every set is first order definable in some structure of the form $\left(\mathrm{V}_{\alpha},\in,p,\beta\right)_{\beta<\alpha}$ with $p\in\mathrm{V}_{\alpha}$”. The following result is classical.
2.14. Myhill-Scott Theorem. The following statements are equivalent for $\mathcal{M}\models\mathrm{ZF}$:
(a) $\mathcal{M}\models\exists p\left(\mathrm{V}=\mathrm{HOD}(p)\right).$
(b) $\mathcal{M}$ carries a definable global well-ordering, i.e., for some $p\in M$ and some set-theoretic formula $\varphi(x,y,z)$, $\mathcal{M}$ satisfies $\varphi(x,y,p)$ well-orders the universe”.
(c) $\mathcal{M}$ carries a definable global choice function, i.e., for some $p\in M$ and some set-theoretic formula $\psi(x,y,z)$, $\mathcal{M}$ satisfies $\psi(x,y,p)$ is the graph of a global choice function”.
3. TALLER CONSERVATIVE ELEMENTARY EXTENSIONS
3.1. Theorem. Every model $\mathcal{M}$ of $\mathrm{ZF}+\exists p\left(\mathrm{V}=\mathrm{HOD}(p)\right)$ (of any cardinality) has a conservative elementary extension $\mathcal{N}$ such that $\mathcal{N}$ is taller than $\mathcal{M}$.
We shall present two proofs of Theorem 3.1, the first is based on a class-sized syntactic construction taking place within a model of set theory; it was inspired by Kaufmann’s proof of the MacDowell-Specker theorem using the Arithmetized Completeness Theorem, as described by Schmerl [Schm-1]. The second one is based on a model-theoretic argument that is reminiscent of the ultrapower proof of the McDowell-Specker theorem (as in [En-5]). The first proof is short and devilish; the second proof is a bit longer but is more transparent due to its combinatorial flavor; it also lends itself to fine-tuning, as indicated in Remark 3.3 and Theorem 3.4.
First proof of Theorem 3.1. We need the following two Facts 1 and 2 below.
[…]
Proof of Fact 1. Since $\mathcal{M}$ has a definable global well-ordering, the Henkin proof of the completeness theorem of first order logic can be applied within $\mathcal{M}$ to construct a Henkinized complete extension $T^{\mathrm{Henkin}}$ of $T$ (in a language extending the language of $T$ by class-many new constant symbols) such that $T$ is definable in $\mathcal{M}$. This in turn allows $\mathcal{M}$ to define $\mathcal{N}$ by reading it off $T^{\mathrm{Henkin}},$ as in the usual Henkin proof of the completeness theorem.
[…]
Proof of Fact 2. This is an immediate consequence of the Reflection Theorem 2.10.
Starting with a model $\mathcal{M}$ of $\mathrm{ZF}+\exists p\left(\mathrm{V}=\mathrm{HOD}(p)\right),$ by Theorem 2.14, $\mathcal{M}$ carries a global definable well-ordering. We will construct an increasing sequence of $\mathcal{L}_{\mathrm{set}}$-structures $\left\langle\mathcal{N}_{k}:k\in\omega\right\rangle$ that satisfies the following properties for each $k\in\omega$:
(1) $\ \ \mathcal{N}_{0}=\mathcal{M}.$
(2) $\ \ \mathcal{M}\preceq_{\Pi_{k+2}}\mathcal{N}_{k}\preceq_{\Pi_{k+1}}\mathcal{N}_{k+1}$.
(3)   There is an ordinal in $\mathcal{N}_{1}$ that is above all of the ordinals of $\mathcal{M}$.
(4) $\ \ \mathcal{N}_{k}$ is a conservative extension of $\mathcal{M}$.
Thus
$\mathcal{M=N}_{0}\preceq_{\Pi_{1},\ \mathrm{cons,\ taller}}\mathcal{N}_{1}\preceq_{\Pi_{2},\ \mathrm{cons}}\mathcal{N}_{2}\preceq_{\Pi_{3},\ \mathrm{cons}}\mathcal{N}_{3,\ \mathrm{cons}}\cdot\cdot\cdot$
and for each $k\in\omega$, $\mathcal{M}\preceq_{\Pi_{n+2},\ \mathrm{cons}}\mathcal{N}_{k}.$ This shows that with the choice of $\mathcal{N}:=\bigcup\limits_{n\in\omega}\mathcal{N}_{n}$, we have $\mathcal{M}\prec_{\mathrm{cons}}\mathcal{N}$; and $\mathcal{N}$ is taller than $\mathcal{M}$. So the proof will be complete once we explain how to recursively build the desired chain $\left\langle\mathcal{N}_{k}:k\in\omega\right\rangle.$
Using the notation of Definition 2.7, let $\mathcal{L}_{\mathrm{set}}^{+}=$ $\mathcal{L}_{\mathrm{set}}\cup\{c_{a}:a\in\mathrm{V}\}$ be the language $\mathcal{L}_{\mathrm{set}}$ of set theory with a constant $c_{a}$ for each $a\in\mathrm{V}$. To build $\mathcal{N}_{1}$ we argue within $\mathcal{M}$: add a new constant $c$ to $\mathcal{L}_{\mathrm{set}}^{+}$ and consider the theory $T_{1}$ defined as follows:
$T_{1}=\left\{\mathrm{Ord}(c)\right\}\cup\{c_{\alpha}\in c:\mathrm{Ord}(\alpha)\}\cup\mathrm{Th}_{\Pi_{3}}($$V$$,\in,c_{a})_{a\in\mathrm{V}}.$
Thus $T_{1}$ is a proper class within $\mathcal{M}$. Arguing within $\mathcal{M}$, and using Fact 2, it is easy to see that $\mathrm{Con}(T_{1})$ holds, and therefore by Fact 1 we can get hold of a proper class model $\mathcal{N}_{1}$ of $T_{1}$. Clearly $\mathcal{M}\prec_{{}_{\Pi_{3}}}\mathcal{N}_{1}$, there is an ordinal in $\mathcal{N}_{1}$ that is above all of the ordinals of $\mathcal{M}$, and $\mathcal{N}_{1}$ is a conservative extension of $\mathcal{M}$ (by the last clause of Fact 1).
Next suppose we have built $\left\langle\mathcal{N}_{i}:1\leq i\leq k\right\rangle$ for some $k\in\omega$ while complying with (1) through (4). Consider the theory $T_{k}$ defined in $\mathcal{M}$ as follows:
$T_{k}=\mathrm{Th}_{\Pi_{k+1}}(\mathcal{N}_{k},c_{b})_{b\in N_{k}}\cup\mathrm{Th}_{\Pi_{k+2}}($$V$$,\in,c_{a})_{a\in\mathrm{V}}.$
Fact 2 together with the inductive assumption that $\mathcal{M}\prec_{\Pi_{k+1}}\mathcal{N}_{k}$ makes it evident that $\mathrm{Con}(T_{k})$ holds in $\mathcal{M}$. Hence by Fact 1 we can get hold of a model $\mathcal{N}_{k+1}$ of $T$ whose elementary diagram is definable in $\mathcal{M}$. This makes it clear that $\left\langle\mathcal{N}_{i}:1\leq i\leq k+1\right\rangle$ meets (1) through (4). $\square$
Second proof of Theorem 3.1. Given $\mathcal{M}\models\mathrm{ZF}+\exists p\left(\mathrm{V}=\mathrm{HOD}(p)\right)$ we will construct an ultrafilter $\mathcal{U}$ over the $\mathcal{M}$-definable subsets of $\mathrm{Ord}^{\mathcal{M}}$ satisfying the following properties:
(1)   Every member of $\mathcal{U}$ is unbounded in $\mathrm{Ord}^{\mathcal{M}}$.
(2)   Given any $\mathcal{M}$-definable $f:\mathrm{Ord}^{\mathcal{M}}\rightarrow\mathcal{M}$, there is a class $X\in\mathcal{U}$ such that either $f\upharpoonright X$ is one-to-one, or the range of $f\upharpoonright X$ is a set (as opposed to a class).
(3)   Given any $n\in\mathbb{\omega}$, the family $\mathcal{U}_{n}=\{X\in\mathcal{U}:X$ is $\Sigma_{n}$-definable in $\mathcal{M}$} is $\mathcal{M}$-definable in the sense that there is a parametric formula $\psi_{n}(\alpha,\beta)$ such that:
$X\in\mathcal{U}_{n}$ iff $\exists\alpha\in\mathrm{Ord}^{\mathcal{M}}$ $X=\{\beta\in\mathrm{Ord}^{\mathcal{M}}:\mathcal{M}\models\psi_{n}(\alpha,\beta)\}.$
We will build $\mathcal{U}$ in $\omega\times\mathrm{Ord}^{\mathcal{M}}$ stages in the following sense: for each $n\in\omega$ and $\alpha\in\mathrm{Ord}^{\mathcal{M}}$ we shall employ external recursion on $n$ and an internal recursion on $\alpha$ to a parametrically $\mathcal{M}$-definable subset $X_{n,\alpha}$ of $\mathrm{Ord}^{\mathcal{M}}$, and then we can define $\mathcal{U}$ as:
$\mathcal{U}=\{X_{n,\alpha}:n\in\omega,\alpha\in\mathrm{Ord}^{\mathcal{M}}\}\cup\mathcal{U}_{-1},$
where $\mathcal{U}_{-1}$ denote the collection of co-bounded parametrically $\mathcal{M}$-definable subsets of $\mathrm{Ord}^{\mathcal{M}}$, i.e., parametrically $\mathcal{M}$-definable subsets $X$ of $\mathrm{Ord}^{\mathcal{M}}$ such that, as viewed in $\mathcal{M}$, the complement of $X$ is a set (as opposed to a proper class). Let $\left\langle f_{n,,\alpha}:\alpha\in\mathrm{Ord}^{\mathcal{M}}\right\rangle$ denote an enumeration within $\mathcal{M}$ of all $\Sigma_{n}$-functions of $\mathcal{M}$ with domain $\mathrm{Ord}^{\mathcal{M}}$. Note that such an $\mathcal{M}$-enumeration exists because:
(a) By Theorem 2.9 for each $n\in\omega$ there is a definable satisfaction predicate for $\Sigma_{n}$-formulae, and
(b) By Theorem 2.14 $\mathcal{M}$ has a definable global well-ordering.
Suppose for some $n\in\omega$ we have constructed $\mathcal{U}_{n}=\{X_{n,\alpha}:\alpha\in\mathrm{Ord}^{\mathcal{M}}\}$ satisfying the following inductive clauses:
$(i)$ $\ \ \mathcal{U}_{n}$ is a parametrically $\mathcal{M}$-definable family of classes in $\mathcal{M}$, i.e., there is a formula $\psi_{n}(\alpha,\beta)$ (with suppressed parameters) such that:
$X\in\mathcal{U}_{n}$ iff $\exists\alpha$ $X=\{\beta\in\mathrm{Ord}^{\mathcal{M}}:\mathcal{M}\models\psi_{n}(\alpha,\beta)\}.$
$(ii)$ $\ \ \mathcal{U}_{n}$ has the finite intersection property in the sense of $\mathcal{M}\mathfrak{,}$ i.e., if $\psi_{n}(\alpha,\beta)$ is as above, then $\mathcal{M}$ satisfies the sentence below:
For all finite sequences of ordinals $\left\langle\alpha_{i}:i<k\right\rangle$ there is $\beta\in\mathrm{Ord}$ such that $\forall i<k\ \psi_{n}(\alpha_{i},\beta)$.
$(iii)$ $\ \ f_{n,\alpha}\upharpoonright X_{n,\alpha}$ is either one-to-one, or the range of $f_{n,\alpha}\upharpoonright X_{n,\alpha}$ is a set.
$(iv)$ $\ \ \mathcal{U}_{-1}\subseteq\mathcal{U}_{n}.$
We shall now describe a definable sequence $\left\langle X_{n+1,\alpha}:\alpha\in\mathrm{Ord}^{\mathcal{M}}\right\rangle$ which continues to satisfy the clauses $(i)$ through $(iii)$ above, and we define $\mathcal{U}_{n+1}:=\left\{X_{n+1,\alpha}:\alpha\in\mathrm{Ord}^{\mathcal{M}}\right\}.$ Since this part of the construction involves an internal recursion within $\mathcal{M}$ we will argue as if we are inside $\mathcal{M}$ and therefore omit reference to $\mathcal{M}$ itself. Suppose we have built $\left\langle X_{n+1,\beta}:\beta<\alpha\right\rangle$ and let $\mathcal{F}_{n+1,\alpha}$ denote:
$\{X_{m,\delta}:m<n+1,\delta<\alpha\},$
and let $\mathcal{F}_{n+1,\alpha}^{\ast}$ denote the set of all finite intersections of $\mathcal{F}_{n+1,\alpha}$. Also, let $\left\langle Y_{\alpha}:\alpha\in\mathrm{Ord}\right\rangle$ be an enumeration of $\mathcal{F}_{n+1,\alpha}^{\ast}$. We now wish to define a class $X_{n+1,\alpha}$ which “takes care of” $f_{n+1,\alpha}$, and which meets every $Y_{\alpha}$ We distinguish the following two cases:
Case 1: There is an ordinal $\gamma$ such that for every ordinal $\alpha$ we have:
$f_{n+1,\alpha}^{-1}\{\beta:\beta<\gamma\}\cap Y_{\alpha}\neq\varnothing.$
Case 2: Not case 1.
If case 1 holds then we shall simply let $X_{n+1,\alpha}=f_{n+1,\alpha}^{-1}\{\beta:\beta<\gamma_{0}\},$ where $\gamma_{0}$ is the first ordinal witnessing the veracity of case 1.
If case 2 holds then we can define a class $S=\{s_{\alpha}:\alpha\in\mathrm{Ord}\}$ by induction on $\alpha$ as follows: suppose we have already built $S$ up to some ordinal $\gamma$ as $\{s_{\delta}:\delta<\gamma\}$ such that:
$s_{\delta}<s_{\delta^{\prime}}$ and $f(s_{\delta})<f(s_{\delta^{\prime}})$ whenever $\delta<\delta^{\prime}<\gamma$,
and
$s_{\delta}\in Y_{\delta}$ for each $\delta<\gamma.$
Let $\theta=\sup\{f(s_{\delta}):\delta<\gamma\}.$ Note that $\theta$ is well-defined because $\mathcal{M}$ satisfies the replacement scheme. Since we know that case 2 holds there must exist some $X$ in $\mathcal{F}_{n+1,\alpha}^{\ast}$ such that:
$f_{n+1,\alpha}^{-1}\{\beta:\beta<\gamma\}\cap X\neq\varnothing$,
therefore we can let $s_{\gamma\text{ }}$denote the first member of $X\cap Y_{\gamma\text{ }}$that is above $\{s_{\delta}:\delta<\gamma\}$. By conditions $(ii)$ and $(iv),$ $X\cap Y_{\gamma\text{ }}$is unbounded in the ordinals hence making it possible to define $s_{\gamma}$. Having defined the class $S$, let $X_{n+1,\alpha}=S$.
We are now done with the construction of $\left\{X_{n+1,\alpha}:\alpha\in\mathrm{Ord}^{\mathcal{M}}\right\}$. We pause to note that even though $\mathcal{U}$ itself in not definable in $\mathcal{M}$, for each fixed $n\in\omega$, the family $\mathcal{U}_{n}=\{X_{n,\alpha}:\alpha\in\mathrm{Ord}^{\mathcal{M}}\}$ is definable in $\mathcal{M}$, which is exactly why $\mathcal{U}$ satisfies condition (3) promised at the beginning of the proof. As in the Łoś proof of the fundamental theorem of ultraproducts (and thanks to the availability of an $\mathcal{M}$-definable global well-ordering) it is routine to check that the definable ultrapower:
$\mathcal{N}=\mathcal{M}^{\mathrm{Ord}^{\mathcal{M}}}/\ \mathcal{U}$,
consisting of $\mathcal{U}$-equivalence classes of $\mathcal{M}$-definable functions from $\mathrm{Ord}^{\mathcal{M}}$ to $M$ forms an elementary extension of $\mathcal{M}$. Also the construction makes it clear that $\mathcal{N}$ is taller than $\mathcal{M}$ since the equivalence class $[\mathrm{id}]_{U}$ exceeds all ordinals of $\mathcal{M}$ (where $\mathrm{id}$ is the identity map on $\mathrm{Ord}^{\mathcal{M}})$, thanks to the fact that $\mathcal{U}_{-1}\subseteq\mathcal{U}$. To verify conservativity of $\mathcal{N}$ over $\mathcal{M}$, suppose we are given a formula $\varphi(x,y)$ and some parameter $s\in N$, and we wish to show that the set $\{m\in M:\mathcal{N}\models\varphi(m,s)\}$ is $\mathcal{M}$-definable. Choose $f$ so that $s=$ the $\mathcal{U}$-equivalence class of $f=[f]_{\mathcal{U}}$ and choose a natural number $j$ large enough so that the formula $\theta(x,z)=\varphi(x,f(z))$ is a $\Sigma_{j}$-formula. Note that for each $m\in M$ the characteristic function of $S_{m}=\{z:\theta(m,z)\}$ is definable by some $\Sigma_{j}$-function. Hence for every $m\in M$ we have:
$S_{m}\in\mathcal{U}$ iff $S_{m}\in\mathcal{U}_{j},$
Therefore, if $\psi(\alpha,\beta)$ denotes the formula defining $\mathcal{U}_{j}$ in $\mathcal{M}$, then for all $m\in M,$ the following holds:
$\mathcal{N}\models\varphi(m,s)$ iff $\mathcal{M}\models\exists\alpha\forall\beta(\psi_{j}(\alpha,\beta)\rightarrow\gamma(\alpha,\beta)).$
$\square$
3.2. Corollary. Work in $\mathrm{ZF}+\exists p\left(\mathrm{V}=\mathrm{HOD}(p)\right)$ in the metatheory, and let $\mathcal{M}$ be an arbitrary model of $\mathrm{ZF}+\exists p\left(\mathrm{V}=\mathrm{HOD}(p)\right).$
(a) For every regular cardinal $\kappa$ there is a conservative elementary extension $\mathcal{N}_{\kappa}$ of $\mathcal{M}$ that is taller than $\mathcal{M}$ and $\mathrm{cf}(\mathrm{Ord}^{\mathcal{N}})=\kappa$ [^12] As indicated by the proof of this part, $\mathrm{ZFC}$ suffices in the metatheory.
(b) There is a proper class model $\mathcal{N}_{\mathrm{Ord}}$ whose elementary diagram is a definable class, $\mathcal{N}_{\mathrm{Ord}}$ is a conservative elementary extension of $\mathcal{M}$ that is taller than $\mathcal{M}$, and $\mathrm{Ord}^{\mathcal{N}_{\mathrm{Ord}}}$ contains a cofinal class of order-type $\mathrm{Ord}$.
Proof. To establish (a), we observe that given a regular cardinal $\kappa$, Theorem 3.1 allows us to build an elementary chain of models $\left\langle\mathcal{M}_{\alpha}:\alpha\in\kappa\right\rangle$ with the following properties:
(1) $\ \ \mathcal{M}_{0}=\mathcal{M}.$
(2)   For each $\alpha\in\kappa$ $\mathcal{M}_{\alpha+1}$ is a conservative elementary extension of $\mathcal{M}$ that is taller than $\mathcal{M}.$
(3)   For limit $\alpha\in\kappa$, $\mathcal{M}_{\alpha}=\bigcup\limits_{\beta\in\alpha}\mathcal{M}_{\beta}.$
The desired model $\mathcal{N}$ is clearly $\mathcal{N}_{\kappa}=\bigcup\limits_{\alpha\in\kappa}\mathcal{M}_{\alpha}.$ This concludes the proof of part (a).
With part (a) at hand, it is evident that in the presence of a global definable well-ordering (available thanks to the veracity of $\exists p\left(\mathrm{V}=\mathrm{HOD}(p)\right)$ in the metatheory), we can construct the desired model $\mathcal{N}_{\mathrm{Ord}}$ as the union of an elementary chain of models $\left\langle\mathcal{M}_{\alpha}:\alpha\in\mathrm{Ord}\right\rangle$ that satisfies (1) and it also satisfies the result of replacing $\kappa$ with Ord in (2) and (3). $\square$
[…]
3.4. Theorem. Every model of $\mathrm{ZF}+\exists p\left(\mathrm{V}=\mathrm{HOD}(p)\right)$ has a cofinal conservative elementary extension that possesses a minimal elementary end extension.
Proof. Let $\mathcal{M}$ be a model of $\mathrm{ZF}+\exists p\left(\mathrm{V}=\mathrm{HOD}(p)\right)$, $\mathcal{N}$ be the model constructed in the second proof of Theorem 3.1, and let $\mathcal{M}^{\ast}$ be the convex hull of $\mathcal{M}$ in $\mathcal{N}$. Note that by Gaifman Splitting Theorem we have:
$\mathcal{M}\preceq_{\mathrm{cof}}\mathcal{M}^{\ast}\prec_{\mathrm{end}}\mathcal{N}$.
The above together with the fact that $\mathcal{N}$ is conservative extension of $\mathcal{M}$ makes it clear that $\mathcal{M}^{\ast}$ is a conservative extension of $\mathcal{M}$ [^14] By Theorem 4.1, $\mathcal{M}^{\ast}$ is a proper elementary extension of $\mathcal{M}$. We claim that $\mathcal{N}$ is a minimal elementary end extension of $\mathcal{M}^{\ast}$. First note that $\mathcal{N}$ is generated by members of $\mathcal{M}$ and the equivalence class of the identity function $[\mathrm{id}]_{\mathcal{U}}$ via the definable terms of $\mathcal{M}$, i.e., the following holds:
For any $m\in M$, there exists some $\mathcal{M}$-definable function $f$ such that $\mathcal{N}\models f([\mathrm{id}]_{\mathcal{U}})=m.$
To establish our claim it is sufficient to show that if $s\in N\backslash M^{\ast}$ then $[\mathrm{id}]_{\mathcal{U}}$ is definable in $\mathcal{N}$ using parameters from $M\cup\{s\}.$ Choose $h$ such that $s=[h]_{\mathcal{U}}$ and recall that by our construction of $\mathcal{U}$ we know that there exists $X\subseteq Y$ with $X\in\mathcal{U}$ such that either $h\upharpoonright X$ is one-to-one, or the image of $X$ under $h$ is a set (and thus of bounded rank). However, since $s$ was chosen to be in $N\backslash M^{\ast}$ we can rule out the latter possibility. Note that since $X\in\mathcal{U}$ the extension of $X$ in $\mathcal{N}$ will contain $[\mathrm{id}]_{\mathcal{U}}$, i.e., if $X$ is defined in $\mathcal{M}$ by $\varphi(x)$ then $\mathcal{N}\models\varphi([\mathrm{id}]_{\mathcal{U}})$. Moreover, by elementarity, since $f$ is one-to-one on $X$ in $\mathcal{M}$, it must remain so in $\mathcal{N}$. Hence we can define $[\mathrm{id}]_{\mathcal{U}}$ to be the unique element of $h^{-1}(b)\cap X.$ This completes the proof. $\square$
3.5. Remark. As Theorem 4.1 of the next section shows, in Theorem 3.1 the model $\mathcal{N}$ cannot be required to fix $\omega^{\mathcal{M}}$ and in particular $\mathcal{N}$ cannot be arranged to end extend $\mathcal{M}$. By putting this observation together with Remark 2.6(k) we can conclude that in the proof of Corollary 3.2, $\omega^{\mathcal{N}_{\kappa}}$ is $\kappa$-like (i.e., its cardinality is $\kappa$ but each proper initial segment of it is of cardinality less than $\kappa$), and $\omega^{\mathcal{N}_{\mathrm{Ord}}}$ is Ord-like (i.e., it is a proper class every proper initial segment of which is isomorphic to an ordinal).
We next address the question of the extent to which the hypothesis in Theorem 3.1 that $\mathcal{M}$ is a model of $\exists p\left(\mathrm{V}=\mathrm{HOD}(p)\right)$ can be wakened. In what follows $\mathrm{AC}$ stands for the axiom of choice, and $\mathrm{DC}$ stands for the axiom of dependent choice.
3.6. Proposition. Suppose $\mathcal{M}$ and $\mathcal{N}$ are models of $\mathrm{ZF}$, $\mathcal{M}\prec_{\mathrm{cons}}\mathcal{N}$, and $\mathcal{N}$ is taller than $\mathcal{M}$. If every set can be linearly ordered in $\mathcal{N}$, then there is an $\mathcal{M}$-definable global linear ordering.
Proof. Suppose $\mathcal{M}$ and $\mathcal{N}$ are as in the assumptions of the proposition. Fix $\alpha\in\mathrm{Ord}^{\mathcal{N}}$ such that $\alpha$ is above all the ordinals of $\mathcal{M}$ and let $R$ denote a linear ordering of $\mathrm{V}_{\alpha}^{\mathcal{N}}$ in $\mathcal{N}$. By conservativity, there is a formula $r(x,y)$ (possibly with parameters from $M$) such that $r(x,y)$ that defines $R\cap M^{2}$ in $\mathcal{M}$ It is clear that $r(x,y)$ defines a global linear ordering in $\mathcal{M}$. $\square$
3.7. Corollary. There are models of $\mathrm{ZFC}$ that have no taller conservative elementary extensions to a model of $\mathrm{ZFC}$.
Proof. This follows from Proposition 3.6, and the well-known fact that there are models of ZFC in which the universe is not definably linearly ordered. [^15] Easton proved (in his unpublished dissertation [Ea]) that assuming Con($\mathrm{ZF}$) there is a model $\mathcal{M}$ of $\mathrm{ZFC}$ that carries no $\mathcal{M}$-definable global choice function for the class of pairs in $\mathcal{M}$; and in particular the universe cannot be definably linearly ordered in $\mathcal{M}$. Easton’s theorem was exposited by Felgner [F, p.231]; for a more recent and streamlined account, see Hamkins’ MathOverflow answer [H]. $\square$
3.8. QuestionCan the conclusion of Proposition 3.6 be improved to $\mathcal{M}\models\exists p\left(\mathrm{V}=\mathrm{HOD}(p)\right)$?
3.9. Remark. If $\mathcal{M}\models\mathrm{ZF}$ and $\mathcal{M}\prec_{\Delta_{0},\ \mathrm{cons}}\mathcal{N}$, then each $k\in\omega^{\mathcal{M}}$ is fixed. To see this, suppose to the contrary that $r\in\mathrm{Ext}_{\mathcal{N}}(k)\backslash\mathrm{Ext}_{\mathcal{M}}(k)$. Consider the set $S=\left\{x\in\mathrm{Ext}_{\mathcal{M}}(k):\mathcal{N}\models x<r\right\}.$ It is easy to see that $S$ has no last element. On the other hand, by the assumption $\mathcal{M}\prec_{\mathrm{cons}}\mathcal{N}$, the set $S=$$\{x\in\mathrm{Ext}_{\mathcal{M}}(k):\mathcal{N}\models x<r\}$ is a definable subset of the predecessors of $r$ and thus by the veracity of the separation scheme in $\mathcal{M}$, $S$ is coded in $\mathcal{M}$ by some element, which makes it clear that $S$ must have a last element, contradiction. This shows that in Theorem 4.1, $\mathcal{N}$ fixes each $k\in\omega^{M}.$ A natural question is whether a taller conservative elementary extension of $\mathcal{M}$ can be arranged to fix $\omega^{\mathcal{M}}$. As we shall see in the next section, this question has a negative answer for models of $\mathrm{ZFC}$ (see Corollary 4.2).
4. Remark. Theorem 4.1 should also be contrasted with Gaifman’s theorem mentioned in Remark 2.5 which bars taller elementary self-extensions of models of (a fragment of) $\mathrm{ZF}$. Thus in Gaifman’s theorem, (a) the condition “$\mathcal{N}$ is an elementary self extension of $\mathcal{M}$” cannot be weakened to “$\mathcal{N}$ is an elementary conservative extension of $\mathcal{M}$”. Also note that in the first proof of Theorem 3.1, each $\mathcal{N}_{k}$ is a self-extension of $\mathcal{M}$ whose associated embedding $j$ is a $\Sigma_{k+2}$-elementary embedding whose image is not cofinal in $\mathcal{M}$.
4FAITHFUL EXTENSIONS
This section contains the proof of Theorem B of the abstract (see Corollary 4.2). Recall from Definition 2.4 and Remark 2.5 that the notion of a faithful extension is a generalization of the notion of a conservative extension. The notion of a $\Sigma_{\omega}$-satisfaction class was defined in Definition 2.7(b).
4.1. Theorem. Suppose $\mathcal{M}$ and $\mathcal{N}$ are both models of $\mathrm{ZFC}$ such that $\mathcal{M}\prec_{\Delta_{0}^{\mathcal{P}},~{}\mathrm{faith}}\mathcal{N}$, $\mathcal{N}$ fixes $\omega^{\mathcal{M}},$ and $\mathcal{N}$ is taller than $\mathcal{M}$. Then:
(a) There is some $\gamma\in\mathrm{Ord}^{\mathcal{N}}$ such that $\mathcal{M}\preceq\mathcal{N}_{\gamma}.$ Thus either $\mathcal{N}$ is a topped rank extension of $\mathcal{M}$, or $\mathcal{M}\prec\mathcal{N}_{\gamma}$. [^16] As pointed out in Remark 4.5, both case of the dichotomy can be realized.
(b) In particular, there is a $\Sigma_{\omega}$-satisfaction class $S$ $\mathcal{M}$ such that $S$ can be written as $D\cap M$, where $D$ is $\mathcal{N}$-definable.
Proof [^17] The proof strategy of Theorem 4.1 is a variant of the proofs of the following results: [En-2, Theorem 1.5], [En-3, Theorem 3.3], [EH, Theorem 2.1], [En-6, Theorem 2.1.3], and [En-7, Lemma 2.19]. A variant of the same strategy is used in the proof of Theorem 4.5. The origins of the strategy can be traced to Kaufmann’s refinement [Kay, Lemma 3.2] of a Skolem hull argument due to Keisler and Silver [KS, Theorem 2.1].. Fix some $\lambda\in\mathrm{Ord}^{\mathcal{N}}\backslash\mathrm{Ord}^{\mathcal{M}}$ and some $\mathcal{N}$-ordinal $\beta>\lambda$. Note that:
(1) $\mathcal{N}_{\beta}\prec_{\Delta_{0}^{\mathcal{P}}}\mathcal{N}.$
[…]
$K_{m}:=\{a\in\mathrm{V}_{\beta}:\mathcal{N}\models a\in\mathrm{Def}($$V$,${}_{\beta},\in,\vartriangleleft,\lambda,m)\},$
where $x\in\mathrm{Def}($$V$,${}_{\beta},\in,\vartriangleleft,\lambda,m)$ is shorthand for the formula of set theory that expresses:
$x$ is definable in the structure (((($V$,${}_{\beta},\in,\vartriangleleft,\lambda,m).$
Note that the only allowable parameters used in a definition of $x$ are $\lambda$ and $m$. Thus, intuitively speaking, $K_{m}$ consists of elements $a$ of $N_{\beta}$ such that within $\mathcal{N}$, $s$ is first order definable in $\left(\mathrm{V}_{\beta},\in,\vartriangleleft,\lambda,m\right)$. Notice that since we are not assuming that $\mathcal{M}$ is $\omega$-standard, an element $a\in K_{m}$ need not be definable in the real world in the structure $\left(\mathcal{N}_{\beta},\vartriangleleft,\lambda,m\right).$ Next we move outside of $\mathcal{N}$ and define $K$ as follows:
$K:=\bigcup\limits_{m\in M}K_{m}.$
[…]
(2) $\ \ \mathcal{M}\subsetneq\mathcal{K}\preceq\mathcal{N}_{\beta}$.
By putting (1) and (2) together we conclude:
(3) $\ \ \mathcal{M}\prec_{\Delta_{0}^{\mathcal{P}}}\mathcal{K}\preceq\mathcal{N}_{\beta}\prec_{\Delta_{0}^{\mathcal{P}}}\mathcal{N}$.
Let $O^{\ast}$ be the collection of ‘ordinals’ of $\mathcal{K}$ that are above the ‘ordinals’ of $\mathcal{M}$, i.e.,
$O^{\ast}=\left\{\gamma\in\mathrm{Ord}^{\mathcal{K}}:\forall\alpha\in\mathrm{Ord}^{\mathcal{M}}\ \mathcal{N}\models\alpha\in\gamma\right\}.$
Clearly $O^{\ast}$ is nonempty since $\lambda\in O$ We now consider the following two cases. As we shall see, Case I leads to the conclusion of the theorem, and Case II is impossible.
Case I. $O^{\ast}$ has a least ordinal (under $\in^{\mathcal{K}}).$
Case II. $O^{\ast}$ has no least ordinal.
Suppose Case I holds and let $\gamma_{0}=\min(O^{\ast})$. Clearly $\gamma_{0}$ is a limit ordinal of $\mathcal{N}$.  By the choice of $\gamma_{0},$ $\mathrm{Ord}^{\mathcal{M}}$ is cofinal in $\mathrm{Ord}^{\mathcal{K}_{\gamma_{0}}}.$ Since by (3) and part (d) of Remark 2.6, $\mathrm{V}_{\alpha}^{\mathcal{M}}=\mathrm{V}_{\alpha}^{\mathcal{K}}$ for each $\alpha\in\mathrm{Ord}^{\mathcal{M}}$, this makes it clear that:
(4) $\ \ \mathcal{M}\preceq_{\Delta_{0}^{\mathcal{P}},\ \mathrm{cof}}\mathcal{K}_{\gamma_{0}}.$
Since $\mathcal{K}_{\gamma_{0}}$ is the convex hull of $\mathcal{M}$ in $\mathcal{K}$, by Gaifman Splitting Theorem, (4) shows:
(5) $\ \ \mathcal{M}\preceq\mathcal{K}_{\gamma_{0}}$.
On the other hand since by (3) $\mathcal{K}\preceq\mathcal{N}_{\beta}$, we have:
(6) $\mathcal{K}_{\gamma_{0}}=\mathrm{V}_{\gamma_{0}}^{\mathcal{K}}\preceq\mathrm{V}_{\gamma_{0}}^{\mathcal{N}_{\beta}}=\mathrm{V}_{\gamma_{0}}^{\mathcal{N}}=\mathcal{N}_{\gamma_{0}}.$
By (5) and (6), $\mathcal{M}\preceq\mathcal{N}_{\gamma_{0}}$. Thus the proof of the theorem will be complete once we show that Case II leads to a contradiction. Within $\mathcal{N}$ let $s_{\beta}$ be the elementary diagram of $\left(\mathrm{V}_{\beta},\in\right)$, i.e., $\mathcal{N}$ views $s_{\beta}$ is the Tarskian satisfaction class for $\left(\mathrm{V}_{\beta},\in\right)$, and let $\Phi:=\bigcup\limits_{m\in M}\Phi_{m}$, where:
$\Phi_{m}:=\{x\in M:\mathcal{N}\models$ $x$ is (the code of) a formula $\varphi(c,c_{m})$ such that $\varphi(c_{\lambda},c_{m})\in s_{\beta}\}\mathrm{.}$
So intuitively speaking, $\Phi$ is the result of replacing $c_{\lambda}$ by $c$ (where $c$ is a fresh constant) in the sentences in the elementary diagram of $\mathcal{N}_{\beta}$ (as computed in $\mathcal{N}$) whose constants are in $\{c_{\lambda}\}\cup\{c_{m}:m\in M\}.$ Since $\mathcal{N}$ need not be $\omega$-standard, the elements of $\Phi$ might be nonstandard formulae.
By the assumption that $\mathcal{N}$ is a faithful extension of $\mathcal{M}$, $\Phi$ is $\mathcal{M}$-amenable. Next let:
$\Gamma:=\left\{t(c,c_{m})\in M:t(c,c_{m})\in\Phi\ \mathrm{and}\ \forall\theta\in\mathrm{Ord\ }\left(t(c,c_{m})>c_{\theta}\right)\in\Phi\right\},$
where $t$ is a definable term in the language $\mathcal{L}_{\mathrm{set}}\cup\{c\}\cup\{c_{m}:m\in M\}.$ Officially speaking, $\Gamma$ consists of syntactic objects $\varphi(c,c_{m},x)$ in $\mathcal{M}$ that satisfy the following three conditions in $(\mathcal{M},\Phi)$:
$(i)\qquad\left[\exists!x\varphi(c,c_{m},x)\right]\in\Phi.$
$(ii)\qquad\left[\forall x\left(\varphi(c,c_{m},x)\rightarrow x\in\mathrm{Ord}\right)\right]\in\Phi.$
$(iii)\qquad\forall\theta\in\mathrm{Ord}\ \left[\forall x\left(\varphi(c,c_{m},x)\rightarrow c_{\theta}\in x\right)\right]\in\Phi.$
Note that $\Gamma$ is definable in $\left(\mathcal{M},\Phi\right)$. Since we are considering Case II, $\left(\mathcal{M},\Phi\right)\models\psi$, where:
$\psi:=\forall t\left(t\in\Gamma\longrightarrow\left[\exists t^{\prime}\in\Gamma\wedge\left((t^{\prime}\in t)\in\Phi\right)\right]\right).$
The veracity of the dependent choice scheme in $\mathcal{M}$ (see part (e) of Definition 2.1), together with the facts that $\Phi$ is $\mathcal{M}$-amenable and $\Gamma$ is definable in $\left(\mathcal{M},\Phi\right)$ make it clear that there is a sequence $s=\left\langle t_{n}:n\in\omega^{\mathcal{M}}\right\rangle$ in $\mathcal{M}$ such that:
$\left(\mathcal{M},\Phi\right)\models\forall n\in\omega\ \left[t_{n}\in\Gamma\wedge\left(\left(t_{n+1}\in t_{n}\right)\in\Phi\right)\right].$
Since $s$ is a countable object in $\mathcal{M}$, and $\mathcal{N}$ fixes $\omega^{M}$ by assumption, $s$ is fixed in the passage from $\mathcal{M}$ to $\mathcal{N}$ by Remark 2.6(c). On the other hand, since $\mathcal{N}$ has a satisfaction predicate $s_{\beta}$ for $\mathcal{N}_{\beta}$, this leads to a contradiction because we have:
$\mathcal{N}\models\left\langle t_{n}^{(\mathrm{V}_{\beta},\in)}(c_{\lambda}):n\in\omega\right\rangle$ is an infinite decreasing sequence of ordinals.
In the above $t_{n}(c_{\lambda})$ is the term obtained by replacing $c$ with $c_{\lambda}$ in $t_{n},$ and $t_{n}^{(\mathrm{V}_{\beta},\in)}(c_{\lambda})$ is the interpretation of $t_{n}(c_{\lambda})$ in $(\mathrm{V}_{\beta},\in)$. $\square$
4.2. Corollary. If $\mathcal{M}\models\mathrm{ZFC}$, $\mathcal{M}\prec_{\Delta_{0}^{\mathcal{P}},\ \mathrm{cons}}\mathcal{N}$, and $\mathcal{M}$ fixes $\omega^{\mathcal{M}},$ then $\mathcal{M}$ is cofinal in $\mathcal{N}$. Thus a conservative elementary extension of $\mathcal{M}$ that fixes $\omega^{\mathcal{M}}$ is a cofinal extension.
Proof. If $\mathcal{M}\models\mathrm{ZFC}$, $\mathcal{M}\prec_{\mathrm{cons}}\mathcal{N}$, and $\mathcal{M}$ is not cofinal in $\mathcal{N}$, then $\mathcal{N}$ is taller than $\mathcal{M}$. But then by part (b) of Theorem 4.1, there is an $\mathcal{M}$-definable $\Sigma_{\omega}$-satisfaction class over $\mathcal{M}$, which contradicts Tarski’s Undefinability of Truth Theorem. $\square$
In Theorem 4.4 below we show that the conclusion of Theorem 4.1 holds if $\mathrm{ZFC}$ is weakened to $\mathrm{ZF,}$ but the assumption that $\mathcal{N}$ fixes $\omega^{\mathcal{M}}$ is strengthened to $\mathcal{M}\subsetneq_{\mathrm{end}}\mathcal{N}$. We first establish a lemma that will come handy in the proof of Theorem 4.4.
4.3. Lemma Suppose $\mathcal{M}$ and $\mathcal{N}$ are models of $\mathrm{ZF}$, and $\mathcal{M}\subsetneq_{\mathrm{end,faithful}}\mathcal{N}$. Then $\mathcal{M}\subsetneq_{\mathrm{rank}}\mathcal{N}$.
Proof. Suppose $\mathcal{M}$ and $\mathcal{N}$ are as in the assumptions of the lemma. Since $\mathcal{N}$ end extends $\mathcal{M}$, in order to show that $\mathcal{N}$ rank extends $\mathcal{M}$ it is sufficient that if $\alpha\in\mathrm{Ord}^{\mathcal{M}}$ and $m\in M$ with $\mathcal{M}\models m=\mathrm{V}_{\alpha}$, then $\mathcal{N}\models m=\mathrm{V}_{\alpha}.$ For this purpose, it is sufficient to verify that $\mathcal{N}$ is a powerset-preserving end extension of $\mathcal{M}$, since the formula expressing $x=\mathrm{V}_{\alpha}$ is $\Sigma_{1}^{\mathcal{P}}$ and is therefore absolute for powerset-preserving end extensions (as noted in Remark 2.6(e)). Towards this goal, let $a\in M$ and suppose $b\in N$ such that:
(1) $\ \ \mathcal{N}\models b\subseteq a.$
Let $X_{b}=\mathrm{Ext}_{\mathcal{N}}(b)\cap\mathrm{Ext}_{\mathcal{M}}(a)=\mathrm{Ext}_{\mathcal{N}}(b).$ Thanks to the assumption that $\mathcal{N}$ is a faithful extension of $\mathcal{M}$, the expansion $(\mathcal{M},X_{b})$ satisfies Separation in the extended language and therefore there is an element $b^{\prime}\in M$ that codes $X_{b}.$ Thus
(2) $\ \ \mathrm{Ext}_{\mathcal{M}}(b^{\prime})=\mathrm{Ext}_{\mathcal{N}}(b)\cap\mathrm{Ext}_{\mathcal{M}}(a),$ and
Since $\mathcal{N}$ end extends $\mathcal{M}$, we have:
[…]
So by (1) and (2),
[…]
By (1), $\mathrm{Ext}_{\mathcal{N}}(b)\subseteq\mathrm{Ext}_{\mathcal{N}}(a),$ so combined with (4) this yields:
(5) $\ \ \mathrm{Ext}_{\mathcal{N}}(b^{\prime})=\mathrm{Ext}_{\mathcal{N}}(b)$,
which by Extensionality makes it clear that $b^{\prime}=b.$square
  • The proof of Theorem 4.4 below is a variant of the proof of Theorem 4.1, and will strike the reader who has worked through the proof of Theorem 4.1 as repetitious, but the proof is a bit more involved here since we do not have access to the axiom of choice in $\mathcal{M}$ or in $\mathcal{N}$. Note that in Theorem 4.4 $\mathcal{N}$ is assumed to end extend $\mathcal{M}$, whereas in Theorem 4.1 $\mathcal{N}$ is only assumed to fix $\omega^{\mathcal{M}}.$ Of course Theorem 4.4 follows from Theorem 4.1 if $\mathcal{M}$ and $\mathcal{N}$ are models of $\mathrm{ZFC}$.
4.4. Theorem. Suppose $\mathcal{M}$ and $\mathcal{N}$ are models of $\mathrm{ZF}$, and $\mathcal{M}\subsetneq_{\mathrm{end,faithful}}\mathcal{N}$. Then:
(a) There is some $\gamma\in\mathrm{Ord}^{\mathcal{N}}\backslash\mathrm{Ord}^{\mathcal{M}}$ such that $\mathcal{M}\preceq\mathcal{N}_{\gamma}.$ Thus either $\mathcal{N}$ is a topped rank extension of $\mathcal{M}$, or $\mathcal{M}\prec\mathcal{N}_{\gamma}$.
(b) In particular, there is a $\Sigma_{\omega}$-satisfaction class $S$ $\mathcal{M}$ such that $S$ can be written as $D\cap M$, where $D$ is $\mathcal{N}$-definable.
Proof. Assume $\mathcal{M}$ and $\mathcal{N}$ are as in the assumptions of the theorem. By Lemma 4.3 we can assume that $\mathcal{N}$ rank extends $\mathcal{M}$. Fix some $\lambda\in\mathrm{Ord}^{\mathcal{N}}\backslash\mathrm{Ord}^{\mathcal{M}}$ and some ordinal $\beta>\lambda$. For each $m\in M$, we can define the following set within $\mathcal{N}$:
$O_{m}:=\left\{\gamma\in\beta:\mathcal{N}\models\gamma\in\mathrm{Def}(\mathrm{V}_{\beta},\in,\lambda,m)\right\}.$
where $x\in\mathrm{Def}($$V$,${}_{\beta},\in,\lambda,m)$ is shorthand for the formula of set theory that expresses:
$x$ is definable in (((($V$,${}_{\beta},\in)$ using at most the parameters $\lambda$ and $m.$
Clearly $\lambda\in O_{m}$ and $O_{m}$ is coded in $\mathcal{N}$. Next let:
$O:=\bigcup\limits_{m\in M}O_{m}.$
Let $O^{\ast}=O\backslash\mathrm{Ord}^{\mathcal{M}}$. Clearly $O^{\ast}$ is nonempty since $\lambda\in O^{\ast}.$ We now consider two cases:
(I) $\ \ O^{\ast}$ has a least element (under $\in^{\mathcal{N}}).$
(II) $\ \ O^{\ast}$ has no least element.
Case I. Let $\gamma_{0}$ be the least element of $O^{\ast}.$ It is easy to see that $\gamma_{0}$ is a limit ordinal in $\mathcal{N}$. We claim that
$\mathcal{M}\preceq\mathcal{N}_{\gamma_{0}}.$
To verify the claim, by the Tarski criterion of elementarity, it suffices to show that if $\mathcal{N}(\gamma_{0})\models\exists y\varphi(y,m)$ for some $m\in M$, then there is some $m_{0}\in M$ such that $\mathcal{N}_{\gamma_{0}}\models\varphi(m_{0},m).$ Let $\gamma_{m}$ be the first ordinal below $\gamma_{0}$ such that:
$\mathcal{N}_{\gamma_{0}}\models\exists y\in\mathrm{V}_{\gamma_{m}}\ \varphi(y,m).$
So $\gamma_{m}$ is definable in $\mathcal{N}_{\beta}$ with parameters $\gamma_{0}$ and $m$, and since $\gamma_{0}\in O$ by assumption, this shows that $\gamma_{m}$ is definable in $\mathcal{N}(\beta)$ with parameter $m.$ So $\gamma_{m}\in O.$ By the fact that $\gamma_{0}$ was chosen to be the least element of $O^{\ast}$, this shows that $\gamma_{m}\in\mathrm{Ord}^{\mathcal{M}},$ which makes it clear that there is some $m_{0}\in M$ such that:
$\mathcal{N}_{\gamma_{0}}\models\varphi(m_{0},m).$
This concludes the verification of $\mathcal{M}\preceq\mathcal{N}_{\gamma_{0}}$. To finish the proof of Theorem 4.4 it suffices to show that Case II is impossible. Suppose to the contrary that $\mathrm{Ord}^{\mathcal{K}}\backslash\mathrm{Ord}^{\mathcal{M}}$ has no least element. Within $\mathcal{N}$ let $s_{\beta}$ be the Tarskian satisfaction class for $\left(\mathrm{V}_{\beta},\in\right)$. For each $m\in M$ let
$\Phi_{m}:=\{\varphi(c,c_{m})\in M:\mathcal{N}\models$ $\varphi(c_{\lambda},c_{m})\in s_{\beta}\}\mathrm{,}$
and let
$\Phi:=\bigcup\limits_{m\in M}\Phi_{m}$.
Note that $\Phi\subseteq M$ and $\Phi$ is $\mathcal{M}$-amenable since $\mathcal{N}$ is a faithful rank extension of $\mathcal{M}$. So intuitively speaking, $\Phi$ is the subset of the elementary diagram of $\mathcal{N}_{\beta}$ (as computed in $\mathcal{N}$) that consists of sentences whose constants are in $\{c\}\cup\{c_{m}:m\in M\}.$ Since $\mathcal{N}$ need not be $\omega$-standard, the elements of $\Phi$ need not be standard formulae. Also note that the constant $c$ is interpreted as $\lambda$ in the right-hand side of the above definition of $\Phi_{m}.$ Now let:
$\Gamma:=\left\{t(c,c_{m})\in M:t(c,c_{m})\in\Phi\ \mathrm{and}\ \forall\theta\in\mathrm{Ord\ }\left(t(c,c_{m})>c_{\theta}\right)\in\Phi\right\},$
where $t$ is a definable function in the language $\mathcal{L}_{\mathrm{set}}\cup\{c\}\cup\{c_{m}:m\in M\}.$ So, officially speaking, $\Gamma$ consists of syntactic objects $\varphi(c,c_{m},x)$ in $\mathcal{M}$ that satisfy the following three conditions in $(\mathcal{M},\Phi)$:
(1) $\ \ \left[\exists!x\varphi(c,c_{m},x)\right]\in\Phi.$
(2) $\ \ \left[\forall x\left(\varphi(c,c_{m},x)\rightarrow x\in\mathrm{Ord}\right)\right]\in\Phi.$
(3) $\ \ \forall\theta\in\mathrm{Ord}\ \left[\forall x\left(\varphi(c,c_{m},x)\rightarrow c_{\theta}\in x\right)\right]\in\Phi.$
Clearly $\Gamma$ is definable in $\left(\mathcal{M},\Phi\right)$. Since we are in Case II, $\left(\mathcal{M},\Phi\right)\models\psi$, where $\psi$ is the sentence that expresses:
$\Gamma\neq\varnothing\wedge\forall t\left(t\in\Gamma\longrightarrow\left[\exists t^{\prime}\in\Gamma\wedge\left((t^{\prime}\in t)\in\Phi\right)\right]\right).$
In contrast to the proof of Theorem 4.1 we cannot at this point conclude that there is a countable descending chain $\left\langle t_{n}:n\in\omega\right\rangle$ in $\mathcal{M}$ since $\mathrm{DC}$ need not hold in $\mathcal{M}$. Instead, we will use the following argument that takes advantage of the Reflection Theorem. Choose $k\in\omega$ such that $\psi$ is a $\Sigma_{k}(\Phi)$-statement, and use the Reflection Theorem in $\left(\mathcal{M},\Phi\right)$ to pick $\mu\in\mathrm{Ord}^{\mathcal{M}}$ such that:
$\left(\mathcal{M}_{\mu},\Phi\cap M_{\mu}\right)\prec_{\Sigma_{k}(\mathrm{X})}\left(\mathcal{M},\Phi\right).$ [^18] Here $\mathrm{X}$ is interpreted by the left-hand side structure as $\Phi\cap M_{\mu}$ and by the right-hand side structure as $\Phi$.
Then $\psi$ holds in $\left(\mathcal{M}_{\mu},\Phi\cap M_{\mu}\right)$. Within $\mathcal{M}$, let
$w=\{v\in\mathrm{V}_{\mu}:v\in\Gamma\}.$
Thus $\left(\mathcal{M},\Phi\cap M\right)$ satisfies $w\neq\varnothing\wedge\forall t\in w\exists t^{\prime}\in w\wedge\left((t^{\prime}\in t)\in\Phi\right).$ $\mathcal{N}$ can evaluate each term $t(c)$ in $w$ as an ordinal $\delta<\beta$, where $\mathcal{N}_{\beta}\models t(c_{\lambda})=\delta$ (here $t(c_{\lambda})$ is the term obtained by replacing all occurrences of the constant $c$ with $c_{\lambda}).$ So we can consider $s\in N$, where:
$\mathcal{N}\models s=\left\{\delta:\exists t(c)\in w:(\mathrm{V}_{\beta},\in)\models t(c_{\lambda}):t(c)\in w\right\}.$
Within $\mathcal{N}$, $s$ is a nonempty of ordinals that has no least element, which of course is a contradiction. This shows that Case II is impossible, thus concluding the proof. $\square$
4.5. Remark. Arguing in $\mathrm{ZFC}$, suppose $\kappa$ is the first strongly inaccessible cardinal, $\lambda$ is the second strongly inaccessible cardinal, $\mathcal{M}=(\mathrm{V}_{\kappa},\in)$ and $\mathcal{N}=(\mathrm{V}_{\lambda},\in)$. Then $\mathcal{N}$ is a proper faithful rank extension of $\mathcal{M}$ in which the first clause of the dichotomy of the conclusion of part (a) of Theorem 4.1 (and 4.4) holds , but not the second one. On the other hand, if $\kappa$ is a weakly compact cardinal, then it is well-known that $(\mathrm{V}_{\kappa},\in)$ has a faithful topless elementary end extension $\mathcal{N}$; in this scenario the second clause of the dichotomy of the conclusion of part (a) of Theorem 4.1 (and 4.4) holds, but not the first one.
4.6. Remark. If the assumptions of Theorem 4.4 are strengthened by adding the assumption that $\mathcal{M}\prec\mathcal{N}$, then as in the proof of Theorem 3.3 of [En-3] we can use ‘Kaufmann’s trick’ to conclude that there is some $\mathcal{K}$ such that $\mathcal{M}\prec\mathcal{K}\preceq\mathcal{N}$ and $\mathcal{K}$ is a topped rank extension of $\mathcal{M}$. This yields a strengthening of Theorem 3.3 of [En-3] by eliminating the assumption that the axiom of choice is true in $\mathcal{M}$ and $\mathcal{N}$.
4.7. Remark. An inspection of the proof of Theorem 4.1 makes it clear that the assumption that $\mathcal{M}\models\mathrm{ZFC}$ can be reduced to $\mathcal{M}\models\mathrm{ZF+DC}$, and the assumption that $\mathcal{N}\models\mathrm{ZFC}$ can be reduced to $\mathcal{N}\models\mathrm{KPR}$ + “every set can be well-ordered”. Here $\mathrm{KPR}$ is as in part (d) of Definition 2.1. Similarly, an inspection of the proof of Theorem 4.4 shows that the assumption that $\mathcal{N}\models\mathrm{ZF}$ can be reduced to $\mathcal{N}\models\mathrm{KPR}$.
4.8. Question. Can Theorem 4.1 and Corollary 4.2 be strengthened by only assuming that $\mathcal{M}$ and $\mathcal{N}$ are models of $\mathrm{ZF}$?
5DEAD-END MODELS
In this section we will establish Theorem C of the abstract, the proof uses many ingredients, including the following one that refines a result obtained independently by Kaufmann and the author who demonstrated Theorem 5.1 for models of $\mathrm{ZF}$ in which the Axiom of Choice holds (see “Added in Proof” of [Kau] and Remark 1.6 of [En-2]).
5.1. Theorem. No model of $\mathrm{ZF}$ has a conservative proper end extension satisfying $\mathrm{ZF}$.
Proof. This follows from putting part (b) of Theorem 4.4 together with Tarski’s Undefinability of Truth Theorem (as in Theorem 2.8). $\square$.□ .
The rest of the section is devoted to presenting results that in conjunction with Theorem 5.1 will allow us to establish Theorem C of the abstract (as Theorem 5.18).
5.2. Definition. Suppose $\mathcal{M}$ is an $\mathcal{L}_{\mathrm{set}}$-structure.
(a) $X\subseteq M$ is piecewise-coded in $\mathcal{M}$ if $\forall a\in M\ \exists b\in M\ \mathrm{Ext}_{\mathcal{M}}(b)=X\cap\mathrm{Ext}_{\mathcal{M}}(a).$
(b) $\mathcal{M}$ is rather classless iff every piecewise-coded subset of $M$ is $\mathcal{M}$-definable.
(c) $\mathcal{M}$ is $\aleph_{1}$-like if $\left|M\right|=\aleph_{1}$ but $\left|\mathrm{Ext}_{\mathcal{M}}(a)\right|\leq\aleph_{0}$ for each $a\in M.$
5.3. Theorem. (Keisler-Kunen [Ke-1], Shelah [Sh]). Every countable model of $\mathrm{ZF}$ has an elementary end extension to an $\aleph_{1}$-like rather classless model.
5.4. Proposition. No rather classless model of $\mathrm{ZF}$ has a proper rank extension to a model of $\mathrm{ZF}$.
Proof. This follows from putting Theorem 4.4 together with the observation that a rank extension of a model of $\mathrm{ZF}$ is a conservative extension, and therefore a faithful extension. $\square$
5.5. Remark. As pointed out in Remark 1.6 of [En-3] it is possible for a rather classless model to have a proper end extension satisfying $\mathrm{ZF}$, since $\aleph_{1}$-like rather classless models exist by Theorem 5.3, and one can use the Boolean-valued approach to forcing to construct set generic extensions of such models.
5.6. Definition. A ranked tree $\tau$ is a two-sorted structure $\tau=(\mathbb{T},\ \leq_{\mathbb{T}},\ \mathbb{L},\ \leq_{\mathbb{L}},\ \rho)$ satisfying the following three properties:
(1) $(\mathbb{T},\ \leq_{\mathbb{T}})$ is a tree, i.e., a partial order such that any two elements below a given element are comparable.
(2) $(\mathbb{L},\ \leq_{\mathbb{L}})$ is a linear order with no last element.
(3) $\rho$ is an order preserving map from $(\mathbb{T},\ \leq_{\mathbb{T}})$ onto $(\mathbb{L},\ \leq_{\mathbb{L}})$ with the property that for each $t\in\mathbb{T},$ $\rho$ maps the set of predecessors of $t$ onto the initial segment of $(\mathbb{L},\ \leq_{\mathbb{L}})$ consisting of elements of $L$ less than $\rho(t).$
5.7. Definition. Suppose $\tau=(\mathbb{T},\ \leq_{\mathbb{T}},\ \mathbb{L},\ \leq_{\mathbb{L}},\ \rho)$ is a ranked tree. A linearly ordered subset $B$ of $\mathbb{T}$ is said to be a branch of $\tau$ if the image of $B$ under $\rho$ is $\mathbb{L}$. The cofinality of $\tau$ is the cofinality of $(\mathbb{L},\ \leq_{\mathbb{L}}).$
[…]
5.9. Theorem. Suppose $\mathcal{M}$ is a countable structure in a countable language.
(a) (Keisler-Kunen [Ke-1], essentially). It is a theorem of $\mathrm{ZFC}+\Diamond_{\omega_{1}}$ that $\mathcal{M}$ can be elementarily extended to a rather branchless model.
(b) (Shelah [Sh]). It is a theorem of $\mathrm{ZFC}$ that $\mathcal{M}$ can be elementarily extended to a rather branchless model.
5.10. Definition. Suppose $(\mathbb{P},\leq_{\mathbb{P}})$ is a poset (partially ordered set).
(a) $(\mathbb{P},\leq_{\mathbb{P}})$ is directed if any given pair of elements of $\mathbb{P}$ has an upper bound.
(b) A subset $F$ of $\mathbb{P}$ is a filter over $(\mathbb{P},\leq_{\mathbb{P}})$ if the sub-poset $(F,\leq_{P})$ is directed.
(c) A filter over $(\mathbb{P},\leq_{\mathbb{P}})$ is maximal if it cannot be properly extended to a filter over $(\mathbb{P},\leq_{\mathbb{P}})$. [^19] Thus a maximal filter can be described as a maximally compatible subset of $(\mathbb{P},\leq_{\mathbb{P}}).$
(d) A subset $C$ of $\mathbb{P}$ is cofinal in $(\mathbb{P},\leq_{\mathbb{P}})$ if $\forall x\in\mathbb{P}\ \exists y\in C\ y\leq_{\mathbb{P}}x.$
5.11. Definition. Suppose $s$ is an infinite set.
(a) $[s]^{<\omega}$ is the directed poset of finite subsets of $s,$ ordered by containment. Note that $[s]^{<\omega}$ is a directed set with no maximum element.
(b) $\mathrm{Fin}(s,2)$ is the poset of finite functions from $s$ into $\{0,1\}$, ordered by containment (where a function is viewed as a set of ordered pairs).
5.12. Example. Given an infinite set $s$, and $a\subseteq s,$ let $\chi_{a}:s\rightarrow 2$ be the characteristic function of $a,$ i.e., $\chi_{a}(x)=1$ iff $x\in a.$ Let $F_{a}=$ finite approximations to $\chi_{a}$. $F_{a}$ is a maximal filter of $\mathrm{Fin}(s,2).$
5.13. Definition. A structure $\mathcal{M}$ is a Rubin model if it has the following two properties:
(a) Every $\mathcal{M}$-definable directed poset $(\mathbb{D},\leq_{\mathbb{D}})$ with no maximum element has a cofinal chain of length $\omega_{1}.$
(b) Given any $\mathcal{M}$-definable poset $\mathbb{P}$, and any maximal filter $F$ of $\mathbb{P}$, if $F$ has a cofinal chain of length $\omega_{1}$, then $F$ is coded in $\mathcal{M}$.
5.14. Remark. If $\tau=(\mathbb{T},\ \leq_{\mathbb{T}},\ \mathbb{L},\ \leq_{\mathbb{L}},\ \rho)$ is a ranked tree, then each branch of $\tau$ is a maximal filter over $(\mathbb{T},\ \leq_{\mathbb{T}}).$ This makes it clear that every Rubin model is rather branchless. Also, a rather branchless model of $\mathrm{ZF}$ is rather classless. To see this consider the ranked tree $\tau_{\mathrm{class}}$ defined within a model $\mathcal{M}$ of $\mathrm{ZF}$ as follows: The nodes of $\tau_{\mathrm{class}}$ are ordered pairs $(s,\alpha)$, where $s\subseteq\mathrm{V}_{\alpha}$, the rank of $(s,\alpha)$ is $\alpha$ and $(s,\alpha)<(t,\beta)$ if $\alpha\in\beta$ and $s=t\cap\mathrm{V}_{\alpha}$. It is easy to see that $\mathcal{M}$ is rather classless iff every branch of $\tau_{\mathrm{class}}^{\mathcal{M}}$ is $\mathcal{M}$-definable. Hence we have the following chain of implications:
Rubin $\Rightarrow$ rather branchless $\Rightarrow$ rather classless.
5.15. Theorem. (Rubin [^20] The attribution to Rubin is informed by fact that the credit for [RS, Lemma 2.3] is explicitly given to Rubin (at the end of the introduction of the paper). [RS, Corollary 2.4]). It is a theorem of $\mathrm{ZFC}+\Diamond_{\omega_{1}}$ that if $\mathcal{M}$ is a countable structure in a countable language, then $\mathcal{M}$ has an elementary extension of cardinality $\aleph_{1}$ that is a Rubin model. [^21] The proof of [RS, Lemma 2.3] has a small gap (it is assumed that $T_{0}\vdash\exists\overline{x}\psi(\overline{x})$ instead of assuming the consistency of $T_{0}+\exists\overline{x}\psi(\overline{x})$). The improved proof was presented in [En-1, Theorem 2.1.3] and is reproduced here in the Appendix.
5.16. Definition. Suppose $\mathcal{M}$ is a model of $\mathrm{ZF}$. $\mathcal{M}$ is weakly Rubin if (a) and (b) below hold:
(a) $\mathcal{M}$ is rather classless (this is the asymptotic case of $(ii)$ below, if one could use $a=\mathrm{V}$).
(b) For every element $a$ of $\mathcal{M}$ that is infinite in the sense of $\mathcal{M}$ the following two statements hold:
$(i)$ $\left([a]^{<\omega}\right)^{\mathcal{M}}$ has a cofinal chain of length $\omega_{1}.$
$(ii)$ If $F$ is a maximal filter of $\mathrm{Fin}^{\mathcal{M}}(a,2)$ and $F$ has a cofinal chain of length $\omega_{1}$, then $F$ is coded in $\mathcal{M}$.
As demonstrated in the Appendix, Schmerl’s strategy of $\Diamond_{\omega_{1}}$-elimination in [Schm-2] (which is based on an absoluteness argument first presented by Shelah [Sh]) can be employed to build weakly Rubin models within ZFC. Thus we have:
5.17. Theorem (Rubin-Shelah-Schmerl) It is a theorem of $\mathrm{ZFC}$ that every countable model of $\mathrm{ZF}$ has an elementary extension to a weakly Rubin model of cardinality $\aleph_{1}$.
We are finally ready to establish the main result of this section on the existence of ‘dead-end’ models. Note that since every generic extension is an end extension (even for ill-founded models), the model $\mathcal{M}$ in Theorem 5.18 has no proper generic extension.
5.18. Theorem. Every countable model $\mathcal{M}_{0}\models\mathrm{ZF}$ has an elementary extension $\mathcal{M}$ of power $\aleph_{1}$ that has no proper end extension to a model $\mathcal{N}\models\mathrm{ZF}$. Thus every consistent extension of $\mathrm{ZF}$ has a model of power $\aleph_{1}$ that has no proper end extension to a model of $\mathrm{ZF}$.
Proof [^22] A more direct proof can be given using the machinery of recursive set functions, but in the interest of conceptual clarity, we opted for the proof presented here.. We begin with a basic fact that will be called upon in the proof.
[…]
$\left(\left[a\right]^{<\omega}\right)^{\mathcal{M}}=\left(\left[a\right]^{<\omega}\right)^{\mathcal{N}}.$
Proof. The assumptions on $\mathcal{M}$ and $\mathcal{N}$ readily imply that $\omega^{\mathcal{M}}=\omega^{\mathcal{N}}$ and $\mathrm{Ord}^{\mathcal{M}}\subseteq_{\mathrm{end}}\mathrm{Ord}^{\mathcal{N}}$. Recall that within $\mathrm{ZF}$, given any set parameter $a$ one can define the inner model $\mathrm{L}(a)$ as
$\bigcup\limits_{\alpha\in\mathrm{Ord}}\mathrm{L}_{\alpha}(a)$,
where $\mathrm{L}_{0}(a)$ is the transitive closure of $\{a\},$ $\mathrm{L}_{\alpha+1}(a)$ is the set of all subsets of $\mathrm{L}_{\alpha}(a)$ that are parametrically definable in $(\mathrm{L}_{\alpha+1}(a),\in),$ and for limit $\alpha,$
$\mathrm{L}_{\alpha}(a)=\bigcup\limits_{\beta\in\mathrm{\alpha}}\mathrm{L}_{\beta}(a).$
It is well-known that for each axiom $\varphi$ of $\mathrm{ZF}$, the statement $\forall a\ \varphi^{\mathrm{L}(a)}$ is provable in $\mathrm{ZF}$, and in particular, $\mathrm{ZF}$ can prove the following statement:
$\forall a\ \exists\alpha\in\mathrm{Ord\ }\left(\exists x\mathrm{\ }(x=\left[a\right]^{<\omega})\right)^{\mathrm{L}_{\alpha}(a)}$ (indeed $\alpha=2$ works).)).) .
On the other hand, the formula expressing $x=\mathrm{L}_{\alpha}(a)$ is provably $\Delta_{1}$ (in parameters $\alpha$ and $a$), so for a fixed $a\in M$ there is for some $\alpha\in\mathrm{Ord}^{\mathcal{M}}$ such that $\left(\left[a\right]^{<\omega}\right)^{\mathrm{L}_{\alpha}^{\mathcal{M}}(a)}=\left(\left[a\right]^{<\omega}\right)^{\mathrm{L}_{\alpha}^{\mathcal{N}}(a)}.$ Thus $\left(\left[a\right]^{<\omega}\right)^{\mathcal{M}}=\left(\left[a\right]^{<\omega}\right)^{\mathcal{N}}$ since
$\left(\left[a\right]^{<\omega}\right)^{\mathcal{M}}=\left(\left[a\right]^{<\omega}\right)^{\mathrm{L}_{\alpha}^{\mathcal{M}}(a)}$ and $\left(\left[a\right]^{<\omega}\right)^{\mathcal{N}}=\left(\left[a\right]^{<\omega}\right)^{\mathrm{L}_{\alpha}^{\mathcal{N}}(a)}.$
[…]
Given a countable model $\mathcal{M}_{0}$ of $\mathrm{ZFC}$, by Theorem 5.17 there is a weakly Rubin model $\mathcal{M}$ that elementary extends $\mathcal{M}_{0}$. By Theorem 5.1, to prove Theorem 5.18 it is sufficient to verify that every end extension $\mathcal{N}$ of $\mathcal{M}$ that satisfies $\mathrm{ZF}$ is a conservative extension. Towards this goal, suppose $\mathcal{M}\subsetneq_{\mathrm{end}}\mathcal{N}\models\mathrm{ZF}$. In light of Remark 2.6(e), it suffices to show that $\mathcal{M}\subsetneq_{\mathrm{end}}^{\mathcal{P}}\mathcal{N}$.
To show $\mathcal{M}\subsetneq_{\mathrm{end}}^{\mathcal{P}}\mathcal{N}$, suppose for $a\in M$ and $s\in N$, $\mathcal{N}\models s\subseteq a$ for some $a\in N.$ We will show that $s\in M.$ By Fact ()($\nabla$)( ∇ ) we may assume that $a$ is infinite as viewed from $\mathcal{M}$. Note that this implies that $\mathcal{M}$ views $[a]^{<\omega}$ as a directed set with no maximum element. Fact ()($\nabla$)( ∇ ), assures us that:
()($\ast$)( ∗ ) $\ \forall m\in M$ $\left([m]^{<\omega}\right)^{\mathcal{M}}=\left([m]^{<\omega}\right)^{\mathcal{N}},$ and
Since $\mathrm{Fin}(m,2)\subseteq[m\times\{0,1\}]^{<\omega}$, ()($\ast$)( ∗ ) implies:
[…]
Let $\chi_{s}\in N$ be the characteristic function of $s$ in $\mathcal{N}$, i.e., as viewed from $\mathcal{N}$, $\chi_{s}:a\rightarrow\{0,1\}$ and $\forall x\in a(x\in s\leftrightarrow\chi_{s}(x)=1)$. Let $F_{s}$
$F_{s}:=$ $\mathrm{Ext}_{\mathcal{N}}\left([\chi_{s}]^{<\omega}\right)^{\mathcal{N}}.$
[…]
$q_{\alpha}:=\mathrm{Ext}_{\mathcal{N}}(\chi_{s}\upharpoonright p_{\alpha})=\mathrm{Ext}_{\mathcal{M}}(\chi_{s}\upharpoonright p_{\alpha}).$
Therefore, by the assumption that $\mathcal{M}$ is weakly Rubin, $F_{s}$ is $\mathcal{M}$-definable and thus coded in $\mathcal{M}$ by some $m\in M$. This makes it clear that $s=\cup m$, this shows that $s\in$ $N$, which concludes the verification that $\mathcal{M}\subsetneq_{\mathrm{end}}^{\mathcal{P}}\mathcal{N}.$ As explained earlier, this is sufficient to establish Theorem 5.18. $\square$
An examination of the proof of Theorem 5.1, together with Remark 4.7, makes it clear that Theorem 5.1 can be refined as follows:
5.19. TheoremNo model of $\mathrm{KPR+\Pi}_{\infty}^{1}$-$\mathrm{DC}$ has a conservative proper end extension to a model of KPR.
The proof of Theorem 5.18, together with Theorem 5.19 allows us to refine Theorem 5.18 as follows:
5.20. TheoremEvery countable model $\mathcal{M}_{0}\models\mathrm{KPR+\Pi}_{\infty}^{1}$-$\mathrm{DC}$ has an elementary extension $\mathcal{M}$ of power $\aleph_{1}$ that has no proper end extension to a model $\mathcal{N}\models\mathrm{KPR.}$ Thus every consistent extension of $\mathrm{KPR+\Pi}_{\infty}^{1}$-$\mathrm{DC}$ has a model of power $\aleph_{1}$ that has no proper end extension to model of $\mathrm{KPR.}$
5.21. QuestionIs there an $\omega$-standard model of $\mathrm{ZF(C)}$ that has no proper end extension to a model of $\mathrm{ZF(C)}$?
5.22. Remark. Weakly Rubin models are never $\omega$-standard (even though $\omega$-standard rather classless models exist in abundance). However, a slight variation of the existence proof of Rubin models shows that a Rubin model $\mathcal{M}$ can be arranged to have an $\aleph_{1}$-like $\omega^{\mathcal{M}}.$
6. ADDENDUM AND CORRIGENDUM TO [En-3]
Theorem 2.2 of [En-3] incorrectly stated (in the notation of the same paper, where Gothic letters are used for denoting structures) that if $\mathfrak{A}\models\mathrm{ZFC}$, $\mathfrak{A}\prec_{\mathrm{cons}}\mathfrak{B}$ and $\mathfrak{B}$ fixes each $a\in\omega^{\mathfrak{A}},$ then $\mathfrak{A}$ is cofinal in $\mathfrak{B}$. As shown in Theorem 3.1 of this paper, it is possible to arrange taller conservative elementary extensions of models of $\mathrm{ZFC}$, thus Theorem 2.2 of [En-3] is false as stated. However, by ‘slightly’ strengthening the condition:
$\mathfrak{B}$ fixes each $a\in\omega^{\mathfrak{A}},$
to
$\mathfrak{B}$ fixes $\omega^{\mathfrak{A}},$
one obtains a true statement, as indicated in Corollary 4.2 of this paper. As pointed out in Remark 3.9, the former condition holds as soon whenever $\mathfrak{A}\prec_{\mathrm{cons}}\mathfrak{B}\models\mathrm{ZF}$.
Several corollaries were drawn from Theorem 2.2 of [En-3]; what follows are the modifications in them, necessitated by the above correction. We assume that the reader has [En-3] handy for ready reference.
  • In Corollary 2.3(a) of [En-3], the assumption that $\mathfrak{B}$ fixes every integer of $\mathfrak{A}$ should be strengthened to: $\mathfrak{B}$ fixes $\omega^{\mathfrak{A}}$.
  • Part (b) of Corollary 2.3 of [En-3] is correct as stated in [En-3]: if $\mathfrak{A}$ is a Rubin model of $\mathrm{ZFC}$ and $\mathfrak{A}\prec\mathfrak{B}$ that fixes every $a\in\omega^{\mathfrak{A}}$, then $\mathfrak{B}$ is a conservative extension of $\mathfrak{A}$. This is really what the proof of part (a) of Corollary 2.3 of [En-3] shows. However, the following correction is needed: towards the end of the proof of part (a) of Corollary 2.3 of [En-3], after defining the set $X$, it must simply be stated as follows: But $X=Y\cap A,$ where $Y$ is the collection of members $a\in A$ satisfying $[\varphi(b,a)\wedge a\in R_{\alpha}]$. Since $\mathfrak{B}$ is an end extension of $\mathfrak{C}$, $Y\in C,$ so by claim ()($\ast$)( ∗ ), $X$ is definable in $\mathfrak{A}$.
  • Corollary 2.4 of [En-3] is correct as stated since it does not rely on the full force of Theorem 2.2 of [En-3] and only relies on the fact that no elementary conservative extension of a model of $\mathrm{ZFC}$ is an end extension. The latter follows from Corollary 4.2 of this paper.
  • The condition stipulating that $\mathfrak{B}$ fixes $\omega^{\mathfrak{A}}$ must be added to condition $(i)$ of Corollary 2.5 of [En-3].
APPENDIX: PROOFS OF THEOREMS 5.15 & 5.17
A.1. Theorem (Rubin-Shelah-Schmerl) It is a theorem of $\mathrm{ZFC}$ that every countable model of $\mathrm{ZF}$ has an elementary extension to a weakly Rubin model of cardinality $\aleph_{1}$.
The proof of Theorem A.1 has two distinct stages: in the first stage we prove within $\mathrm{ZFC}+\Diamond_{\omega_{1}}$ that every countably infinite structure in a countable language has an elementary extension to a Rubin model of cardinality $\aleph_{1}$ (Theorem 5.15), and then in the second stage we use the result of the first stage together with a forcing-and-absoluteness argument to establish the theorem. [^23] To my knowledge it is not known whether Theorem 5.15 (existence of Rubin elementary extensions) can be proved in $\mathrm{ZFC}$ alone. However, as shown here, $\mathrm{ZFC}$ can prove that every countable model of $\mathrm{ZF}$ has an elementary extension to a weakly Rubin model. See also Remark A.7.
Stage 1 of the Proof of Theorem A.1
Fix a $\Diamond_{\omega_{1}}$ sequence $\left\langle S_{\alpha}:\alpha<\omega_{1}\right\rangle$.  Given a countable language $\mathcal{L}$, and countably infinite $\mathcal{L}$-structure model $\mathcal{M}_{0}$, assume without loss of generality that $M_{0}=\alpha_{0}\in\omega_{1}.$ We plan to inductively build two sequences $\left\langle\mathcal{M}_{\alpha}:\alpha<\omega_{1}\right\rangle$, and $\left\langle\mathcal{O}_{\alpha}:\alpha<\omega_{1}\right\rangle.$ The first is a sequence of approximations to our final model $\mathcal{M}.$ The second sequence, on the other hand, keeps track of the increasing list of ‘obligations’ we need to abide by throughout the construction of the first sequence. More specifically, each $\mathcal{O}_{\alpha}$ will be of the form $\{\{V_{n}^{\alpha},\ W_{n}^{\alpha}\}:n\in\mathbb{\omega}\}$, where $\{V_{n}^{\alpha},\ W_{n}^{\alpha}\}$ is pair of disjoint subsets $M_{\alpha}$ that are inseparable in $\mathcal{M}_{\alpha}$ and should be kept inseparable in each $\mathcal{M}_{\beta}$, for all $\beta\geq\alpha.$ Here we say that two disjoint subsets $V$ and $W$ of a model $\mathcal{N}$ are inseparable in $\mathcal{N}$ if there is no $\mathcal{N}$-definable $X\subseteq N$ such that $V\subseteq X,$ and $W\cap X=\varnothing.\mathfrak{\vskip 6.0pt plus 2.0pt minus 2.0pt}$
We only need to describe the construction of these two sequences for stages $\alpha+1$ for limit ordinals $\alpha$ since:
  • $\mathcal{O}_{0}:=\mathcal{\emptyset}$.
  • For limit $\alpha$, $\mathcal{M}_{\alpha}:=\bigcup\limits_{\beta<\alpha}\mathcal{M}_{\beta}$ and $\mathcal{O}_{\alpha}:=\bigcup\limits_{\beta<\alpha}\mathcal{O}_{\beta}$.
  • For nonlimit $\alpha$ , $\mathcal{M}_{\alpha+1}:=\mathcal{M}_{\alpha}$ and $\mathcal{O}_{\alpha+1}:=\mathcal{O}_{\alpha}$.
At stage $\alpha+1,$ where $\alpha$ is a limit ordinal, we have access to a model $\mathcal{M}_{\alpha}$ (where $M_{\alpha}\in\omega_{1}$), and a collection $\mathcal{O}_{\alpha}$ of inseparable pairs of subsets of $M_{\alpha}$. We now look at $S_{\alpha}$, and consider two cases: either $S_{\alpha}$ is parametrically undefinable in $\mathfrak{A}_{\alpha},$ or not. In the latter case we “do nothing” and define $\mathcal{M}_{\alpha+1}:=\mathcal{M}_{\alpha}$ and $\mathcal{O}_{\alpha+1}:=\mathcal{O}_{\alpha}.$ But if the former is true, we augment our list of obligations via:
$\mathcal{O}_{\alpha+1}:=\mathcal{O}_{\alpha}\cup\{\left\{S_{\alpha},\ M_{\alpha}\backslash S_{\alpha}\right\}\}.\mathfrak{\vskip 6.0pt plus 2.0pt minus2.0pt}$
Notice that if $S_{\alpha}$ is parametrically undefinable in $\mathcal{M}_{\alpha},$ then $\{S_{\alpha},\ M_{\alpha}\backslash S_{\alpha}\}$ is inseparable in $\mathcal{M}_{\alpha}.$ Then we use Lemma A.2 below to build an elementary extension $\mathcal{M}_{\alpha+1}$ of $\mathcal{M}_{\alpha}$ that satisfies the following three conditions:
(1) For each $\{V,\ W\}\in\mathcal{O}_{\alpha+1}$, $V$ and $W$ are inseparable in $\mathcal{M}_{\alpha+1}$.
(2) For every $\mathcal{M}_{\alpha}$-definable directed set $\mathbb{D}$ with no last element, there is some $d^{\ast}\in\mathbb{D}^{\mathcal{M}_{\alpha+1}}$ such that for each $d\in\mathbb{D}^{\mathcal{M}_{\alpha}},$ $\mathcal{N}\models\left(d<_{\mathbb{D}_{n}}d^{\ast}\right).$
(3) $M_{\alpha+1}=M_{\alpha}+\omega$ (ordinal addition).
This concludes the description of the sequences $\left\langle\mathcal{M}_{\alpha+1}:\alpha<\omega_{1}\right\rangle$, and $\left\langle\mathcal{O}_{\alpha}:\alpha<\omega_{1}\right\rangle.$
Let $\mathcal{M}:=\bigcup\limits_{\alpha<\omega_{1}}\mathcal{M}_{\alpha},$ and $\mathcal{O}:=\bigcup\limits_{\alpha<\omega_{1}}\mathcal{O}_{\alpha}$; note that:
$(4)$ For each $\{V,\ W\}\in\mathcal{O}$, $V$ and $W$ are inseparable in $\mathcal{M}$.
We now verify that $\mathcal{M}$ is a Rubin model. Suppose, on the contrary, that for some $\mathcal{M}$-definable partial order $\mathbb{P}$, there is a maximal filter $F\subseteq\mathbb{P}$ that is not $\mathcal{M}$-definable, and $F$ contains a cofinal $\omega_{1}$-chain $E$. By a standard Löwenheim-Skolem argument there is some limit $\alpha<\omega_{1}$ such that
$(5)$ $\ \ (\mathcal{M}_{\alpha},S_{\alpha})\prec(\mathcal{M},S),$ where $S_{\alpha}=S\cap\alpha.$
In particular,
(6) $S_{\alpha}$ is an undefinable subset of $\mathcal{M}_{\alpha}.$
Since $S_{\alpha}$ is a countable subset of some $\mathcal{M}$-definable partial order $\mathbb{P}$, and $E$ is chain of length $\omega_{1}$ that is cofinal in $\mathbb{P}$, there is some $e\in\mathbb{P}$ such that $p<_{\mathbb{P}}e$ for each $p\in S_{\alpha}.$ Fix some $\beta<\omega_{1}$ such that $e\in M_{\beta}.$ We claim that the formula $\varphi(x)=(x<_{P}e)$ separates $S_{\alpha}$ and $\mathbb{P}^{M_{\alpha}}\backslash S_{\alpha}$ in $\mathcal{M}_{\alpha}.$ This is easy to see since if $q\in\mathbb{P}^{M_{\alpha}}\backslash S_{\alpha}$, then by (5), $S_{\alpha}$ is a maximal filter on $\mathbb{P}^{M_{\alpha}}$, and therefore there is some $p\in S_{\alpha}$ such that $\mathcal{M}_{\alpha}\models$$p$ and $q$ have no common upper bound”, which by (5) implies the same to be true in $\mathcal{M}$, in turn implying that $\lnot(q<_{\mathbb{P}}e),$ as desired. We have arrived at a contradiction since on one hand, based on (2) and (6), the formula $\varphi(x)$ witnesses the separability of $S_{\alpha}$ and $M_{\alpha}\backslash S_{\alpha}$ within $\mathcal{M},$ and on the other hand $\{S_{\alpha},\ M_{\alpha}\backslash S_{\alpha}\}\in\mathcal{O}$ by (5), and therefore (4) dictates that $S_{\alpha}$ and $M_{\alpha}\backslash S_{\alpha}$ are inseparable in $\mathcal{M}.$ Thus the proof of Theorem A.1 will be complete once we establish Lemma A.2 below.
A.2. Lemma. Suppose $\mathcal{M}$ is a countable $\mathcal{L}$-structure, where $\mathcal{L}$ is countable. Let $\{\{V_{n},W_{n}\}:n\in\mathbb{\omega}\}$ is a countable list of $\mathcal{M}$-inseparable pairs of subsets of $M$, and $\left\{\left(\mathbb{D}_{n},\leq_{\mathbb{D}_{n}}\right):n\in\mathbb{\omega}\right\}$ be an enumeration of $\mathcal{M}$-definable directed sets with no last element. Then there exists an elementary extension $\mathcal{N}$ of $\mathcal{M}$ that satisfies the following two properties:
(a) $V_{n}$ and $W_{n}$ remain inseparable in $\mathcal{N}$ for all $n\in\mathbb{\omega}$.
(b) For each $n\in\mathbb{\omega}$ there is some $d_{n}\in N$ such that $\mathcal{N}\models d_{n}\in\mathbb{D}_{n}$ (i.e., $d_{n}$ satisfies the formula that defines $\mathbb{D}_{n}$ in $\mathcal{M}$), and for each $d\in\mathbb{D}_{n},$ $\mathcal{N}\models\left(d<_{\mathbb{D}_{n}}d_{n}\right).$
Let $\mathcal{L}^{+}$ be the language $\mathcal{L}$ augmented with constants $\left\{d_{n}:n\in\mathbb{\omega}\right\}\cup\{m:m\in M\}$, and consider the $\mathcal{L}^{+}$-theory $T$ below:
$T:=\mathrm{Th}(\mathcal{M},m)_{m\in M}+\left\{d_{n}\in\mathbb{D}_{n}:n\in\mathbb{\omega}\right\}\cup\left\{d<_{\mathbb{D}_{n}}d_{n}:n\in\mathbb{\omega},\text{\ }d\in\mathbb{D}_{n}\right\}.$
$T$ is readily seen to be consistent since it is finitely satisfiable in $\mathcal{M}\mathfrak{.}$ Moreover, it is clear that if $\mathcal{N}\models T$, then $\mathcal{M}\prec\mathcal{N}$ and $\mathcal{N}$ satisfies condition (b) of the theorem. To arrange a model of $T$ in which conditions (a) also holds requires a delicate omitting types argument. First, we need a pair of preliminary lemmas. The proof of Lemma A.3 is routine and left to the reader, and Lemma A.4 is an immediate corollary of Lemma A.3.
A.3. Lemma. The following two conditions are equivalent for a sentence $\varphi(d_{k_{1}},\cdot\cdot\cdot,d_{k_{n}})$ of $\mathcal{L}^{+}$.
$(i)$   $T\vdash\varphi(d_{k_{1}},\cdot\cdot\cdot,d_{k_{n}})$.
$(ii)$ $\ \ \mathcal{M}\models\exists r_{1}\in\mathbb{D}_{k_{1}}\forall s_{1}\in\mathbb{D}_{k_{1}}\cdot\cdot\cdot\exists r_{b}\in D_{k_{n}}\forall s\in\mathbb{D}_{k_{n}}\left[\left(\bigwedge\limits_{1\leq i\leq n}r_{i}<_{\mathbb{D}_{k_{n}}}s_{i}\right)\rightarrow\varphi(s_{1},\cdot\cdot\cdot,s_{n})\right].$
A.4. Lemma. The following two conditions are equivalent for a sentence $\varphi(d_{k_{1}},\cdot\cdot\cdot,d_{k_{n}})$ of $\mathcal{L}^{+}$.
$(i)$   $T+\varphi(d_{k_{1}},\cdot\cdot\cdot,d_{k_{n}})$ is consistent.
$(ii)$ $\ \ \mathcal{M}\models\forall r_{1}\in\mathbb{D}_{k_{1}}\exists s_{1}\in\mathbb{D}_{k_{1}}\cdot\cdot\cdot\forall r_{b}\in\mathbb{D}_{k_{n}}\exists s\in\mathbb{D}_{k_{n}}\left[\left(\bigwedge\limits_{1\leq i\leq n}r_{i}<_{\mathbb{D}_{k_{n}}}s_{i}\right)\wedge\varphi(s_{1},\cdot\cdot\cdot,s_{n})\right].$
We are now ready to carry out an omitting types argument to complete the proof of Lemma A.2. For each formula $\psi(y,\overline{x})$ of $\mathcal{L}^{+}$, and each $n\in\omega$, consider the following $p$-type formulated in the language $\mathcal{L}^{+}$, where $\overline{x}=\left\langle x_{1},...,x_{p}\right\rangle$ is a $p$-tuple of variables:
$\Sigma_{n}^{\psi}(\overline{x}):=\{\psi(a,\overline{x}):a\in V_{n}\}\cup\{\lnot\psi(b,\overline{x}):b\in W_{n}\}.$
Note that $\Sigma_{n}^{\psi}$ expresses that $\psi(y,\overline{x})$ separates $V_{n}$ and $W_{n}.$
A.5. Lemma. $\Sigma_{n}^{\psi}$ is locally omitted by $T$ for each formula $\psi(y,\overline{x})$, and each $n\in\mathbb{\omega}$.
Proof. Suppose to the contrary that there is a formula $\theta(\overline{x},d_{k_{1}},\cdot\cdot\cdot,d_{k_{n}})$ of $\mathcal{L}^{+}$ and some $n\in\omega$ such that (1) through (3) below hold:
(1) $T+\exists\overline{x}\ \theta(\overline{x},c)$ is consistent.
(2) For all $a\in V_{n},$ $T\vdash\forall\overline{x}\ \left[\theta(\overline{x},d_{k_{1}},\cdot\cdot\cdot,d_{k_{n}})\rightarrow\psi(a,\overline{x})\right].$
(3) For all $b\in W_{n},$ $T\vdash\forall\overline{x}\ \left[\theta(\overline{x},d_{k_{1}},\cdot\cdot\cdot,d_{k_{n}})\rightarrow\lnot\psi(b,\overline{x})\right].$
Invoking Lemmas A.3 and A.4, (1) through (3) translate to $(1^{\prime})$ through $(3^{\prime})$ below:
$(1^{\prime})$ $\ \mathcal{M}\models\forall r_{1}\in\mathbb{D}_{k_{1}}\exists s_{1}\in\mathbb{D}_{k_{1}}\cdot\cdot\cdot\forall r_{b}\in\mathbb{D}_{k_{n}}\exists s_{n}\in\mathbb{D}_{k_{n}}\left[\left(\bigwedge\limits_{1\leq i\leq n}r_{i}<_{D_{n}}s_{i}\right)\wedge\exists\overline{x}\ \theta(\overline{x},\overline{s})\right],$ where $\overline{s}=\left\langle s_{1},...,s_{n}\right\rangle.$
$(2^{\prime})$ For all $a\in V_{n}$, $\mathcal{M}\models\lambda(a),$ where:
$\lambda(\mathbf{v}):=\forall\overline{x}\ \exists r_{1}\in\mathbb{D}_{k_{1}}\forall s_{1}\in\mathbb{D}_{k_{1}}\cdot\cdot\cdot\exists r_{b}\in\mathbb{D}_{k_{n}}\forall s_{n}\in\mathbb{D}_{k_{n}}\left[\left(\bigwedge\limits_{1\leq i\leq n}r_{i}<_{D_{n}}s_{i}\right)\rightarrow\left(\theta(\overline{x},\overline{s})\rightarrow\psi(\mathbf{v},\overline{x})\right)\right].$
$(3^{\prime})$ For all $b\in W_{n},\ \mathcal{M}\models\gamma(b)$, where:
$\gamma(\mathbf{w}):=\forall\overline{x}\ \exists r_{1}\in\mathbb{D}_{k_{1}}\forall s_{1}\in\mathbb{D}_{k_{1}}\cdot\cdot\cdot\exists r_{b}\in\mathbb{D}_{k_{n}}\forall s_{n}\in\mathbb{D}_{k_{n}}\left[\left(\bigwedge\limits_{1\leq i\leq n}r_{i}<_{D_{n}}s_{i}\right)\rightarrow\left(\theta(\overline{x},\overline{s})\rightarrow\lnot\psi(\mathbf{w},\overline{x})\right)\right].$
Let $\Lambda:=\{m\in M:\mathcal{M}\models\lambda(m)\}$, and $\Gamma:=\{m\in M:\mathcal{M}\models\gamma(m)\}$, and observe that $V_{n}\subseteq\Lambda$ by $(2^{\prime})$ and $W_{n}\subseteq\Gamma$ by $(3^{\prime}$). To arrive at a contradiction, we will show that $\Lambda\cap\Gamma=\emptyset$, which implies that $V_{n}$ and $W_{n}$ are separable in $\mathcal{M}$. To this end, suppose to the contrary that for some $m\in M,$
(4) $\mathcal{M}\models\lambda(m)\wedge\gamma(m).$
It is easy to see, using the fact that each $\left(\mathbb{D}_{n},\leq_{\mathbb{D}_{n}}\right)$ is a directed that (4) implies:
(5)))) $\mathcal{M}\models\exists r_{1}\in\mathbb{D}_{k_{1}}\forall s_{1}\in\mathbb{D}_{k_{1}}\cdot\cdot\cdot\exists r_{b}\in\mathbb{D}_{k_{n}}\forall s\in\mathbb{D}_{k_{n}}\left[\left(\bigwedge\limits_{1\leq i\leq n}r_{i}<_{\mathbb{D}_{n}}s_{i}\right)\rightarrow\left(\psi(m,\overline{x})\wedge\lnot\psi(m,\overline{x})\right)\right].$
Since each $\left(\mathbb{D}_{n},\leq_{\mathbb{D}_{n}}\right)$ is a directed set with no last element, (5) yields a contradiction, so the proof is complete. $\square$ (Lemma A.5)
Proof of Lemma A.2. Put Lemma A.5 together with the Henkin-Orey omitting types theorem [CK, Theorem 2.2.9] to conclude that there exists a model $\mathcal{N}$ of $T$ that satisfies properties (a) and (b). $\square$
Stage 2 of the Proof of Theorem A.1
In this stage we employ the result of the first stage together with some set-theoretical considerations to establish the Rubin-Shelah-Schmerl Theorem. The main idea here is a variant of the one used by Schmerl, [Schm-2], which itself based on a method of $\diamondsuit_{\mathbb{\aleph}_{1}}$-elimination introduced by Shelah [Sh]. This stage has three steps.
Step 1 of Stage 2. It is well-known that there is an $\omega_{1}$-closed notion $\mathbb{Q}_{1}$ in the universe $\mathrm{V}$ of $\mathrm{ZFC}$ such that the $\mathbb{Q}_{1}$-generic extension $\mathrm{V}^{\mathbb{Q}_{1}}$ of $\mathrm{V}$ satisfies $\mathrm{ZFC}+\diamondsuit_{\mathbb{\aleph}_{1}}$ [Ku, Theorem 8.3] (note that $\aleph_{1}$ is absolute between $\mathrm{V}$ and $\mathrm{V}^{\mathbb{Q}_{1}}$ since $\mathbb{Q}_{1}$ is $\omega_{1}$-closed). Thus by Step 1, given a model $\mathcal{M}$ in the universe $\mathrm{V}$ of $\mathrm{ZF}$, the forcing extension $\mathrm{V}^{\mathbb{Q}_{1}}$ satisfies “there is an elementary extension $\mathcal{N}$ of $\mathcal{M}$ such that $\mathcal{N}$ is a Rubin model”. In $\mathrm{V}^{\mathbb{Q}_{1}}$, since $\mathcal{N}$ is a Rubin model, for each $s\in N$ that is infinite in the sense of $\mathcal{N}$, there is a subset $C_{s}=\left\{c_{\alpha}^{s}:\alpha\in\omega_{1}\right\}$ of $\mathcal{N}$ such that $\left\langle c_{\alpha}^{s}:\alpha\in\omega_{1}\right\rangle$ is an $\omega_{1}$-chain that is cofinal in the directed set $\left([s]^{<\omega}\right)^{\mathcal{M}}.$
  • Let $\mathbb{P}_{s}$ be the suborder of $\left(\mathrm{Fin}(s,2)\right)^{\mathcal{M}}$ consisting of partial $\mathcal{M}$-finite functions $f$ from $s$ into $2$ such that $\mathrm{dom}(f)\in C_{s}$. Note that $\mathbb{P}_{s}$ is a tree-order (the predecessors of any given element are comparable).
  • $\mathbb{P}_{s}$ can be turned into a ranked-tree $\tau_{s}$ by defining $\rho(f)=\alpha$ if $\mathrm{dom}(f)=c_{\alpha}$. The ranked tree $\tau_{s}$ has the key property that each maximal filter of $\left(\mathrm{Fin}(s,2)\right)^{\mathcal{M}}$ is uniquely determined by a branch of $\tau_{s}.$
  • Let $C$ be the binary predicate on $\mathcal{N}$ defined by $C(s,x)$ iff $x\in C_{s}$, and let
$\mathcal{N}^{+}=\left(\mathcal{N},C\right)$.
Note that the $\mathcal{N}^{+}$-definable ranked trees include trees of the form $\tau_{s}$, as well as the ranked tree $\tau_{\mathrm{class}}^{\mathcal{N}}$ (whose branches uniquely determine piecewise coded subsets of $\mathcal{N}$, as noted in Remark 5.14).
  • This gives us $\aleph_{1}$-many ranked trees ‘of interest’ in $\mathrm{V}^{\mathbb{Q}_{1}},$ namely $\tau_{\mathrm{class}}^{\mathcal{N}}$ together with $\tau_{s}$ for each infinite set $s$ of $\mathcal{N}$.
Step 2 of Stage 2. The concept of a weakly specializing function plays a key role in this step. Given a ranked tree $(\mathbb{T},\ \leq_{\mathbb{T}},\ \mathbb{L},\ \leq_{\mathbb{L}},\ \rho)$, we say that $f:\mathbb{T\rightarrow\omega}$ weakly specializes $(\mathbb{T},\ \leq_{\mathbb{T}})$ if $f$ has the following property:
If $x\leq_{\mathbb{T}}y$ and $x\leq_{\mathbb{T}}z$ and $f(x)=f(y)=f(z),$ then $y$ and $z$ are $\leq_{\mathbb{T}}$-comparable.
The following lemma captures the essence of a weakly specializing function for our purposes.
A.6. Lemma. Suppose $\tau=(\mathbb{T},\ \leq_{\mathbb{T}},\ \mathbb{L},\ \leq_{\mathbb{L}},\ \rho)$ is a ranked tree, where $(\mathbb{L},\ \leq_{\mathbb{L}})$ has cofinality $\omega_{1}$, and $f$ weakly specializes $(\mathbb{T},\ \leq_{\mathbb{T}})$. Then every branch of $\tau$ is parametrically definable in the expansion $\left(\tau,f\right)$ of $\tau.$
Proof. Suppose $B$ is a branch of $\tau$. Then since ($\mathbb{L},\ \leq_{\mathbb{L}})$ has cofinality $\omega_{1},$ there is a subset $B_{0}$ of $B$ of order type $\omega_{1}$ that is cofinal in $B$. Therefore there is some unbounded $S\subseteq\omega_{1}$ such that $f$ is constant on $\left\{f(b):b\in B_{0}\right\}\}.$ Fix some element $b_{0}\in B_{0}$ and consider subset $X$ of $\mathbb{T}$ that is defined in $\left(\tau,f\right)$ by the formula:
$\varphi(x,b_{0}):=[x\leq_{\mathbb{T}}b_{0}\vee(b_{0}\leq_{\mathbb{T}}x\wedge f(b_{0})=f(x))].$
By the defining property of weakly specializing functions, $X$ is linearly ordered and is cofinal in $B$. This makes it clear that $B$ is definable in $\left(\tau,f\right)$ as the downward closure of $X$, i.e., by the formula $\psi(x,b_{0}):=\left[\exists y(\varphi(y,b_{0})\wedge x\leq_{\mathbb{T}}y\right].$square
By a remarkable theorem, due independently to Baumgartner [Bau] and Shelah [Sh], given any ranked tree $\tau$ of cardinality $\aleph_{1}$ and of cofinality $\omega_{1}$, there is a c.c.c. notion of forcing $\mathbb{Q}_{\tau}$ such that $\mathrm{V}^{\mathbb{Q}_{1}\ast\mathbb{Q}_{\tau}}:=\left(\mathrm{V}^{\mathbb{Q}_{1}}\right)^{\mathbb{Q}_{\tau}}$ contains a function $f_{\tau}$ such that weakly specializes $\tau.$ [^24] This result generalizes the fact, established by Baumgartner, Malitz, and Reinhardt [BMR] that every $\omega_{1}$-Aronszajn tree can be turned into a special Aronszajn tree in a c.c.c. forcing extension of the universe. The forcing condition in $\mathbb{Q}_{\tau}$ consists of all finite attempts to build a function that weakly specializes $\tau$. Moreover, given any collection $\{\tau_{\alpha}:\alpha\in\kappa\}$ of ranked trees of cofinality $\omega_{1}$, the finite support product $\mathbb{Q}$ of $\left\{\mathbb{Q}_{\tau_{\alpha}}:\alpha\in\kappa\right\}$ is also known to have the c.c.c. property [^25] It is well-known that the finite support product of a given collection of forcing notions $\left\{\mathbb{P}_{\alpha}:\alpha\in\kappa\right\}$ has the c.c.c. property iff every finite sub-product of $\left\{\mathbb{P}_{\alpha}:\alpha\in\kappa\right\}$ has the c.c.c. property., as indicated in [CZ, Corollary 3.3]262626[CZ, Corollary 3.3] is overtly about specializing $\omega_{1}$-Aronszajn trees, but as shown by Baumgartner [Bau, Section 7], given a tree $\tau$ of height $\omega_{1}$ with at most $\aleph_{1}$ many branches, there is an Aronszajn subtree $\tau_{0}$ of $\tau$ of height $\omega_{1}$, and moreover, any function $f_{0}$ that specializes $\tau_{0}$ can be canonically extended to a function $f$ that specializes $\tau.$ For more detail, see the proof of [Sw, Lemma 3.5].. Note that for each $\alpha\in\kappa$, $\mathbb{Q}$ forces the existence of a function $f_{\tau_{\alpha}}$ that weakly specializes $\tau_{\alpha}.$
  • Thus by choosing $\{\tau_{\alpha}:\alpha\in\omega_{1}\}$ to be the collection of ranked trees that are parametrically definable in $\mathcal{N}^{+}$, there is a function $g(s,x)$ in $\mathrm{V}^{\mathbb{Q}_{1}\ast\mathbb{Q}}:=\left(\mathrm{V}^{\mathbb{Q}_{1}}\right)^{\mathbb{Q}}$ such that for each infinite set $s$ of $\mathcal{N}$, $g(s,x)$ weakly specializes $\tau_{s}$ ($\tau_{s}$ was defined in Step 1 of Stage 2), and a function $h$ that weakly specializes $\tau_{\mathrm{class}}^{\mathcal{N}}.$
The next lemma shows that forcing with $\mathbb{Q}$ does not add new branches to ranked trees in the ground model that have cofinality $\omega_{1}$. Note that when the lemma below is applied to the case when V is replaced by $\mathrm{V}^{\mathbb{Q}_{1}}$, it shows that the ample supply of rather branchless models in $\mathrm{V}^{\mathbb{Q}_{1}}$ remain rather branchless after forcing with $\mathbb{Q}$, i.e., in $\mathrm{V}^{\mathbb{Q}_{1}\ast\mathbb{Q}}.$
A.7. Lemma. Suppose $\tau=(\mathbb{T},\ \leq_{\mathbb{T}},\ \mathbb{L},\ \leq_{\mathbb{L}},\ \rho)$ is a ranked tree, where $(\mathbb{L},\ \leq_{\mathbb{L}})$ has cofinality $\omega_{1}.$ If $B$ is a branch of $\tau$ in $\mathrm{V}^{\mathbb{Q}}$, then $B\in\mathrm{V}$.
Proof. Suppose $B$ is a branch of $\tau$ in $\mathrm{V}^{\mathbb{Q}}$, and let $f$ be the $\mathbb{Q}$-generic function in $\mathrm{V}^{\mathbb{Q}}$ that specializes $\tau.$ Since $\mathbb{Q}$ does not collapse $\omega_{1}$ and the rank order of $\tau$ is assumed to have cofinality $\omega_{1}$ in $\mathrm{V}$, there is for some $k\in\omega$ such that:
(1) $\ \ \mathrm{V}^{\mathbb{Q}}\models$$f^{-1}\{k\}\cap B$ is unbounded in $B$”.
In $\mathrm{V}^{\mathbb{Q}},$ let $x_{0}\in B$ with $f(x_{0})=k.$ Coupled with (1) this shows that there is some $q_{0}\in\mathbb{Q}$ such that:
(2) $\ \ \ q_{0}\Vdash$ $\left[f(x_{0})=k\wedge\exists y>_{\mathbb{T}}x_{0}(y\in B\wedge f(y)=k)\right]$.
Let $B_{0}$ be the subset of $\mathbb{T}$ defined in $\mathrm{V}$ as:
$B_{0}=\left\{y\in\mathbb{T}:\exists q\in\mathbb{Q\ }(q\supseteq q_{0}\wedge q\Vdash\left[y>_{\mathbb{T}}x_{0}\wedge y\in B\wedge f(y)=k\right]\right\}.$
We wish to show:
$(\ast)\ \ \mathrm{V}^{\mathbb{Q}}\models$ $B_{0}$ is a cofinal subset of $B$.
To verify ()($\ast$)( ∗ ), suppose $b_{0}\in B$. By (1), for some $y\in\mathbb{T}$, $\mathrm{V}^{\mathbb{Q}}$ satisfies “$y>_{\mathbb{T}}x_{0}$, $y\in B$, and $f(y)=k$. Therefore there is some $\mathbb{Q}$-condition $q\supseteq q_{0}$ such that:
[…]
Since $B_{0}\in\mathrm{V}$, ()($\ast$)( ∗ ) makes it clear that $B\in\mathrm{V}$ as well, since $B$ can be defined in $\mathrm{V}$ as the downward closure of $B_{0}.$square
  • Let $\Omega$ be the collection consisting of the ranked tree $\tau_{\mathrm{set}}^{\mathcal{N}}$ together with the family of ranked trees $\left\{\tau_{s}:N\models\left|s\right|\geq\aleph_{0}\right\}.$ Since $\mathcal{N}$ is a Rubin model in $\mathrm{V}$, each of the trees in $\Omega$ are rather branchless in $\mathrm{V}$, and Lemma A.7 assures us the ranked trees in $\Omega$ remain rather branchless in $\mathrm{V}^{\mathbb{Q}}$. Lemma A.6, on the other hand, assures us that each branch of the ranked trees in $\Omega$ has a simple definition in $\mathrm{V}^{\mathbb{Q}}.$
  • Let $\mathcal{L}_{\mathrm{set}}^{+}$ be the finite language obtained by augmenting the usual language $\mathcal{L}_{\mathrm{set}}$ of set theory with extra symbols for denoting the binary relation $C$, the binary function $g,$ and the unary function $f$, and let $\mathcal{L}$ be the countable language that is the result of augmenting $\mathcal{L}_{\mathrm{set}}^{+}$ with constants for each element of $\mathcal{M}$. The content of the above bullet item allows us to write a single sentence $\psi$ in $\mathcal{L}_{\mathsf{\omega}_{\mathsf{1}}\mathsf{,\omega}}(\mathsf{Q})$ (where $\mathsf{Q}$ is the quantifier “there exist uncountably many”) that captures the following salient features of the structure $\left(\mathcal{N}^{+},C,g,f\right)$ that hold in $\mathrm{V}^{\mathbb{Q}}$:
  1. (1)
    The $\mathcal{L}_{\mathrm{set}}$-reduct $\left(\mathcal{N}^{+},C,g,f\right)$ elementary extends $\mathcal{M}$.
  2. (2)
    For each infinite $s$, $\tau_{s}$ has uncountable cofinality, is rather branchless, and $g(s,x)$ weakly specializes $\tau_{s}$.
  3. (3)
    $\tau_{\mathrm{set}}^{\mathcal{N}}$ has uncountable cofinality, is rather branchless, and $f$ weakly specializes $\tau_{\mathrm{set}}^{\mathcal{N}}$.
More specifically, $\psi$ is the conjunction of the following sentences:
  • $\psi_{1}=$ the conjunction of the countably many sentences in the elementary diagram of $\mathcal{M}.$
  • $\psi_{2}=$ the sentence expressing (2) above.
  • $\psi_{3}=$ the sentence expressing (3) above.
We elaborate on how to write $\psi_{2}$, a similar idea is used to write $\psi_{3}.$ The quantifier Q allows us to express that $\tau_{s}$ has uncountable cofinality, and the range of the specializing function $g(s,x)$ is countable. To expresses the rather branchless feature of $\tau_{s}$, let $\psi(x,y)$ be as in the proof of Lemma A.6. Note that in the notation of Lemma A.6, for any $b\in\mathbb{T}$$\{x\in\mathbb{T}_{s}:\psi(x,b)\}$ is linearly ordered by $\leq_{\mathbb{T}}.$ By Lemma A.6, in $\mathrm{V}^{\mathbb{Q}},$ if $s$ is infinite and $B$ is a branch of $\tau_{s}$, then there is some $b_{0}\in B$ such that $\psi(x,b_{0})$ defines $B$.  Together with the fact that $\tau_{s}$ is rather branchless in $\mathrm{V}^{\mathbb{Q}}$ (thanks to Lemma A.7) the $\mathcal{L}_{\mathsf{\omega}_{\mathsf{1}}\mathsf{,\omega}}$-formula below (no need for $\mathsf{Q}$ here) expresses the rather branchless feature of $\tau_{s}$. In the formula below $\mathbb{T}_{s}$ is the set of the nodes of the ranked tree $\tau_{s}$, and “unbounded rank” is short for “unbounded $\tau_{s}$-rank”.
$\forall b\in\mathbb{T}_{s}\left[\{x\in\mathbb{T}_{s}:\psi(x,b)\}\ \mathrm{has\ unbounded\ rank\ }\rightarrow\bigvee\limits_{\theta(x,y)\in\mathcal{L}_{\mathrm{set}}^{+}}\exists y\ \forall x\ \left(\psi(x,b)\leftrightarrow\theta(x,y)\right)\right].$
Step 3 of Stage 2. Let $\psi$ be the sentence described above. Since $\psi$ has a model in $\mathrm{V}^{\mathbb{Q}_{1}\ast\mathbb{Q}}$, and by Keisler’s completeness theorem for $\mathcal{L}_{\omega_{1},\omega}(\mathsf{Q})$ [Ke-2], the consistency of sentences in $\mathcal{L}_{\omega_{1},\omega}(\mathsf{Q})$ is absolute for extensions of $\mathrm{V}$ that do not collapse $\aleph_{1},$ we can conclude that $\psi$ has a model in $\mathrm{V}$ since $\aleph_{1}$ is absolute between $\mathrm{V}$ and $\mathrm{V}^{\mathbb{Q}_{1}\ast\mathbb{Q}}$ (recall that $\mathbb{Q}_{1}$ is $\omega_{1}$-closed and $\mathbb{Q}$ is c.c.c.).)).) . Thanks to Lemmas A.6 and A.7, the $\mathcal{L}_{\mathrm{set}}$-reduct of any model of $\psi$ is a weakly Rubin model that elementarily extends $\mathcal{M}$ (and the cardinality of the model can be arranged to be $\aleph_{1}$ by a Löwenheim-Skolem argument). This concludes the proof of Theorem A.1. $\square$
A.7. Remark. Schmerl [Schm-2] used a similar argument as the one used in the proof of Theorem A.1 to show that every countable model $\mathcal{M}$ of $\mathrm{ZFC}$ has an elementary extension $\mathcal{N}$ of cardinality $\aleph_{1}$ such that every $\omega^{\mathcal{N}}$-complete ultrafilter over $\mathcal{N}$ is definable (and thus coded in $\mathcal{N}$). The proof of Theorem A.1 can be dovetailed with Schmerl’s proof so as to show that every countable model $\mathcal{M}$ of $\mathrm{ZFC}$ has an elementary extension $\mathcal{N}$ of cardinality $\aleph_{1}$ that is weakly Rubin and also has the additional property that every $\omega^{\mathcal{N}}$-complete ultrafilter over $\mathcal{N}$ is definable.