### v6

# Models of Set Theory: Extensions and Dead-ends

Abstract.
This paper is a contribution to the study of extensions of arbitrary models of $\mathsf{ZF}$ (Zermelo-Fraenkel set theory), with no regard to countability or well-foundedness of the models involved. Our main results include the theorems below; in Theorems A and B, $\mathcal{N}$ is said to be a conservative elementary extension of $\mathcal{M}$ if $\mathcal{N}$ elementarily extends $\mathcal{M}$, and the intersection of every $\mathcal{N}$-definable set with the universe of $\mathcal{M}$ is $\mathcal{M}$-definable (parameters allowed). In Theorem B, $\mathsf{ZFC}$ is the result of augmenting $\mathsf{ZF}$ with the axiom of choice.
Theorem A. Every model $\mathcal{M}$ of $\mathsf{ZF}+\exists p\left(\mathrm{V}=\mathrm{HOD}(p)\right)$ has a conservative elementary extension $\mathcal{N}$ that contains an ordinal above all of the ordinals of $\mathcal{M}$.
Theorem B. If $\mathcal{N}$ is a conservative elementary extension of a model $\mathcal{M}$ of $\mathsf{ZFC}$, and $\mathcal{N}$ has the same natural numbers as $\mathcal{M}$, then $\mathcal{M}$ is cofinal in $\mathcal{N}$.
Theorem C. Every consistent extension of $\mathrm{ZFC}$ has a model $\mathcal{M}$ of power $\aleph_{1}$ such that $\mathcal{M}$ has no proper end extension to a model of $\mathsf{ZF}$.
1. INTRODUCTION
This paper concerns the model theory of Zermelo-Fraenkel ($\mathsf{ZF}$) set theory. More specifically, it is about various types of extensions of arbitrary models of $\mathsf{ZF}$, as opposed to countable or well-founded ones. This is a topic that I have worked on intermittently for over four decades. Indeed, Theorem C of the abstract answers the following question that was initially posed in my 1984 doctoral dissertation [En-1, Question 1.1.12]:
[…]
Note that if there is a proper class of strongly inaccessible cardinals, then every well-founded model of $\mathsf{ZF}$ can be properly end extended to a model of $\mathsf{ZF}$. Furthermore, by a well-known result due to Keisler and Morley [KM], every model of $\mathsf{ZF}$ whose class of ordinals has countable cofinality has a proper elementary end extension. The motivation for the above question emerges from the comparative study of the model theory of $\mathsf{ZF}$ and $\mathsf{PA}$ (Peano Arithmetic). Prima facie, $\mathsf{ZF}$ and $\mathsf{PA}$ are unrelated; after all $\mathsf{PA}$ axiomatizes basic intuitions concerning the familiar arithmetical structure of natural numbers, whereas $\mathsf{ZF}$ codifies the laws governing Cantor’s mysterious universe of sets, a vastly more complex structure. However, thanks to a clever coding idea, introduced first by Ackermann [A], $\mathsf{ZF}$ and $\mathsf{PA}$ turn out to be intimately connected at a formal level: $\mathsf{PA}$ is bi-interpretable with the theory $\mathsf{ZF}_{\mathrm{fin}}+\mathsf{TC}$, where $\mathsf{ZF}_{\mathrm{fin}}$ is obtained from $\mathsf{ZF}$ by replacing the axiom of infinity by its negation, and $\mathsf{TC}$ expresses “the transitive closure of every set exists”. [^2] The role of $\mathsf{TC}$ was elucidated in [ESV] by showing that $\mathsf{TC}$ cannot be dropped in this bi-interpretability result. The details of this bi-interpretability result are worked out by Kaye and Wong in [KW]. Their work shows that indeed the two theories are definitionally equivalent (aka synonymous), i.e., they have a common definitional extension. In this light it is natural to compare and contrast the model theoretic behavior of models of $\mathsf{ZF}$ and $\mathsf{PA}$ in order to elucidate the role of the axiom of infinity in Zermelo-Fraenkel set theory. A central result in the model theory of $\mathsf{PA}$ is the MacDowell-Specker Theorem [MS], and its refinement (independently) by Phillips [P] and Gaifman [G-2], that states that every model of PA (of any cardinality) has a proper conservative elementary end extension. Using the aforementioned bi-interpretation between $\mathsf{PA}$ and $\mathsf{ZF}$, the refinement of the MacDowell-Specker Theorem is equivalent to:
[…]
In the above, $\mathcal{N}$ is said to be a conservative elementary extension of $\mathcal{M}$ if $\mathcal{N}$ elementarily extends $\mathcal{M}$, and the intersection of every $\mathcal{N}$-definable set with $M$ is $\mathcal{M}$-definable (parameters allowed). It is known that Theorem $\triangledown$ becomes false if $\mathsf{ZF}_{\mathrm{fin}}+\mathsf{TC}$ is replaced by $\mathsf{ZF}$, but the answer to Question $\heartsuit$ (Theorem 5.18) remained open until now. [^3] As shown by Keisler and Silver [KS], if $\kappa$ is the first inaccessible cardinal, then $\left(\mathrm{V}_{\kappa},\in\right)$ does not have a proper elementary end extension. More generally, the results of Kaufmann [Kan] and the author [En-2] showed every consistent extension of $\mathsf{ZFC}$ has a model of power $\aleph_{1}$ that has no proper rank extension to another model of $\mathsf{ZFC}$. For an exposition of the general framework for the analogues of the MacDowell-Specker Theorem for set theory, see [En-5].
The foremost motivation in writing this paper was to present the rather intricate details of the solution to the above 40-year old question to my younger self. Another motivation was to publicize what can be described as a long overdue ‘Additions and Corrections’ to my 1987 paper [En-3], namely Theorems A and B of the abstract (Theorem 3.1 and Corollary 4.2 of the paper, respectively).
Section 2 is devoted to preliminaries needed for Sections 3, 4, and 5 in which the main results are presented. Section 3 contains the proof of Theorem A of the abstract (and closely related material). Section 4 contains results on faithful extensions (a generalization of conservative extensions), including Theorem 4.1 that yields Theorem B of the abstract (as Corollary 4.2). A key result in Section 4 is Theorem 4.4 which is used in Section 5 in conjunction with other machinery to establish Theorem C of the abstract. Section 6 is an Addendum and Corrigendum to my aforementioned 1987 paper [En-3]. The Appendix presents the proof of a theorem due to the collective effort of Rubin, Shelah, and Schmerl that plays a key role in the proof of Theorem C of the abstract.
The main results of the paper were presented online in February 2025 at the CUNY set theory seminar. I am grateful to Vika Gitman and Gunter Fuchs and other participants of the seminar for helpful feedback. Thanks also to Corey Switzer for illuminating discussions about the forcing argument used in the proof of Theorem A.1 (in the Appendix), and to Junhong Chen for catching notational inconsistencies of a previous draft. I am especially indebted to the anonymous referee for meticulous and extensive feedback that was instrumental in shaping the paper in its current form.
2. PRELIMINARIES
Here we collect the basic definitions, notations, conventions, and results that are relevant for the subsequent sections.
2.1. Definitions and Basic Facts (Languages and theories of sets).
  1. (a)
    $\mathcal{L}_{\mathrm{set}}=\{=,\in\}$ is the language of $\mathsf{ZF}$-set theory. We treat $\mathsf{ZF}$ as being axiomatized as usual, except that instead of including the scheme of replacement among the axioms of $\mathsf{ZF}$, we include the schemes of separation and collection, as in [CK, Appendix A]. Thus, in our setup the axioms of Zermelo set theory $\mathsf{Z}$ are obtained by removing the scheme of collection from the axioms of $\mathsf{ZF}$. More generally, for $\mathcal{L}\supseteq\mathcal{L}_{\mathrm{set}}$ we construe $\mathsf{ZF}(\mathcal{L})$ to be the natural extension of $\mathsf{ZF}$ in which the schemes of separation and collection are extended to $\mathcal{L}$-formulae; similarly we will use $\mathsf{Z}(\mathcal{L})$ to denote Zermelo set theory over $\mathcal{L}$, i.e., as the result of extending $\mathsf{Z}$ with the $\mathcal{L}$-separation scheme $\mathsf{Sep}(\mathcal{L})$. Officially speaking, $\mathsf{Sep}(\mathcal{L})$ consists of the universal closures of $\mathcal{L}$-formulae of the form
    $\forall v\exists w\forall x(x\in w\longleftrightarrow x\in v\wedge\varphi(x,\vec{y})).$
    Thus $\mathsf{ZF}(\mathcal{L})$ is the result of augmenting $\mathsf{Z}(\mathcal{L})$ with the $\mathcal{L}$-collection scheme $\mathsf{Coll}(\mathcal{L})$, which consists of the universal closures of $\mathcal{L}$-formulae of the form:
[…]
    When X is a new predicate, we write $\mathcal{L}_{\mathrm{set}}(\mathrm{X})$ instead of $\mathcal{L}_{\mathrm{set}}\cup\{\mathrm{X\}}$, and similarly we write $\mathsf{ZF}(\mathrm{X})$, $\mathsf{Sep}(\mathrm{X})$, $\mathsf{Coll}(\mathrm{X})$, etc. instead of $\mathsf{ZF}(\mathcal{L})$, $\mathsf{Sep}(\mathcal{L})$, $\mathsf{Coll}(\mathcal{L})$, etc. (respectively) when $\mathcal{L}=\mathcal{L}_{\mathrm{set}}(\mathrm{X})$.
  2. (b)
    For $n\in\omega$, we employ the common notation ($\Sigma_{n},$ $\Pi_{n},$ $\Delta_{n}$) for the Levy hierarchy of $\mathcal{L}_{\mathrm{set}}$-formulae, as in the standard references in modern set theory by Kunen [Ku], Jech [J] and Kanamori [Kan]. In particular, $\Delta_{0}=\Sigma_{0}=\Pi_{0}$ corresponds to the collections of $\mathcal{L}_{\mathrm{set}}$-formulae all of whose quantifiers are bounded. In addition, we will also consider the Takahashi hierarchy of formulae ($\Sigma_{n}^{\mathcal{P}},$ $\Pi_{n}^{\mathcal{P}},$ $\Delta_{n}^{\mathcal{P}}$) (introduced in [T], and further studied by Mathias’ [M-1]), where $\Delta_{0}^{\mathcal{P}}$ is the smallest class of $\mathcal{L}_{\mathrm{set}}$-formulae that contains all atomic formulae, and is closed both under Boolean connectives and under quantification in the form $Qx\in y$ and $Qx\subseteq y$ where $x$ and $y$ are distinct variables, and $Q$ is $\exists$ or $\forall$. The classes $\Sigma_{1}^{\mathcal{P}},\Pi_{1}^{\mathcal{P}}$, etc. are defined inductively from the class $\Delta_{0}^{\mathcal{P}}$ in the same way that the formula-classes $\Sigma_{1},\Pi_{1}$, etc. are defined from $\Delta_{0}$-formulae.
  3. (c)
    For $n\in\omega$. and $\mathcal{L}\supseteq\mathcal{L}_{\mathrm{set}},$ the Levy hierarchy can be naturally extended to $\mathcal{L}$-formulae ($\Sigma_{n}(\mathcal{L}),$ $\Pi_{n}(\mathcal{L}),$ $\Delta_{n}(\mathcal{L})$), where $\Delta_{0}(\mathcal{L})$ is the smallest family of $\mathcal{L}$-formulae that contains all atomic $\mathcal{L}$-formulae and is closed under Boolean operations and bounded quantification.
  4. (d)
    $\mathsf{KPR}$ (Kripke-Platek with ranks) is the subtheory of $\mathsf{ZF}$ whose axioms consist of $\mathsf{KP}$ plus an axiom that asserts that for all ordinals $\alpha,$ $\{x:\mathrm{rank}(x)<\alpha\}$ exists (where $\mathrm{rank}(x)$ is the usual rank function on sets). Following recent practice (initiated by Mathias [M-1]), the foundation scheme of $\mathsf{KP}$ is only limited to $\Pi_{1}$-formulae; thus in this formulation $\mathsf{KP}$ can prove the scheme of $\in$-induction for $\Sigma_{1}$-formulae. In contrast to Barwise’s $\mathsf{KP}$ in [Bar], which includes the full scheme of foundation, our version of $\mathsf{KP}$ is finitely axiomatizable. Thus $\mathsf{KPR}$ is finitely axiomatizable.
  5. (e)
[…]
  6. (f)
    For $\mathcal{L}\supseteq\mathcal{L}_{\mathrm{set}}$, the dependent choice scheme, denoted $\Pi_{\infty}^{1}$-$\mathsf{DC}(\mathcal{L})$, consists of the universal closure of $\mathcal{L}$-formulae of the following form:
    $\forall x\exists y\ \varphi(x,y,\vec{z})\longrightarrow\forall x\exists f\ \left[f:\omega\rightarrow\mathrm{V},\ f(0)=x,\mathrm{and}\ \forall n\in\omega\ \varphi(f(n),f(n+1),\vec{z})\right].$
    In the presence of $\mathsf{ZF}(\mathrm{\mathcal{L}})\mathrm{,}$ thanks to the Reflection Theorem (Theorem 2.10), $\Pi_{\infty}^{1}$-$\mathsf{DC}\mathrm{(\mathcal{L})}$ is equivalent to the single sentence $\mathsf{DC}$ below:
    $\forall r\ \forall a\ \left[\begin{array}[]{c}\left(\forall x\in a\ \exists y\in a\ \left\langle x,y\right\rangle\in r\right)\longrightarrow\vskip 6.0pt plus 2.0pt minus 2.0pt\\ \left(\forall x\in a\ \exists f\ \left(f:\omega\rightarrow\mathrm{V},\ f(0)=x,\mathrm{and}\ \forall n\in\omega\ \left\langle f(n),f(n+1)\right\rangle\in r\right)\right)\end{array}\right].$
    Note that $\mathsf{DC}$ is provable in $\mathsf{ZFC}$.
2.2. Definitions and basic facts. (Model theoretic concepts) In what follows we make the blanket assumption that $\mathcal{M}$, $\mathcal{N}$, etc. are $\mathcal{L}$-structures, where $\mathcal{L}\supseteq\mathcal{L}_{\mathrm{set}}$.
  1. (a)
    We follow the convention of using $M$, $M^{\ast},$ $M_{0}$, etc. to denote (respectively) the universes of structures $\mathcal{M}$, $\mathcal{M}^{\ast},$ $\mathcal{M}_{0},$ etc. We denote the membership relation of $\mathcal{M}$ by $\in^{\mathcal{M}}$; thus an $\mathcal{L}_{\mathrm{set}}$-structure $\mathcal{M}$ is of the form $(M,\in^{\mathcal{M}})$.
  2. (b)
    $\mathrm{Ord}^{\mathcal{M}}$ is the class of ordinals in the sense of $\mathcal{M}$, i.e.,
    $\mathrm{Ord}^{\mathcal{M}}:=\left\{m\in M:\mathcal{M}\models\mathrm{Ord}(m)\right\},$
    where $\mathrm{Ord}(x)$ expresses “$x$ is transitive and is well-ordered by $\in$”. More generally, for a formula $\varphi(\vec{x})$ with parameters from $M$, where $\vec{m}=\left(x_{1},\cdot\cdot\cdot,x_{k}\right)$, we write $\varphi^{\mathcal{M}}$ for:
    $\left\{\vec{m}\in M^{k}:\mathcal{M\models\varphi}\left(m_{1},\cdot\cdot\cdot,m_{k}\right)\right\}.$
    A subset $D$ of $M^{k}$ is $\mathcal{M}$-definable if it is of the form $\varphi^{\mathcal{M}}$ for some choice of $\varphi.$ We write $\mathbb{\omega}^{\mathcal{M}}$ for the set of finite ordinals (i.e., natural numbers) of $\mathcal{M}$, and $\mathbb{\omega}$ for the set of finite ordinals in the real world, whose members we refer to as metatheoretic natural numbers. $\mathcal{M}$ is said to be $\omega$-standard if $\left(\mathbb{\omega},\mathbb{\in}\right)^{\mathcal{M}}\cong\left(\mathbb{\omega},\mathbb{\in}\right).$ For $\alpha\in\mathrm{Ord}^{\mathcal{M}}$ we often use $\mathcal{M}_{\alpha}$ to denote the substructure of $\mathcal{M}$ whose universe is
    $M_{\alpha}:=\left\{m\in M:\mathcal{M}\models m\in\mathrm{V}_{\alpha}\right\},$
    where $\mathrm{V}_{\alpha}$ is defined as usual as $\{x:\mathrm{rank}(x)<\alpha\}$, where $\mathrm{rank}(x)=\sup\{\mathrm{rank}(y)+1:y\in x\}.$
(c) For $c\in M$, $\mathrm{Ext}_{\mathcal{M}}(c)$ is the $\mathcal{M}$-extension of $c$, i.e., $\mathrm{Ext}_{\mathcal{M}}(c):=\{m\in M:m\in^{\mathcal{M}}c\}.$
(d) We say that $X\subseteq M$ is coded (in $\mathcal{M)}$ if there is some $c\in M$ such that $\mathrm{Ext}_{\mathcal{M}}(c)=X.$
(e) Suppose $\mathcal{M}\models\mathsf{ZF}$. We say that $X\subseteq M$ is $\mathcal{M}$-amenable if $(\mathcal{M},X)\models\mathsf{ZF}(\mathrm{X})$; here $X$ is the interpretation of $\mathrm{X}$.
2.3. Definitions. (‘geometric shapes’ of extensions). Suppose $\mathcal{L}\supseteq\mathcal{L}_{\mathrm{set}}$, $\mathcal{M}$ and $\mathcal{N}$ are $\mathcal{L}$-structures, and $\mathcal{M}$ a submodel [^6] The notion of a submodel here is the usual one in model theory; i.e., $\mathcal{M}\subseteq\mathcal{N}$ means $M\subseteq N$ and the interpretation of each nonlogical symbol of $\mathcal{L}$ in $\mathcal{M}$ is the restriction to $M$ of the corresponding of interpretation in $\mathcal{N}$ (in particular, $\mathrm{Ext}_{\mathcal{M}}(a)\subseteq\mathrm{Ext}_{\mathcal{N}}(a)$ for all $a\in M).$ of $\mathcal{N}$ (written $\mathcal{M}\subseteq\mathcal{N)}.$
  1. (a)
    $\mathcal{M}^{\ast}$ is the convex hull of $\mathcal{M}$ in $\mathcal{N}$ if $M^{\ast}=\bigcup\limits_{a\in M}\mathrm{Ext}_{\mathcal{N}}(a).$
  2. (b)
    Suppose $a\in M.$ $\mathcal{N}$ fixes $a$ if $\mathrm{Ext}_{\mathcal{M}}(a)=\mathrm{Ext}_{\mathcal{N}}(a)$, and $\mathcal{N}$ enlarges $a$ if $\mathrm{Ext}_{\mathcal{M}}(a)\subsetneq\mathrm{Ext}_{\mathcal{N}}(a).\vskip 3.0pt plus 1.0pt minus 1.0pt$
  3. (c)
    $\mathcal{N}$ end extends $\mathcal{M}$ (written $\mathcal{M}\subseteq_{\mathrm{end}}\mathcal{N}$), if $\mathcal{N}$ fixes every $a\in M$. End extensions are also referred to in the literature as transitive extensions, and in the old days as outer extensions.
  4. (d)
    $\mathcal{N}$ is a powerset-preserving end extension of $\mathcal{M}$ (written $\mathcal{M}\subseteq_{\mathrm{end}}^{\mathcal{P}}\mathcal{N}$) if $(i)$ $\mathcal{M}\subseteq_{\mathrm{end}}\mathcal{N}$, and $(ii)$ for if $a\in M$, $b\in N$, and $\mathcal{N}\models(b\subseteq a)$, then $b\in M$.
  5. (e)
[…]
  6. (f)
    $\mathcal{N}$ is a cofinal extension of $\mathcal{M}$ (written $\mathcal{M}\subseteq_{\mathrm{cof}}\mathcal{N)}$ if for every $b\in N$ there is some $a\in M$ such that $b\in\mathrm{Ext}_{\mathcal{N}}(a)$, i.e., $\mathcal{N}$ is the convex hull of $\mathcal{M}$ in $\mathcal{N}$.
  7. (g)
    $\mathcal{N}$ is taller than $\mathcal{M}$ (written $\mathcal{M}\subseteq_{\mathrm{taller}}\mathcal{N)}$ if there is some $b\in M$ such that $M\subseteq\mathrm{Ext}_{\mathcal{N}}(b).$
2.4. Definitions. (‘logical shapes’ of extensions). Suppose $\mathcal{L}\supseteq\mathcal{L}_{\mathrm{set}}$, and $\mathcal{M}$ and $\mathcal{N}$ are $\mathcal{L}$-structures such that $\mathcal{M}$ is a submodel of $\mathcal{N}$.
  1. (a)
    For a subset $\Gamma$ of $\mathcal{L}$-formulae, $\mathcal{M}$ is a $\Gamma$-elementary submodel of $\mathcal{N}$ (written $\mathcal{M}\preceq_{\Gamma}\mathcal{N}$) if for all $n$-ary formulae $\gamma\in\Gamma$ and for all $a_{1},\cdot\cdot\cdot,a_{n}$ in $M$, $\mathcal{M}\models\gamma(a_{1},\cdot\cdot\cdot,a_{n})$ iff $\mathcal{N}\models\gamma(a_{1},\cdot\cdot\cdot,a_{n}).$ $\mathcal{M}$ is an elementary submodel of $\mathcal{N}$ (written $\mathcal{M}\preceq\mathcal{N}$) if $\mathcal{M}\preceq_{\Gamma}\mathcal{N}$ for $\Gamma=$ $\mathcal{L}$-formulae. We say that $\mathcal{M}$ is an elementary submodel of $\mathcal{N}$ if $\mathcal{M}\preceq_{\Gamma}\mathcal{N}$ for the set $\Gamma$ of all $\mathcal{L}$-formulae (written $\mathcal{M}\preceq\mathcal{N}$). For a given $\gamma\in\Gamma$ we say that $\gamma$ is absolute between $\mathcal{M}$ and $\mathcal{N}$ if $\mathcal{M}\preceq_{\{\gamma\}}\mathcal{N}$. As usual, we write $\mathcal{M}\prec_{\Gamma}\mathcal{N}$ if $\mathcal{M}\preceq_{\Gamma}\mathcal{N}$ and $\mathcal{M}\subsetneq\mathcal{N}$ (similarly for $\mathcal{M}\prec\mathcal{N}$).
  2. (b)
    $\mathcal{N}$ is a self-extension of $\mathcal{M}$ (written $\mathcal{M}\subseteq_{\mathrm{self}}\mathcal{N}$) if $\mathcal{N}$ is interpretable in $\mathcal{M}$ (i.e., $\mathcal{N}$ is isomorphic to a model $\mathcal{N}^{\ast}$ whose universe $N^{\ast}$ of and the interpretation $R^{\mathcal{N}^{\ast}}$ of each $R\in\mathcal{L}$, are all $\mathcal{M}$-definable), and additionally, there is an $\mathcal{M}$-definable embedding $j$ of $\mathcal{M}$ into $\mathcal{N}^{\ast}$. In this context, $\mathcal{N}$ is an elementary self-extension of $\mathcal{M}$ if the map $j$ is an elementary embedding.
  3. (c)
    $\mathcal{N}$ is a conservative extension of $\mathcal{M}$ (written $\mathcal{M}\subseteq_{\mathrm{cons}}\mathcal{N}$) if for every $\mathcal{N}$-definable $D$, $M\cap D$ is $\mathcal{M}$-definable.
  4. (d)
    For $\mathcal{M}\models\mathsf{ZF}$, $\mathcal{N}$ is a faithful extension of $\mathcal{M}$ (written $\mathcal{M}\subseteq_{\mathrm{faith}}\mathcal{N)}$, if for every $\mathcal{N}$-definable $D$, $M\cap D$ is $\mathcal{M}$-amenable (as in Definition 2.2(e)). [^8] The notion of a faithful extension was first introduced and studied in the context of elementary extensions in [En-3].
2.5. Remark. The following are readily verifiable:
  1. (a)
    If $\left\langle\mathcal{N}_{k}:k\in\omega\right\rangle$ is a chain of models extending $\mathcal{M}$ whose union is $\mathcal{N}$, and $\mathcal{M}\preceq_{\Pi_{k,\mathrm{cons}}}\mathcal{N}_{k}\preceq_{\Pi_{k}}\mathcal{N}_{k+1}$ for each $k\in\omega,$ then $\mathcal{M}\preceq_{\mathrm{cons}}\mathcal{N}$. [^9] This fact readily follows from Theorem 2.11. It comes handy in the first proof of Theorem 3.1. Here $\mathcal{M}\preceq_{\Pi_{k,\mathrm{cons}}}\mathcal{N}_{k}$ is shorthand for: $\mathcal{M}\preceq_{\Pi_{k}}\mathcal{N}_{k}$ and $\mathcal{M}\preceq_{{}_{\mathrm{cons}}}\mathcal{N}_{k}$.
  2. (b)
    $\left(\mathcal{M}\subseteq_{\mathrm{self}}\mathcal{N}\right)\Rightarrow\left(\mathcal{M}\subseteq_{\mathrm{cons}}\mathcal{N}\right).\vskip 3.0pt plus 1.0pt minus 1.0pt$
  3. (c)
    Assuming $\mathcal{M}\models\mathsf{ZF}$, $\left(\mathcal{M}\subseteq_{\mathrm{cons}}\mathcal{N}\right)\Rightarrow\left(\mathcal{M}\subseteq_{\mathrm{faith}}\mathcal{N}\right).$
The implications in (b) and (c) are not reversible. More specifically, (b) is not reversible since, e.g., if $\mathcal{N}$ is a self-extension of $\mathcal{M},$ then $\left|M\right|=\left|N\right|,$ but it is possible to build a conservative elementary extension $\mathcal{N}$ of a model $\mathcal{M}$ with $\left|M\right|<\left|N\right|,$ e.g., if $\mathcal{U}$ is an $\mathcal{M}$-ultrafilter, then the ultrapower formation modulo $\mathcal{U}$ can be iterated along any linear order (of arbitrary cardinality) and results in a conservative elementary extension of $\mathcal{M}$. In particular, it is possible to arrange $\mathcal{M}\subseteq_{\mathrm{cons}}\mathcal{N}$ such that $M$ is countable but $N$ is uncountable. Another way to see that $(i)$ is not reversible is to note that by a theorem of Gaifman [G-1] every elementary self-extension of a model $\mathcal{M}$ of the $\mathsf{ZF}$ is a cofinal extension. [^10] Indeed Gaifman’s theorem is stated for models $\mathcal{M}$ of the fragment $\mathsf{ZR}$ of $\mathsf{ZF}$, where $\mathsf{ZR}$ is the result of extending $\mathsf{Z}$ (Zermelo set theory) with an axiom that every element is a member of some set of the form $\mathrm{V}_{\alpha}.$ However, by Theorem 3.1 there are models of $\mathsf{ZF}$ that have taller conservative elementary extensions. To see that (c) is not reversible, note that if $\kappa$ and $\lambda$ are strongly inaccessible cardinals with $\kappa<\lambda$, then $\left(\mathrm{V}_{\kappa},\in\right)$ is faithfully extended by $\left(\mathrm{V}_{\lambda},\in\right)$, but since the truth predicate for $\left(\mathrm{V}_{\kappa},\in\right)$ is parametrically definable in $\left(\mathrm{V}_{\lambda},\in\right)$, by Tarski’s undefinability of truth $\left(\mathrm{V}_{\lambda},\in\right)$ is not a conservative extension of $\left(\mathrm{V}_{\kappa},\in\right)$.
It is also noteworthy that the most common examples of definable extensions in the context of models of set theory are elementary extensions. However, the Boolean-valued approach to forcing provides a wealth of definable extensions that are not elementary. More specifically, if $\mathcal{M}$ is a model of $\mathsf{ZFC}$, $\mathbb{B}$ is an element of $\mathcal{M}$ that is a complete Boolean algebra from the point of view of $\mathcal{M}$, and $\mathcal{U}$ is an ultrafilter on $\mathbb{B}$ that is in $\mathcal{M}$, then there is an $\mathcal{M}$-definable embedding $j$ of $\mathcal{M}$ into the model $\mathcal{M}^{\mathbb{B}}/\mathcal{U}$ (where $\mathcal{M}^{\mathbb{B}}$ is the $\mathbb{B}$-valued forcing extension of the universe, as calculated in $\mathcal{M}$). However, in typical cases $\mathcal{M}$ and $\mathcal{M}^{\mathbb{B}}/\mathcal{U}$ are not elementarily equivalent, e.g., we can always arrange $\mathbb{B}$ so that the truth-value of the continuum hypothesis changes in the transition between the $\mathcal{M}$ and $\mathcal{M}^{\mathbb{B}}/\mathcal{U}$. Furthermore, even though $\mathcal{M}$ and $\mathcal{M}^{\mathbb{B}}/\mathcal{U}$ might be elementarily equivalent in some cases, in general the embedding $j$ is not elementary if $\mathbb{B}$ is a nontrivial notion of forcing, since the statement “there is a filter over $\mathbb{B}$ that meets all the dense subsets of $\mathbb{B}$” holds in $\mathcal{M}^{\mathbb{B}}/\mathcal{U}$, but not in $\mathcal{M}$ (assuming $\mathbb{B}$ is a nontrivial notion of forcing).
2.6. Remark. Suppose $\mathcal{M}$ and $\mathcal{N}$ are $\mathcal{L}_{\mathrm{set}}$-structures with $\mathcal{M}\subseteq\mathcal{N}\mathrm{.}$ The following are readily verifiable from the definitions involved:
  1. (a)
    For each $n\in\omega,$ $\mathcal{M}\preceq_{\Sigma_{n}}\mathcal{N}$ iff $\mathcal{M}\preceq_{\Pi_{n}}\mathcal{N}.\vskip 3.0pt plus 1.0pt minus 1.0pt$
  2. (b)
    $\mathcal{M}\subseteq_{\mathrm{end}}\mathcal{N}$, implies $\mathcal{M}\preceq_{\Delta_{0}}\mathcal{N}$; and $\mathcal{M}\preceq_{\Delta_{0}}\mathcal{N}$ implies $\mathcal{M}\preceq_{\Delta_{1}}\mathcal{N}$.
  3. (c)
    Suppose $\mathcal{M}\preceq_{\Delta_{0}}\mathcal{N}$, $\mathcal{N}$ fixes $a\in M$, and there is bijection $f$ in $\mathcal{M}$ between $a$ and some $b\in M$. Then $\mathcal{N}$ fixes $b$ as well (by $\Delta_{0}$-elementarity, $f$ remains a bijection between $a$ and $b$ as viewed from $\mathcal{N}$).
  4. (d)
    Suppose $\mathcal{M}$ and $\mathcal{N}$ are models of $\mathsf{KPR}$, $\mathcal{M}\preceq_{\Delta_{0}^{\mathcal{P}}}\mathcal{N}$, and $\alpha\in\mathrm{Ord}^{M}$. Then $\mathrm{V}_{\alpha}^{\mathcal{M}}=\mathrm{V}_{\alpha}^{\mathcal{N}}$; this follows from the fact that the formula expressing that $x$ is an ordinal and $y=\mathrm{V}_{x}$ as a $\Sigma_{1}^{\mathcal{P}}$-formula in $\mathsf{KPR}$ (using the usual recursive definition of the $\mathrm{V}_{\alpha}$-hierarchy).
  5. (e)
    If $\mathcal{M}\subseteq_{\mathrm{end}}^{\mathcal{P}}\mathcal{N}$, then $\mathcal{M}\preceq_{\Delta_{0}^{\mathcal{P}}}\mathcal{N}$. On the other hand, since the formula expressing “$x$ is an ordinal and $y=\mathrm{V}_{x}$” can be written as a $\Sigma_{1}^{\mathcal{P}}$-formula, it is absolute between models $\mathcal{M}$ and $\mathcal{N}$ such that $\mathcal{M}\subseteq_{\mathrm{end}}^{\mathcal{P}}\mathcal{N}$. Thus for models $\mathcal{M}$ and $\mathcal{N}$ of $\mathsf{KPR}$:
    $\mathcal{M}\subseteq_{\mathrm{end}}^{\mathcal{P}}\mathcal{N}\Rightarrow\mathcal{M}\subseteq_{\mathrm{rank}}\mathcal{N}.$
    As pointed out by the referee, by [M-1, Proposition 6.17], the converse of the above implication also holds provided $\mathcal{M}\models\mathsf{Most}$ and $\mathcal{N}\models\mathsf{Mac}$.
  6. (f)
    If $\mathcal{M}\preceq_{\Delta_{0}}\mathcal{N}$, and $\mathcal{M}^{\ast}$ is the convex hull of $\mathcal{M}$ in $\mathcal{N}$, then $\mathcal{M}\subseteq_{\mathrm{cof}}\mathcal{M}^{\ast}\subseteq_{\mathrm{end}}\mathcal{N}$. If $\mathcal{M}$ and $\mathcal{N}$ are furthermore assumed to be models of $\mathsf{KPR}$, and $\mathcal{M}\preceq_{\Delta_{0}^{\mathcal{P}}}\mathcal{N}$, then (e) above implies that $\mathcal{M}\subseteq_{\mathrm{cof}}\mathcal{M}^{\ast}\subseteq_{\mathrm{rank}}\mathcal{N}$.
  7. (g)
    If both $\mathcal{M}$ and $\mathcal{N}$ are models of $\mathsf{KPR}$, and $\mathcal{M}\preceq_{\Delta_{0}}\mathcal{N}$, then the statement $\mathcal{N}$ is taller than $\mathcal{M}$” is equivalent to “there is some $\gamma\in\mathrm{Ord}^{\mathcal{N}}$ that exceeds each $\alpha\in\mathrm{Ord}^{\mathcal{M}}$”.
  8. (h)
    Generic extensions of models of $\mathsf{ZF}$ are end extensions but not rank extensions since they have the same ordinals. However, by (f) above, powerset preserving end extensions of models of $\mathsf{KPR}$ are rank extensions, thus for models $\mathcal{M}$ and $\mathcal{N}$ of $\mathsf{KPR}$ we have:
[…]
    On the other hand, since $\mathcal{P}(x)=y$ is $\Pi_{1}$-statement, we have:
    $\mathcal{M}\preceq_{\Sigma_{1},\mathrm{end}}\mathcal{N}\Longrightarrow\mathcal{M}\subseteq_{\mathrm{end}}^{\mathcal{P}}\mathcal{N}.$
    The converse of the above implication need not be true for arbitrary models $\mathcal{M}$ and $\mathcal{N}$ of $\mathsf{KPR}$. The referee has offered the following counterexample: let $\mathcal{N}$ be a model of $\mathsf{ZFC}$ that is $\omega$-standard but whose $\omega_{1}$ is nonstandard, and let $\mathcal{M}$ be the well-founded part of $\mathcal{N}$. Then $\mathcal{M}\models\mathsf{KPR}$, and the property of being equinumerous to a von Neumann ordinal is a $\Sigma_{1}$-property of $\mathcal{P}^{\mathcal{N}}(\omega)$ that holds in $\mathcal{N}$ but fails in $\mathcal{M}$. However, if $\mathcal{M}\models\mathsf{Most}$ and $\mathcal{N}\models\mathsf{Mac}$, then by [M-1, Proposition 6.17] the converse of the above implication holds.
  9. (i)
    As we shall see in Lemma 4.3, faithful end extensions of models of $\mathsf{ZF}$ are rank extensions.
  10. (g)
    Thanks to the provability of the induction scheme (over natural numbers) in $\mathsf{ZF}$, if $\mathcal{N}$ is a faithful extension of a model $\mathcal{N}$ (where $\mathcal{M}$ is a model of $\mathsf{ZF}$), then $\mathcal{N}$ fixes each element of $\omega^{\mathcal{M}}$ (but $\mathcal{N}$ need not fix $\omega^{\mathcal{M}}$, e.g., consider the case when $\mathcal{N}$ is an internal ultrapower of $\mathcal{M}$ modulo a nonprincipal ultrafilter over $\omega^{\mathcal{M}}$).
2.7. Definition. Reasoning within $\mathsf{KP}$, for each object $a$ in the universe of sets, let $\dot{a}$ be a constant symbol denoting $a$ (where the map $a\mapsto\dot{a}$ is $\Delta_{1}$, e.g., $\dot{a}=\left\langle 3,a\right\rangle$ as in Devlin’s monograph [D]). Let $\mathrm{Sent}_{\mathcal{L}_{\mathrm{set}}}(x)$ be the $\mathcal{L}_{\mathrm{set}}$-formula that defines the proper class $\mathrm{Sent}_{\mathcal{L}_{\mathrm{set}}}$ of sentences in the language
$\mathcal{L}^{+}=$ $\mathcal{L}\cup\{\dot{a}:a\in\mathrm{V}\}$,
[…]
  1. (a)
    Given $k\in\omega$, $S$ is a $\Sigma_{k}$-satisfaction class for $\mathcal{M}$ if $\left(\mathcal{M},S\right)\models\mathrm{Sat}(k,\mathrm{S})$, where $\mathrm{Sat}(k,\mathrm{S})$ is the universal generalization of the conjunction of the axioms $(i)$ through $(iv)$ below. We assume that first order logic is formulated using only the logical constants $\left\{\lnot,\vee,\exists\right\}.$
    1. $(i)$
      $\left[\left(\mathrm{S}\left(\dot{x}=\dot{y}\right)\leftrightarrow x=y\right)\wedge\left(\mathrm{S}\left(\dot{x}\in\dot{y}\right)\leftrightarrow x\in y\right)\right].$
    2. $(ii)$
      $\left[\mathrm{Sent}_{\mathcal{L}_{\mathrm{set}}}(k,\varphi)\wedge\left(\varphi=\lnot\psi\right)\right]\rightarrow\left[\mathrm{S}(\varphi)\leftrightarrow\lnot\mathrm{S}\mathsf{(}\psi\mathsf{)}\right]\mathsf{.}$
    3. $(iii)$
      $\left[\mathrm{Sent}_{\mathcal{L}_{\mathrm{set}}}(k,\varphi)\wedge\left(\varphi=\psi_{1}\vee\psi_{2}\right)\right]\rightarrow\left[\mathrm{S}(\varphi)\leftrightarrow\left(\mathrm{S}(\psi_{1})\vee\mathrm{S}(\psi_{2})\right)\right]\mathsf{.}$
    4. $(iv)$
      $\left[\mathrm{Sent}_{\mathcal{L}_{\mathrm{set}}}(k,\varphi)\wedge\left(\varphi=\exists v\ \psi(v)\right)\right]\rightarrow\left[\mathrm{S}(\varphi)\leftrightarrow\exists x\ \mathrm{S}\mathsf{(}\psi\mathsf{(}\dot{x}\mathsf{))}\right].$
  2. (b)
    $S$ is a full satisfaction class for $\mathcal{M}$ if $\left(\mathcal{M},S\right)\models\mathrm{Sat}(k,\mathrm{S})$ for every $k\in\omega$.
2.8. Tarski’s Definability and Undefinability of Truth TheoremsSuppose $\mathcal{M}$ is an $\mathcal{L}_{\mathrm{set}}$-structure.
  1. (a)
    (Definability/Codability of Truth) If $\mathcal{M}$ is a structure coded as an element $m$ of a model $\mathcal{N}$ of a sufficiently strong [^12] There are two canonical fragments of $\mathsf{ZF}$ that are ‘sufficiently strong’ for this purpose, namely: (1) $\mathsf{KP}$ (see Definition 2.1(d)), as shown by Friedman, Lu, and Wong in [FLW, Lemma 4.1]. (2) The fragment $\mathsf{M}_{0}$ of $\mathsf{Z}$ (see Definition 2.1(e)), as shown by Mathias [M-1, Proposition 3.10]. Note, however, that much weaker systems suffice if one only wishes to have a set theory within which the Tarskian satisfaction relation of every internal set structure is definable, as opposed to: coded as a set. One such weak system is $\mathsf{DS}$ (for ‘Devlin strengthened’), which is shown by Mathias [M-2, Proposition 10.37] to be capable of defining the Tarskian satisfaction predicate for set structures. fragment of $\mathsf{ZF}$, then there is $s\in N$ such that $\mathcal{N}$ views $s$ as the Tarskian satisfaction relation on $m$, written $\mathcal{N}\models\mathrm{sat}(s,m).$ [^13] In other words, within $\mathcal{N}$, $s$ is the set consisting of ordered pairs $(\varphi(\vec{x}),\vec{a})$ such that $m\models\varphi\left(\vec{x}/\vec{a}\right).$ Moreover, $\mathrm{sat}(x,y)$ is $\Delta_{1}$ in $\mathcal{N}.$
  2. (b)
    (Undefinability of Truth) If $S$ is a full satisfaction class for $\mathcal{M}$, then $S$ is not $\mathcal{M}$-definable.
It is a well-known result of Levy that if $\mathcal{M}\models\mathsf{ZF}\mathrm{,}$ then there is a $\Delta_{0}$-satisfaction class for $\mathcal{M}$ that is definable in $\mathcal{M}$ both by a $\Sigma_{1}$-formula and a $\Pi_{1}$-formula (see [J, p. 186] for a proof). This makes it clear that for each $n\geq 1,$ there is a $\Sigma_{n}$-satisfaction class for $\mathcal{M}$ that is definable in $\mathcal{M}$ by a $\Sigma_{n}$-formula. Levy’s result extends to models of $\mathsf{ZF}\mathrm{(}\mathcal{L}\mathrm{)}$ if $\mathcal{L}$ is finite as follows:
[…]
The Reflection Theorem is often formulated as a theorem scheme of $\mathsf{ZF}$ (e.g., as in [J]), the proof strategy of the Reflection Theorem applies equally well to $\mathsf{ZF}(\mathcal{L})$ for finite extensions $\mathcal{L}$ of $\mathcal{L}_{\mathrm{set}}$.
2.10. Montague-Vaught-Levy Reflection Theorem. Let $\mathcal{L}$ be a finite extension of $\mathcal{L}_{\mathrm{set}}.$ For each $n\in\omega$, $\mathsf{ZF}(\mathcal{L})$ proves that there are arbitrarily large ordinals $\gamma$ such that the submodel of the universe determined by VVγ is a $\Sigma_{n}(\mathcal{L})$-elementary submodel of the universe. [^15] This is expressible with the help of $\mathrm{Sat}_{\Sigma_{n}(\mathcal{L})}.$
2.11. $\mathbf{\Sigma}_{\mathbf{n}}$-Elementary Chains Theorem. Suppose $n\in\omega,$ $\mathcal{L}\supseteq\mathcal{L}_{\mathrm{set}}$, $(I,<_{I})$ is a linear order, and $\left\{\mathcal{M}_{i}:i\in I\right\}$ is a collection of $\mathcal{L}$-structures such that $\mathcal{M}_{i}\ \mathcal{\preceq}_{\Sigma_{n}(\mathcal{L})}\ \mathcal{M}_{i^{\prime}}$ whenever $i<_{I}i^{\prime}.$ Then we have:
$\mathcal{M}_{i}\ \mathcal{\preceq}_{\Sigma_{n}(\mathcal{L})}\ \bigcup\limits_{i\in I}\mathcal{M}_{i}.$
The following is a special case of a result of Gaifman [G-1, Theorem 1, p.54]. In the statement and the proof, $\mathcal{M}^{\ast}\prec_{{}_{\Delta_{0}(\mathcal{L}),\mathrm{end}}}\mathcal{N}$ is shorthand for the conjunction of $\mathcal{M}^{\ast}\prec_{\Delta_{0}(\mathcal{L)}}\mathcal{N}$ and $\mathcal{M}^{\ast}\prec_{\mathrm{end}}\mathcal{N}$.
2.12. Gaifman Splitting Theorem. Let $\mathcal{L}\supseteq\mathcal{L}_{\mathrm{set}}.$ Suppose $\mathcal{M}\models\mathsf{ZF}(\mathcal{L})$, $\mathcal{N}$ is an $\mathcal{L}$-structure with $\mathcal{M}\prec_{\Delta_{0}(\mathcal{L})}\mathcal{N}$, and $\mathcal{M}^{\ast}$ is the convex hull of $\mathcal{M}$ in $\mathcal{N}$. Then the following hold:
  1. (a)
    $\mathcal{M}\preceq_{\mathrm{cof}}\mathcal{M}^{\ast}\preceq_{\Delta_{0}(\mathcal{L}),\mathrm{end}}\mathcal{N}$.
  2. (b)
    $\mathcal{M\prec N}\Rightarrow\mathcal{M}\preceq_{\mathrm{cof}}\mathcal{M}^{\ast}\preceq_{\mathrm{end}}\mathcal{N}$.
Proof. To simplify the notation, we present the proof for $\mathcal{L}=\mathcal{L}_{\mathrm{set}}$ since the same reasoning handles the general case. In our proof we use the notation $\varphi^{v}$, where $\varphi$ is an $\mathcal{L}_{\mathrm{set}}$-formula and $v$ is a parameter, to refer to the $\Delta_{0}$-formula obtained by restricting all the quantifiers of $\varphi$ to the elements of $v$. A straightforward induction on the complexity of formulae shows that:
(1) If $\mathcal{K}$ is an $\mathcal{L}_{\mathrm{set}}$-structure, and $m\in K$, then $\mathcal{K}\models\varphi^{m}$ iff $(m,\in)^{\mathcal{K}}\models\varphi$.
Note that in the above, $\varphi$ is allowed to contain parameters in $\mathrm{Ext}_{\mathcal{K}}(m).$ Suppose $\mathcal{M}\prec_{\Delta_{0}}\mathcal{N},$ where $\mathcal{M}\models\mathsf{ZF.}$ Note that we are not assuming that $\mathcal{N}\models\mathsf{ZF}$. Putting (1) together with $\mathcal{M}\prec_{\Delta_{0}}\mathcal{N}$ implies that $(m,\in)^{\mathcal{M}}\preceq(m,\in)^{\mathcal{M}^{\ast}}$ for all $m\in M.$ On the other hand, the definition of $\mathcal{M}^{\ast}$ makes it clear that $(m,\in)^{\mathcal{N}}=(m,\in)^{\mathcal{M}^{\ast}}$ if $m\in M.$ Putting all this together, we have:
(2) For all $m\in M$, $(m,\in)^{\mathcal{M}}\preceq(m,\in)^{\mathcal{M}^{\ast}}=(m,\in)^{\mathcal{N}}$.
It is easy to see, using the definition of $\mathcal{M}^{\ast}$, that $\mathcal{M}\subseteq_{\mathrm{cof}}\mathcal{M}^{\ast}\subseteq_{\mathrm{end}}\mathcal{N}$, and in particular $\mathcal{M}^{\ast}\preceq_{\Delta_{0}}\mathcal{N}$ by Remark 2.6(b). Thus the proof of (a) is complete once we show that $\mathcal{M}\preceq_{\Sigma_{n}}\mathcal{M}^{\ast}$ for each $n\in\omega.$ Towards this aim, given $n\in\omega$, we can invoke the Reflection Theorem in $\mathcal{M}$ to get hold of some $U\subseteq\mathrm{Ord}^{\mathcal{M}}$ such that:
(3) $(\mathcal{M},U)$ satisfies “$U$ is unbounded in $\mathrm{Ord},$ and $\left[\left(\mathrm{V}_{\alpha},\in\right)\prec_{\Sigma_{n}}\left(\mathrm{V},\in\right)\right]$ for every $\alpha\in U$.
For each $\alpha\in U$ let $m_{\alpha}\in M$ such that $m_{\alpha}=\mathrm{V}_{\alpha}^{\mathcal{M}}.$ By (3), we have:
(4) If $\alpha,\beta\in U$ with $\alpha<\beta$, then $(m_{\alpha},\in)^{\mathcal{M}}\prec_{\Sigma_{n}}(m_{\beta},\in)^{\mathcal{M}}.$
Next, we observe that if $a,b$ are in $M$ with $(a,\in)^{\mathcal{M}}\prec_{\Sigma_{n}}(b,\in)^{\mathcal{M}}$, then for each $k$-ary $\Sigma_{n}$-formula $\sigma(\vec{x})$, $\mathcal{M}$ satisfies the $\Delta_{0}$-statement $\delta_{\sigma}(a,b),$ where:
$\delta_{\sigma}(a,b):=\forall\vec{x}\in a\ \left[\sigma^{a}(\vec{x})\leftrightarrow\sigma^{b}(\vec{x})\right]$,
where the prefix $\forall\vec{x}\in a$ is shorthand for $\forall x_{0}\in a\cdot\cdot\cdot\forall x_{k-1}\in a.$ Therefore, by putting (4) and (2) together with the assumption $\mathcal{M}\prec_{\Delta_{0}}\mathcal{N}$ we can conclude:
(5) If $\alpha,\beta\in U$ with $\alpha<\beta$, then $(m_{\alpha},\in)^{\mathcal{M}^{\ast}}\prec_{\Sigma_{n}}(m_{\beta},\in)^{\mathcal{M}^{\ast}}.$
On the other hand, since $U$ is unbounded in $\mathrm{Ord}^{\mathcal{M}}$, the definition of $m_{\alpha}$, together with the fact that $\mathcal{M}\subseteq_{\mathrm{cof}}\mathcal{M}^{\ast}$, imply:
[…]
Thanks to (2), (5), (6), and Theorem 2.11 ($\Sigma_{n}$-Elementary Chains Theorem) we can conclude:
(7) If $\alpha\in U$, then $(m_{\alpha},\in)^{\mathcal{M}}\preceq(m_{\alpha},\in)^{\mathcal{M}^{\ast}}\preceq_{\Sigma_{n}}\mathcal{M}^{\ast}\mathrm{.}$
We can conclude that $\mathcal{M}\preceq_{\Sigma_{n}}\mathcal{M}^{\ast}$ by (4), (7), and Theorem 2.11. This concludes the proof of (a).
To prove (b), suppose $\mathcal{M}\prec\mathcal{N}.$ With (a) at hand, we only need to show that $\mathcal{M}^{\ast}\preceq_{\Sigma_{n}}\mathcal{N}$ for each $n\in\omega$. We observe that if $\alpha\in U,$ then by (5) $(m_{\alpha},\in)^{\mathcal{M}^{\ast}}\prec_{\Sigma_{n}}\mathcal{M}^{\ast},$ and by (2), $(m_{\alpha},\in)^{\mathcal{M}^{\ast}}=(m_{\alpha},\in)^{\mathcal{N}}.$ On the other hand, the assumption $\mathcal{M}\prec\mathcal{N}$ together with (3) makes it clear that if $\alpha\in U,$ $(m_{\alpha},\in)^{\mathcal{N}}\prec_{\Sigma_{n}}\mathcal{N}$. Thus $\mathcal{M}^{\ast}\preceq_{\Sigma_{n}}\mathcal{N}$ by Theorem 2.11. $\square$
2.13. Remark. In general, if $\mathcal{M}\subseteq\mathcal{N}$, where $\mathcal{M}$ and $\mathcal{N}$ are both models of $\mathsf{ZF}$, the condition $\mathrm{Ord}^{\mathcal{M}}$ is cofinal in $\mathrm{Ord}^{\mathcal{N}}$” does not imply that $\mathcal{M}$ is cofinal in $\mathcal{N}$ (for example consider $\mathcal{M}$ and $\mathcal{N}$ where $\mathcal{N}$ is a set-generic extension of $\mathcal{M}$). However, the conjunction of $\Sigma_{1}$-elementarity and $\mathrm{Ord}^{\mathcal{M}}$ is cofinal in $\mathrm{Ord}^{\mathcal{N}}$”, where $\mathcal{M}\models\mathsf{ZF}$ (or even $\mathsf{KPR}$) and $\mathcal{N}$ is an $\mathcal{L}_{\mathrm{set}}$-structure implies that $\mathcal{M}$ is cofinal in $\mathcal{N}$, since the statement $(y\in\mathrm{Ord}\wedge x=\mathrm{V}_{y})$ can be written as a $\Pi_{1}$-statement within $\mathsf{KPR}$, and therefore it is absolute between $\mathcal{M}$ and $\mathcal{N}$ if $\mathcal{M}\prec_{\Sigma_{1}}\mathcal{N}.$ [^16] As noted in [Kan, Lemma 0.2, p5], the rank function is $\Delta_{1}^{\mathsf{ZF}}$, and this fact can be used to show that $(y\in\mathrm{Ord}\wedge x=\mathrm{V}_{y})$ can be written as a $\Pi_{1}$-statement within $\mathsf{ZF}$; an examination of the proof makes it clear that $\mathsf{KPR}$ suffices for this purpose. This is how elementary embeddings between inner models are usually formalized within $\mathsf{ZF}$, as in Kanamori’s monograph [Kan, Proposition 5.1].
Recall that for models of $\mathsf{ZF}$, the $\mathcal{L}_{\mathrm{set}}$-sentence $\exists p\left(\mathrm{V}=\mathrm{HOD}(p)\right)$ expresses: “there is some $p$ such that every set is first order definable in some structure of the form $\left(\mathrm{V}_{\alpha},\in,p,\beta\right)_{\beta<\alpha}$ with $p\in\mathrm{V}_{\alpha}$”. The following result is classical.
2.14. Myhill-Scott Theorem [MS]The following statements are equivalent for $\mathcal{M}\models\mathsf{ZF}$:
  1. (a)
    $\mathcal{M}\models\exists p\left(\mathrm{V}=\mathrm{HOD}(p)\right).$
  2. (b)
    $\mathcal{M}$ carries a definable global well-ordering, i.e., for some $p\in M$ and some set-theoretic formula $\varphi(x,y,z)$, $\mathcal{M}$ satisfies $\varphi(x,y,p)$ well-orders the universe”.
  3. (c)
    $\mathcal{M}$ carries a definable global choice function, i.e., for some $p\in M$ and some set-theoretic formula $\psi(x,y,z)$, $\mathcal{M}$ satisfies $\psi(x,y,p)$ is the graph of a global choice function”.
3. TALLER CONSERVATIVE ELEMENTARY EXTENSIONS
3.1. Theorem. Every model $\mathcal{M}$ of $\mathsf{ZF}+\exists p\left(\mathrm{V}=\mathrm{HOD}(p)\right)$ (of any cardinality) has a conservative elementary extension $\mathcal{N}$ such that $\mathcal{N}$ is taller than $\mathcal{M}$.
We shall present two proofs of Theorem 3.1, the first one is based on a class-sized syntactic construction taking place within a model of set theory; it was inspired by Kaufmann’s proof of the MacDowell-Specker theorem using the Arithmetized Completeness Theorem, as described by Schmerl [Schm-1]. [^17] As noted by Phillips [P] and Gaifman [G-2], the McDowell-Specker proof lends itself to a fine-tuning that ensures that the conservative elementary extension $\mathcal{N}$ of a given model $\mathcal{M}$ has the minimality property, i.e., there is no $\mathcal{K}$ such that $\mathcal{M}\prec\mathcal{K}\prec\mathcal{N}$. However, Kaufmann’s slick proof does not seem to lend itself to this embellishment. The second one is based on a model-theoretic construction that is reminiscent of the original ultrapower proof [^18] The original proof by MacDowell and Specker of their result uses the concept of a “finitely additive 2-valued measure”, which is an alternative formulation of the notion of an ultrafilter. of the MacDowell-Specker theorem [MS]. The first proof is short and devilish; the second proof is a bit longer but is more transparent due to its combinatorial flavor; it also lends itself to a refinement, as indicated in Remark 3.3 and Theorem 3.4.
First proof of Theorem 3.1. We need the following two Facts 1 and 2 below.
FACT 1: Suppose $\mathcal{M}$ is a model of $\mathsf{ZF}$ that carries an $\mathcal{M}$-definable global well-ordering, and $T$ is an $\mathcal{M}$-definable class of first order sentences such that $\mathcal{M}$ satisfiesTT is a consistent first order theory”. Then there is a model $\mathcal{N}\models T^{\mathrm{st}}$ such that the elementary diagram of $\mathcal{N}$ is $\mathcal{M}$-definable, where $T^{\mathrm{st}}$ is the collection of sentences in $T$ with standard shape [^19] More specifically, $\varphi\in T^{\mathrm{st}}$ iff $\varphi\in T$ and there is some standard formula $\varphi^{\ast}$ such that $\mathcal{M}$ believes that $\varphi$ is the result of substituting constants for the free variables of $\varphi^{\ast}.$.
Proof of Fact 1. Since $\mathcal{M}$ has a definable global well-ordering, the Henkin proof of the completeness theorem of first order logic can be applied within $\mathcal{M}$ to construct a Henkinized complete extension $T^{\mathrm{Henkin}}$ of $T$ (in a language extending the language of $T$ by class-many new constant symbols) such that $T$ is definable in $\mathcal{M}$. This in turn allows $\mathcal{M}$ to define $\mathcal{N}$ by reading it off $T^{\mathrm{Henkin}},$ as in the usual Henkin proof of the completeness theorem.
[…]
Proof of Fact 2. This is an immediate consequence of the Reflection Theorem 2.10.
Starting with a model $\mathcal{M}$ of $\mathsf{ZF}+\exists p\left(\mathrm{V}=\mathrm{HOD}(p)\right),$ by Theorem 2.14, $\mathcal{M}$ carries a global definable well-ordering. We will construct an increasing sequence of $\mathcal{L}_{\mathrm{set}}$-structures $\left\langle\mathcal{N}_{k}:k\in\omega\right\rangle$ that satisfies the following properties for each $k\in\omega$:
  1. (1)
    $\mathcal{N}_{0}=\mathcal{M},$
  2. (2)
    $\mathcal{M}\preceq_{\Pi_{k+2}}\mathcal{N}_{k}\preceq_{\Pi_{k+1}}\mathcal{N}_{k+1}$.
  3. (3)
    There is some $\alpha\in\mathrm{Ord}^{\mathcal{N}_{1}}$ that is above each $\beta\in\mathrm{Ord}^{\mathcal{M}}$, thus $\mathcal{M}\preceq_{\mathrm{taller}}\mathcal{N}_{k}$ for all $k\geq 1$.
  4. (4)
    $\mathcal{N}_{k}$ is a self-extension of $\mathcal{M}$, and in particular $\mathcal{M}\preceq_{\mathrm{cons}}\mathcal{N}_{k}$. [^20] See Definition 2.4(b) and Remark 2.5(a).
In other words:
$\mathcal{M=N}_{0}\preceq_{\Pi_{3},\mathrm{taller}}\mathcal{N}_{1}\preceq_{\Pi_{2}}\mathcal{N}_{2}\preceq_{\Pi_{3}}\mathcal{N}_{3}\$...,
and for each $k\in\omega$, $\mathcal{M}\preceq_{\Pi_{k+2},\mathrm{cons,taller}}\mathcal{N}_{k}.$
In light of part (a) of Remark 2.5, this shows that $\mathcal{M}\prec_{\mathrm{cons,taller}}\mathcal{N}$, where $\mathcal{N}:=\bigcup\limits_{k\in\omega}\mathcal{N}_{k}$. So the proof will be complete once we explain how to recursively build the desired chain $\left\langle\mathcal{N}_{k}:k\in\omega\right\rangle.$
Using the notation of Definition 2.7, let $\mathcal{L}_{\mathrm{set}}^{+}=$ $\mathcal{L}_{\mathrm{set}}\cup\{\dot{a}:a\in\mathrm{V}\}$ be the language $\mathcal{L}_{\mathrm{set}}$ of set theory with a constant $\dot{a}$ for each $a\in\mathrm{V}$. To build $\mathcal{N}_{1}$ we argue within $\mathcal{M}$: add a new constant $c$ to $\mathcal{L}_{\mathrm{set}}^{+}$ and consider the theory $T_{1}$ defined as follows:
$T_{1}=\left\{\mathrm{Ord}(c)\right\}\cup\{\dot{\alpha}\in c:\mathrm{Ord}(\alpha)\}\cup\mathrm{Th}_{\Pi_{3}}($$V.,\in,\dot{a})_{a\in\mathrm{V}}$.
Thus $T_{1}$ is a proper class within $\mathcal{M}$. Arguing within $\mathcal{M}$, and using Fact 2, it is easy to see that $\mathrm{Con}(T_{1})$ holds, and therefore by Fact 1 we can get hold of a proper class model $\mathcal{N}_{1}$ of $T_{1}$ such that $\mathcal{N}_{1}$ is a self-extension of $\mathcal{M}$. Thus $\mathcal{M}\prec_{{}_{\Pi_{3,\mathrm{cons}}}}\mathcal{N}_{1}$, and there is an ordinal in $\mathcal{N}_{1}$ that is above all of the ordinals of $\mathcal{M}$.
Next suppose we have built $\left\langle\mathcal{N}_{i}:1\leq i\leq k\right\rangle$ for some $k\in\omega$ while complying with (1) through (4). Consider the theory $T_{k}$ defined in $\mathcal{M}$ as follows:
$T_{k+1}=\mathrm{Th}_{\Pi_{k+1}}(\mathcal{N}_{k},b)_{b\in N_{k}}\cup\mathrm{Th}_{\Pi_{k+2}}($$V.,\in,a)_{a\in\mathrm{V}}$.
Fact 2 together with the inductive assumption that $\mathcal{M}\prec_{\Pi_{k+1}}\mathcal{N}_{k}$ makes it evident that $\mathrm{Con}(T_{k+1})$ holds in $\mathcal{M}$. Hence by Fact 1 we can get hold of a model $\mathcal{N}_{k+1}$ of $T$ such that $\mathcal{N}_{k+1}$ is a self-extension of $\mathcal{M}$. This makes it clear that $\left\langle\mathcal{N}_{i}:1\leq i\leq k+1\right\rangle$ meets (1) through (4). $\square$
Second proof of Theorem 3.1. Given $\mathcal{M}\models\mathsf{ZF}+\exists p\left(\mathrm{V}=\mathrm{HOD}(p)\right),$ let $\mathbb{B}$ be the Boolean algebra consisting of $\mathcal{M}$-definable subsets of $\mathrm{Ord}^{\mathcal{M}}$ (ordered by $\subseteq$). In Stage 1 of the proof we will build an appropriate ultrafilter $\mathcal{U}$ on $\mathbb{B},$ and in Stage 2 we will verify that the definable ultrapower of $\mathcal{M}$ modulo $\mathcal{U}$, denoted $\mathcal{M}^{\mathrm{Ord}^{\mathcal{M}}}/\ \mathcal{U}$, is a conservative elementary extension of $\mathcal{M}$ that is taller than $\mathcal{M}$.
STAGE 1: The following facts come handy in this stage:
Fact 1. By Theorem 2.9, for each $n\in\omega$ there is a definable satisfaction predicate for $\Sigma_{n}$-formulae.
Fact 2. By Theorem 2.14, $\mathcal{M}$ has a parametrically definable global well-ordering.
To facilitate the construction of $\mathcal{U}$, let us introduce some conventions:
  1. $(i)$
    Given $\mathfrak{X}\subseteq\mathcal{P}(M)$, we say that $\mathfrak{X}$ is $\mathcal{M}$-listable, if that there is a parametric formula $\psi(\alpha,\beta)$ such that for all $X\subseteq M$ we have:
$X\in\mathfrak{X}$ iff $\exists\alpha\in\mathrm{Ord}^{\mathcal{M}}$ $X=\{m\in M:\mathcal{M}\models\psi(\alpha,m)\}.$
  1. In the above context, if $\psi$ is a $\Sigma_{n}$-formula, $\mathfrak{X}$ is said to be $\Sigma_{n}$-listable in $\mathcal{M}$. Moreover, we say that $\left\langle X_{\alpha}:\alpha\in\mathrm{Ord}^{\mathcal{M}}\right\rangle$ is an $\mathcal{M}$-list of $\mathfrak{X}$ if $X_{\alpha}=\{m\in M:\mathcal{M}\models\psi(\alpha,m)\}$ for all $\alpha\in\mathrm{Ord}^{\mathcal{M}}.$
  2. $(ii)$
    Given an $\mathcal{M}$-listable $\mathfrak{X}\subseteq\mathbb{B}$, we say that $\mathfrak{X}$ is $\mathcal{M}$-thick if every $\mathcal{M}$-finite intersection of elements of $\mathfrak{X}$ is nonempty. More precisely, $\mathfrak{X}$ is $\mathcal{M}$-thick if $\mathcal{M}$ satisfies the following, where $\psi$ is a listing formula for $\mathfrak{X}$:
$\forall k\in\omega$ $\forall\left\langle\alpha_{i}:i<k\right\rangle\in\mathrm{Ord}^{k}\ \exists\gamma\in\mathrm{Ord}\ \forall i<k\ \psi(\alpha_{i},\gamma).$
  1. $(iii)$
    For each $n\in\omega$, let $\mathbb{B}_{n}$ be the collection of $X\in\mathbb{B}$ such that $X$ is $\Sigma_{n}$-definable in $\mathcal{M}$ (parameters allowed). Note that by Facts 1 and 2, $\mathbb{B}_{n}$ is $\mathcal{M}$-listable for each $n\in\omega$.
  2. $(iv)$
[…]
We wish to construct an ultrafilter $\mathcal{U}$ $\subseteq$ $\mathbb{B}$ that satisfies the following properties:
  1. $(1)$
    $\mathcal{U}$ is an ultrafilter on $\mathbb{B}$.
  2. $(2)$
    $\mathcal{F\subseteq U}$.
  3. $(3)$
    Given any $\mathcal{M}$-definable $f:\mathrm{Ord}^{\mathcal{M}}\rightarrow\mathcal{M}$, there is some $P\in\mathcal{U}$ such that either $f\upharpoonright P$ is one-to-one, or the range of $f\upharpoonright P$ is a set (as opposed to a class).
  4. $(4)$
    Given any $n\in\mathbb{\omega}$, $\mathcal{U}\cap\mathbb{B}_{n}$ is $\mathcal{M}$-listable. This condition is equivalent to: for any $\mathcal{M}$-list $\left\langle X_{\alpha}:\alpha\in\mathrm{Ord}^{\mathcal{M}}\right\rangle$, $\left\{\alpha\in\mathrm{Ord}^{\mathcal{M}}:X_{\alpha}\in\mathcal{U}\right\}\in\mathbb{B}.$
The desired ultrafilter $\mathcal{U}$ will be defined as $\bigcup\limits_{n\in\omega}\mathcal{U}_{n}$, where $\mathcal{U}_{n+1}$ will be defined from $\mathcal{U}_{n}$ using an internal recursion within $\mathcal{M}$ of length $\mathrm{Ord}^{\mathcal{M}}$. Thus, intuitively speaking, $\mathcal{U}$ will be constructed in $\omega\times\mathrm{Ord}^{\mathcal{M}}$-stages.
  • In what follows, for $n\in\omega$, we fix an $\mathcal{M}$-list $\left\langle f_{n,\alpha}:\alpha\in\mathrm{Ord}^{\mathcal{M}}\right\rangle$ of all parametrically definable $\Sigma_{n}$-functions [^21] The the notion of $\mathcal{M}$-listability can be naturally extended to family $\mathcal{Y}$ of subsets of $M^{2}$, since $\mathcal{Y}$ can be coded as a family $\mathfrak{X}_{\mathcal{Y}}$ of subsets of $M$ with the help of a canonical pairing function. of $\mathcal{M}$ with domain $\mathrm{Ord}^{\mathcal{M}}$, and an $\mathcal{M}$-list $\left\langle S_{n,\alpha}:\alpha\in\mathrm{Ord}^{\mathcal{M}}\right\rangle$ of all elements of $\mathbb{B}_{n}$. Such $\mathcal{M}$-lists exist by Facts 1 and 2 listed at the beginning of Stage 1.
Let $\mathcal{U}_{0}$ $\mathcal{F}$. Thus $\mathcal{U}_{0}$ is $\mathcal{M}$-listable and $\mathcal{M}$-thick. We now describe the inductive construction of $\mathcal{U}_{n+1}$ from $\mathcal{U}_{n}$. So suppose $n\in\omega$, and we have $\mathcal{U}_{n}\subseteq\mathbb{B}$ that satisfies the clauses $\mathbb{C}_{1}(n)$ through $\mathbb{C}_{3}(n)$ below. Note that conditions $\mathbb{C}_{2}(n)$ and $\mathbb{C}_{3}(n)$ only kick in for $n\geq 1$ and ensure that $\mathcal{U}_{n}$ abides by certain $\Sigma_{n-1}$-obligations for $n\geq 1$.
  1. $\mathbb{C}_{1}(n):$
    $\mathcal{U}_{n}$ is $\mathcal{M}$-listable and $\mathcal{M}$-thick.
  2. $\mathbb{C}_{2}(n):$
    If $n\geq 1$, then $\forall\alpha\in\mathrm{Ord}^{\mathcal{M}}\ \exists X\in\mathcal{U}_{n}$ ($f_{n-1,\alpha}\upharpoonright X$ is one-to-one, or the range of $f_{n-1,\alpha}\upharpoonright X$ is a set).
  3. $\mathbb{C}_{3}(n):$
[…]
The following lemma is crucial for the construction of $\mathcal{U}_{n+1}\supseteq\mathcal{U}_{n}$ such that $\mathcal{U}_{n+1}$ satisfies $\mathbb{C}_{1}(n+1)$ through $\mathbb{C}_{3}(n+1).$
Lemma 3.1.1. Suppose $n\in\omega$, $\mathcal{F}\subseteq\mathfrak{X}\subseteq\mathbb{B},$ $\mathfrak{X}$ is $\Sigma_{n}$-listable in $\mathcal{M}$, and $\mathfrak{X}$ is $\mathcal{M}$-thick. Then there is a large enough $k\in\omega$ that satisfies the following two conditions:
  1. (a)
    Given any $A\subseteq\mathrm{Ord}^{\mathcal{M}}$ that is $\Sigma_{m}$-definable in $\mathcal{M}$, let $\widehat{A}\in\mathbb{B}$ be defined by: $\widehat{A}=A$ if $\mathfrak{X\cup\{}A\}$ is $\mathcal{M}$-thick, and otherwise $\A\widehat{A}:=\mathrm{Ord}^{\mathcal{M}}\backslash$ A. Then $\mathfrak{X\cup\{}\widehat{A}\}$ is $\mathcal{M}$-thick, and $\widehat{A}$ is $\Sigma_{p}$-definable in $\mathcal{M}$ for $p=\max\{m,n\}+k$.
  2. (b)
    Given any $f:\mathrm{Ord}^{\mathcal{M}}\rightarrow M$ that is $\Sigma_{m}$-definable in $\mathcal{M}$, there is some $P\in\mathbb{B}$ that is $\Sigma_{p}$-definable in $\mathcal{M}$ for $p=\max\{m,n\}+k$ such that $\mathfrak{X\cup\{}P\}$ is $\mathcal{M}$-thick, and $f\upharpoonright$ $P$ is either one-to-one, or the range of $f\upharpoonright$ $P$ is coded as an element of $\mathcal{M}$.
Proof. (a) is easy to see, so we leave the proof to the reader. Our proof of (b) will not explicitly specify the bound $k$, instead we focus on the combinatorics, which is uniform enough to yield the existence of such a bound. Given an $\mathcal{M}$-definable $f:\mathrm{Ord}^{\mathcal{M}}\rightarrow M$, we distinguish two cases:
Case 1: There is some $\gamma\in\mathrm{Ord}^{\mathcal{M}}$ such that $\mathfrak{X}\cup\{Y_{\gamma}\}$ is $\mathcal{M}$-thick, where
$Y_{\gamma}=\{\alpha\in\mathrm{Ord}^{\mathcal{M}}:\mathcal{M}\models f(\alpha)<\gamma\}.$
Case 2: Not case 1.
If case 1 holds, then we let $P:=Y_{\gamma_{0}}$, where $\gamma_{0}$ is the first ordinal in $\mathrm{Ord}^{\mathcal{M}}$ witnessing the veracity of case 1. Clearly the range of $f\upharpoonright$ $P$ is coded as an element of $\mathcal{M}$.
If case 2 holds, then let $\mathfrak{X}^{\ast}:=$ the set of all $\mathcal{M}$-finite intersections of $\mathfrak{X}$. Note that $\mathfrak{X}^{\ast}$ is $\mathcal{M}$-listable, so we can let $\left\langle D_{\eta}:\eta\in\mathrm{Ord}^{\mathcal{M}}\right\rangle$ be an $\mathcal{M}$-list of $\mathfrak{X}^{\ast}$. It is also clear that the $\mathcal{M}$-thickness of $\mathfrak{X}$ is inherited by $\mathfrak{X}^{\ast}$. We will construct an $\mathcal{M}$-definable set $P=\{p_{\eta}:\eta\in\mathrm{Ord}^{\mathcal{M}}\}$ by recursion (within $\mathcal{M}$) such that the restriction of $f$ to $P$ is one-to-one, and such that $\mathfrak{X}\cup\{P\}$ is $\mathcal{M}$-thick. Suppose we have already built $P$ up to some $\gamma\in\mathrm{Ord}^{\mathcal{M}}$ as $\{p_{\eta}:\eta<\gamma\}$ such that the following two conditions hold in $\mathcal{M}$:
$(\ast)$ $p_{\eta}<p_{\eta^{\prime}}$ and $f(p_{\eta})<f(p_{\eta^{\prime}})$ whenever $\eta<\eta^{\prime}<\gamma$.
$(\ast\ast)$ $p_{\eta}\in$ $D_{\eta}$ for all $\eta<\gamma.$
Within $\mathcal{M}$, let $\theta=\sup\{f(p_{\eta}):\eta<\gamma\}.$ Note that $\theta$ is well-defined because $\mathcal{M}$ satisfies the replacement scheme. Since we know that case 2 holds, there must exist some $D_{\delta}\in\mathfrak{X}^{\ast}$ such that:
$D_{\delta}\cap Y_{\theta+1}=\varnothing$, where $Y_{\theta+1}:=\{\alpha\in\mathrm{Ord}^{\mathcal{M}}:\mathcal{M}\models f(\alpha)<\theta+1\}.$
Therefore within $\mathcal{M}$ can choose $p_{\gamma}$ to be the first member of $D_{\delta}\cap D_{\gamma\text{ }}$that is above $\{s_{\eta}:\eta<\gamma\}$. Since $\mathfrak{X}$ is $\mathcal{M}$-thick and $\mathcal{F}\subseteq\mathfrak{X}$, $D_{\delta}\cap D_{\gamma\text{ }}$is unbounded in $\mathrm{Ord}^{\mathcal{M}},$ hence $s_{\gamma}$ is well-defined. Thus we have constructed an $\mathcal{M}$-definable set $P=\{p_{\eta}:\eta\in\mathrm{Ord}^{\mathcal{M}}\}$ such that $p_{\eta}\in$ $D_{\eta}$ for all $\eta\in\mathrm{Ord}^{\mathcal{M}}$ (so $\mathfrak{X}\cup\{P\}$ is $\mathcal{M}$-thick), and $f$ is strictly monotone increasing on $P$. This concludes the proof of part (b) of Lemma 3.1.1. $\square$
Using Lemma 3.1.1, it is now straightforward to use transfinite recursion within $\mathcal{M}$ of length $\mathrm{Ord}^{\mathcal{M}}$ to construct a family $\mathcal{U}_{n+1}\subseteq\mathbb{B}$ that extends $\mathcal{U}_{n}$ and which satisfies clauses $\mathbb{C}_{1}(n+1)$ through $\mathbb{C}_{3}(n+1)$. $\mathcal{U}_{n+1}$ will be of the form:
$\mathcal{U}_{n}\ \cup\bigcup\limits_{\alpha\in\mathrm{Ord}^{\mathcal{M}}}\{A_{\alpha},P_{\alpha}\}$,
where $A_{\alpha}$ and $P_{\alpha}$ are defined within $\mathcal{M}$ by transfinite recursion on $\alpha$. Suppose for some $\alpha\in\mathrm{Ord}^{\mathcal{M}}$ we have built subsets $A_{\beta}$ and $P_{\beta}$ of $\mathbb{B}$ for each $\beta<\alpha,$ and $\mathcal{U}_{n}\cup\bigcup\limits_{\beta<\alpha}\{A_{\beta},P_{\beta}\}$ is $\mathcal{M}$-thick. Now we can use part (a) of Lemma 3.1.1 to construct $A_{\alpha}\in\mathbb{B}$ such that $\mathcal{U}_{n}\cup\bigcup\limits_{\beta<\alpha}\{A_{\beta},P_{\beta}\}\cup\{A_{\alpha}\}$ is $\mathcal{M}$-thick and $A_{\alpha}=S_{n,\alpha}$ or $A_{\alpha}=\mathrm{Ord}^{\mathcal{M}}\backslash S_{n,\alpha}.$ In the next step, using part (b) of Lemma 3.1.1, we can construct $P_{\alpha}\in\mathbb{B}$ which ‘takes care of’ $f_{n,\alpha}$ (i.e., the restriction of $f_{n,\alpha}$ to $P_{\alpha}$ is either one-to-one, or its range is a set) and which has the property that $\mathcal{U}_{n}\cup\bigcup\limits_{\beta<\alpha}\{A_{\beta},P_{\beta}\}\cup\{A_{\alpha},P_{\alpha}\}$ is $\mathcal{M}$-thick. This concludes the recursive construction of $A_{\alpha}$ and $P_{\alpha}$ for $\alpha\in\mathrm{Ord}^{\mathcal{M}}$, thus we can define:
$\mathcal{U}_{n+1}:=\mathcal{U}_{n}\ \cup\bigcup\limits_{\alpha\in\mathrm{Ord}^{\mathcal{M}}}\{A_{\alpha},P_{\alpha}\}$.
The construction makes it clear that $\mathcal{U}_{n+1}$ satisfies $\mathbb{C}_{1}(n+1)$ through $\mathbb{C}_{3}(n+1)$. This, in turn, makes evident that for
$\mathcal{U}:=\bigcup\limits_{n\in\omega}\mathcal{U}_{n}$,
$\mathcal{U}$ has properties (1) through (4) promised at the beginning of Stage 1. Note that $\mathcal{U}$ satisfies condition (4) since $\mathcal{U}_{n+1}$ and $\mathbb{B}_{n}$ are both $\mathcal{M}$-listable, and by clause $\mathbb{C}_{3}(n+1)$, $\mathcal{U}\cap\mathbb{B}_{n}\subseteq\mathcal{U}_{n+1}.$ This concludes Stage 1.
STAGE 2: Let $\mathcal{N}:=\mathcal{M}^{\mathrm{Ord}^{\mathcal{M}}}/\ \mathcal{U}$ be the definable ultrapower of $\mathcal{M}$ modulo $\mathcal{U}$ whose universe consists of $\mathcal{U}$-equivalence classes $f]_{\mathcal{U}}$ of $\mathcal{M}$-definable functions $f:\mathrm{Ord}^{\mathcal{M}}\rightarrow M$. Thanks to the availability of an $\mathcal{M}$-definable global well-ordering, the following Łoś-style result can be readily verified.
Theorem 3.1.2. For all $k$-ary formulae $\varphi(x_{0},\cdot\cdot\cdot,x_{k-1})$ of $\mathcal{L}_{\mathrm{set}}$ and all $k$-tuples of $\mathcal{M}$-definable functions $\left\langle f_{i}:i<k\right\rangle$ from $\mathrm{Ord}^{\mathcal{M}}$ to $M$, we have:
$\mathcal{N}\models\varphi([f_{0}]_{\mathcal{U}},\cdot\cdot\cdot,[f_{k-1}]_{\mathcal{U}})$ iff $\left\{\alpha\in\mathrm{Ord}^{\mathcal{M}}:\mathcal{M}\models\varphi\left(f_{0}(\alpha),\cdot\cdot\cdot,f_{k-1}(\alpha)\right)\right\}\in\mathcal{U}$.
For each $m\in M$ let $\widetilde{m}$: $\mathrm{Ord}^{\mathcal{M}}\rightarrow\{m\}$. Theorem 3.1.2 assures us that the map $j:\mathcal{M}\rightarrow\mathcal{N}$ given by $j(m)=[\widetilde{m}]_{\mathcal{U}}$ is an elementary embedding, thus by identifying $\mathcal{M}$ with the image of $j$, we can construe $\mathcal{N}$ as an elementary extension of $\mathcal{M}$. Let $\mathrm{id}:\mathrm{Ord}^{\mathcal{M}}\rightarrow\mathrm{Ord}^{\mathcal{M}}$ be the identity map. By property (2) of $\mathcal{U}$ specified in Stage 1, every element of $\mathcal{U}$ is unbounded in $\mathrm{Ord}^{\mathcal{M}}.$ So by Theorem 3.1.2, $\mathcal{N}$ is taller than $\mathcal{M}$ since $\mathrm{id}]_{\mathcal{U}}$ exceeds all ordinals of $\mathcal{M}$.
It remains to verify that $\mathcal{N}$ is a conservative extension of $\mathcal{M}$. Given a formula $\psi(x,y_{0},\cdot\cdot\cdot,y_{k-1})$ and parameters $s_{0},\cdot\cdot\cdot,s_{k-1}$ in $N$, let:
$Y=\left\{m\in M:\mathcal{N}\models\varphi([\widetilde{m}]_{\mathcal{U}},s_{0},\cdot\cdot\cdot,s_{k-1})\right\}.$
We wish to show that $Y$ is $\mathcal{M}$-definable. Choose $f_{0},\cdot\cdot\cdot,f_{k-1}$ so that $s_{i}=[f_{i}]_{\mathcal{U}}$ for all $i<p$ and consider the formula $\theta(x,\alpha)$ defined below:
$\theta(x,\alpha):=\alpha\in\mathrm{Ord}\wedge\varphi(x,f_{0}(\alpha),\cdot\cdot\cdot,f_{k-1}(\alpha))$.
Pick $n\in\omega$ large enough so that $\theta(x,\alpha)$ is a $\Sigma_{n}$-formula. For each $m\in M,$ let:
$S_{m}:=\{\alpha\in\mathrm{Ord}^{\mathcal{M}}:\mathcal{M}\models\theta(m,\alpha)\}$.
Note that the choice of $n$, $S_{m}\in\mathbb{B}_{n}$ for each $m\in M$, so $S_{m}\in\mathcal{U}$ iff $S_{m}\in\mathcal{U}\cap\mathbb{B}_{n}$ for each $m\in M$. On the other hand, by Theorem 3.1.2 we have:
$m\in Y$ iff $S_{m}\in\mathcal{U}$ for each $m\in M$.
So we can conclude that $m\in Y$ iff $S_{m}\in\mathcal{U}\cap\mathbb{B}_{n}$ for each $m\in M.$ Since $\mathcal{U}\cap\mathbb{B}_{n}$ is $\mathcal{M}$-listable by property (4) of $\mathcal{U}$ specified in Stage 1, we can write $\mathcal{U}\cap\mathbb{B}_{n}$ as an $\mathcal{M}$-list $\left\langle X_{\eta}:\eta\in\mathrm{Ord}^{\mathcal{M}}\right\rangle$ given by some formula $\psi(x,y).$ Thus the following holds for all $m\in M$, which shows that $Y$ is $\mathcal{M}$-definable.
$m\in Y$ iff $\exists\eta\left(S_{m}=X_{\eta}\right)$ iff $\mathcal{M}\models\exists\eta\ \forall\alpha\left(\theta(m,\alpha)\leftrightarrow\psi(\eta,m)\right).$
$\square$
3.2. Corollary. Work in $\mathsf{ZF}+\exists p\left(\mathrm{V}=\mathrm{HOD}(p)\right)$ in the metatheory, and let $\mathcal{M}$ be an arbitrary model of $\mathsf{ZF}+\exists p\left(\mathrm{V}=\mathrm{HOD}(p)\right).$
  1. (a)
    For every regular cardinal $\kappa$ there is a conservative elementary extension $\mathcal{N}_{\kappa}$ of $\mathcal{M}$ that is taller than $\mathcal{M}$ and $\mathrm{cf}(\mathrm{Ord}^{\mathcal{N}})=\kappa$. [^22] As indicated by the proof of this part, $\mathsf{ZFC}$ suffices as the metatheory for this part.
  2. (b)
    There is a proper class model $\mathcal{N}_{\mathrm{Ord}}$ whose elementary diagram is a definable class such that $(1)$ $\mathcal{M\prec}_{\mathrm{cons,taller}}\mathcal{N}_{\mathrm{Ord}}$, and $(2)$ $\mathrm{Ord}^{\mathcal{N}_{\mathrm{Ord}}}$ contains a cofinal class of order-type $\mathrm{Ord}$.
Proof. To establish (a), we observe that given a regular cardinal $\kappa$, Theorem 3.1 allows us to build an elementary chain of models $\left\langle\mathcal{M}_{\alpha}:\alpha\in\kappa\right\rangle$ with the following properties:
[…]
(2) For each $\alpha\in\kappa$ $\mathcal{M}_{\alpha+1}$ is a conservative elementary extension of $\mathcal{M}$ that is taller than $\mathcal{M}.$
(3) For limit $\alpha\in\kappa$, $\mathcal{M}_{\alpha}=\bigcup\limits_{\beta\in\alpha}\mathcal{M}_{\beta}.$
The desired model $\mathcal{N}$ is clearly $\mathcal{N}_{\kappa}=\bigcup\limits_{\alpha\in\kappa}\mathcal{M}_{\alpha}.$ This concludes the proof of part (a).
With part (a) at hand, it is evident that in the presence of a global definable well-ordering (available thanks to the veracity of $\exists p\left(\mathrm{V}=\mathrm{HOD}(p)\right)$ in the metatheory), we can construct the desired model $\mathcal{N}_{\mathrm{Ord}}$ as the union of an elementary chain of models $\left\langle\mathcal{M}_{\alpha}:\alpha\in\mathrm{Ord}\right\rangle$ that satisfies (1) and it also satisfies the result of replacing $\kappa$ with $\mathrm{Ord}$ in (2) and (3). $\square$
[…]
3.4. Theorem. Every model of $\mathsf{ZF}+\exists p\left(\mathrm{V}=\mathrm{HOD}(p)\right)$ has a cofinal conservative elementary extension that possesses a minimal elementary end extension.
Proof. Let $\mathcal{M}$ be a model of $\mathsf{ZF}+\exists p\left(\mathrm{V}=\mathrm{HOD}(p)\right)$, $\mathcal{N}$ be the model constructed in the second proof of Theorem 3.1, and let $\mathcal{M}^{\ast}$ be the convex hull of $\mathcal{M}$ in $\mathcal{N}$. Note that by Gaifman Splitting Theorem we have:
$\mathcal{M}\preceq_{\mathrm{cof}}\mathcal{M}^{\ast}\prec_{\mathrm{end}}\mathcal{N}$.
The above together with the fact that $\mathcal{N}$ is conservative extension of $\mathcal{M}$ makes it clear that $\mathcal{M}^{\ast}$ is a conservative extension of $\mathcal{M}$. [^24] By Theorem 4.1, $\mathcal{M}^{\ast}$ is a proper elementary extension of $\mathcal{M}$. We claim that $\mathcal{N}$ is a minimal elementary end extension of $\mathcal{M}^{\ast}$. First note that $\mathcal{N}$ is generated by members of $\mathcal{M}$ and the $\mathcal{U}$-equivalence class of the identity function $\mathrm{id}]_{\mathcal{U}}$ via the definable terms of $\mathcal{M}$, i.e., the following holds:
For any $a\in N$, there exists some $\mathcal{M}$-definable function $f$ such that $\mathcal{N}\models f([\mathrm{id}]_{\mathcal{U}})=a.$
To verify that $\mathcal{N}$ is a minimal elementary end extension of $\mathcal{M}^{\ast}$ it is sufficient to show that if $s\in N\backslash M^{\ast}$ then $\mathrm{id}]_{\mathcal{U}}$ is definable in $\mathcal{N}$ using parameters from $M\cup\{s\}.$ Choose $h$ such that $s=[h]_{\mathcal{U}}$ and recall that by our construction of $\mathcal{U}$ we know that there exists $X\subseteq\mathrm{Ord}^{\mathcal{M}}$ with $X\in\mathcal{U}$ such that either $h\upharpoonright X$ is one-to-one, or the image of $X$ under $h$ is a set (and thus of bounded $\mathcal{M}$-rank). However, since $s$ was chosen to be in $N\backslash M^{\ast}$ we can rule out the latter possibility. Note that since $X\in\mathcal{U}$ the extension of $X$ in $\mathcal{N}$ will contain $\mathrm{id}]_{\mathcal{U}}$, i.e., if $X$ is defined in $\mathcal{M}$ by $\varphi(x)$, then $\mathcal{N}\models\varphi([\mathrm{id}]_{\mathcal{U}})$. Moreover, by elementarity, since $h$ is one-to-one on $X$ in $\mathcal{M}$, it must remain so in $\mathcal{N}$. Hence, in light of the fact that $\mathcal{N}\models h([\mathrm{id}]_{\mathcal{U}})=s$, we can define $\mathrm{id}]_{\mathcal{U}}$ within $\mathcal{N}$ to be the unique element of $h^{-1}(s)\cap X.$ This completes the proof. $\square$
3.5. Remark. As Theorem 4.1 of the next section shows, in Theorem 3.1 the model $\mathcal{N}$ cannot be required to fix $\omega^{\mathcal{M}}$ and in particular $\mathcal{N}$ cannot be arranged to end extend $\mathcal{M}$. By putting this observation together with Remark 2.6(j) we can conclude that in the proof of Corollary 3.2, $\omega^{\mathcal{N}_{\kappa}}$ is $\kappa$-like (i.e., its cardinality is $\kappa$ but each proper initial segment of it is of cardinality less than $\kappa$), and $\omega^{\mathcal{N}_{\mathrm{Ord}}}$ is $\mathrm{Ord}$-like (i.e., it is a proper class every proper initial segment of which forms a set).
We next address the question of the extent to which the hypothesis in Theorem 3.1 that $\mathcal{M}$ is a model of $\exists p\left(\mathrm{V}=\mathrm{HOD}(p)\right)$ can be weakened. In what follows $\mathrm{AC}$ stands for the axiom of choice, and $\mathsf{DC}$ stands for the axiom of dependent choice.
3.6. Proposition. Suppose $\mathcal{M}$ and $\mathcal{N}$ are models of $\mathsf{ZF}$, $\mathcal{M}\prec_{\mathrm{cons}}\mathcal{N}$, and $\mathcal{N}$ is taller than $\mathcal{M}$. If every set can be linearly ordered in $\mathcal{N}$, then there is an $\mathcal{M}$-definable global linear ordering.
Proof. Suppose $\mathcal{M}$ and $\mathcal{N}$ are as in the assumptions of the proposition. Fix $\alpha\in\mathrm{Ord}^{\mathcal{N}}$ such that $\alpha$ is above all the ordinals of $\mathcal{M}$ and let $R$ denote a linear ordering of $\mathrm{V}_{\alpha}^{\mathcal{N}}$ in $\mathcal{N}$. By conservativity, there is a formula $r(x,y)$ (possibly with parameters from $M$) such that $r(x,y)$ that defines $R\cap M^{2}$ in $\mathcal{M}$ . It is clear that $r(x,y)$ defines a global linear ordering in $\mathcal{M}$. $\square$
3.7. Corollary. There are models of $\mathsf{ZFC}$ that have no taller conservative elementary extensions to a model of $\mathsf{ZFC}$.
Proof. This follows from Proposition 3.6, and the well-known fact that there are models of $\mathsf{ZFC}$ in which the universe is not definably linearly ordered. [^25] Easton proved (in his unpublished dissertation [Ea]) that assuming $\mathrm{Con}(\mathsf{ZF})$ there is a model $\mathcal{M}$ of $\mathsf{ZFC}$ that carries no $\mathcal{M}$-definable global choice function for the class of pairs in $\mathcal{M}$; and in particular the universe cannot be definably linearly ordered in $\mathcal{M}$. Easton’s theorem was exposited by Felgner [F, p.231]; for a more recent and streamlined account, see Hamkins’ MathOverflow answer [H]. $\square$
3.8. QuestionCan the conclusion of Proposition 3.6 be improved to $\mathcal{M}\models\exists p\left(\mathrm{V}=\mathrm{HOD}(p)\right)$?
[…]
3.10. Remark. Theorem 4.1 should also be contrasted with Gaifman’s theorem mentioned in Remark 2.5 which bars taller elementary self-extensions of models of (a fragment of) $\mathsf{ZF}$. Thus in Gaifman’s theorem, (a) the condition $\mathcal{N}$ is an elementary self extension of $\mathcal{M}$” cannot be weakened to $\mathcal{N}$ is an elementary conservative extension of $\mathcal{M}$”. Also note that in the first proof of Theorem 3.1, each $\mathcal{N}_{k}$ is a self-extension of $\mathcal{M}$ whose associated embedding $j$ is a $\Sigma_{k+2}$-elementary embedding whose image is not cofinal in $\mathcal{M}$.
4FAITHFUL EXTENSIONS
This section contains the proof of Theorem B of the abstract (see Corollary 4.2). Recall from Definition 2.4 and Remark 2.5 that the notion of a faithful extension is a generalization of the notion of a conservative extension. The notion of a full satisfaction class was defined in Definition 2.7(b).
4.1. Theorem. Suppose $\mathcal{M}$ and $\mathcal{N}$ are both models of $\mathsf{ZFC}$ such that $\mathcal{M}\prec_{\Delta_{0}^{\mathcal{P}},\mathrm{faith}}\mathcal{N}$, $\mathcal{N}$ fixes $\omega^{\mathcal{M}},$ and $\mathcal{N}$ is taller than $\mathcal{M}$. Then:
  1. (a)
    There is some $\gamma\in\mathrm{Ord}^{\mathcal{N}}$ such that $\mathcal{M}\preceq\mathcal{N}_{\gamma}.$ Thus either $\mathcal{M=N}_{\gamma}$, in which case $\mathcal{N}$ is a topped rank extension of $\mathcal{M}$, or $\mathcal{M}\prec\mathcal{N}_{\gamma}$. [^26] As pointed out in Remark 4.5, both cases of the dichotomy can be realized.
  2. (b)
    In particular, there is a full satisfaction class $S$ for $\mathcal{M}$ such that $S$ can be written as $D\cap M$, where $D$ is $\mathcal{N}$-definable.
Proof [^27] The proof strategy of Theorem 4.1 is a variant of the proofs of the following results: [En-2, Theorem 1.5], [En-3, Theorem 3.3], [EH, Theorem 2.1], [En-6, Theorem 2.1.3], and [En-7, Lemma 2.19]. A variant of the same strategy is used in the proof of Theorem 4.5. The origins of the strategy can be traced to Kaufmann’s refinement [Kau, Lemma 1.4] of a Skolem hull argument due to Keisler and Silver [KS, Theorem 2.1].. Fix some $\mathcal{N}$-ordinal $\lambda\in\mathrm{Ord}^{\mathcal{N}}$ that dominates each $\mathcal{M}$-ordinal, and some $\mathcal{N}$-ordinal $\beta>\lambda$. Note that:
(1) $\mathcal{N}_{\beta}\prec_{\Delta_{0}^{\mathcal{P}}}\mathcal{N}.$
[…]
$K_{m}:=\{a\in\mathrm{V}:\mathcal{N}\models a\in\mathrm{Def}(\mathrm{V}_{\beta},\in,\vartriangleleft,\lambda,m)\},$
where $x\in\mathrm{Def}(\mathrm{V}_{\beta},\in,\vartriangleleft,\lambda,m)$ is shorthand for the formula of set theory that expresses:
$x$ is definable in the structure $(\mathrm{V}_{\beta},\in,\vartriangleleft,\lambda,m).$
Note that the only allowable parameters used in a definition of $x$ are $\lambda$ and $m$. Thus, intuitively speaking, within $\mathcal{N}$ the set $K_{m}$ consists of elements $a$ of $\mathrm{V}_{\beta}$ such that $a$ is first order definable in $\left(\mathrm{V}_{\beta},\in,\vartriangleleft,\lambda,m\right)$. Also note that since we are not assuming that $\mathcal{M}$ is $\omega$-standard, an element $a\in K_{m}$ need not be definable in the structure $\left(\mathcal{N}_{\beta},\vartriangleleft,\lambda,m\right)$ in the real world. Next we move outside of $\mathcal{N}$ and define $K$ as follows:
$K:=\bigcup\limits_{m\in M}K_{m}.$
[…]
By putting (1) and (2) together we conclude:
[…]
Let $O^{\ast}$ be the collection of ‘ordinals’ of $\mathcal{K}$ that are above the ‘ordinals’ of $\mathcal{M}$, i.e.,
$O^{\ast}=\left\{\gamma\in\mathrm{Ord}^{\mathcal{K}}:\forall\alpha\in\mathrm{Ord}^{\mathcal{M}}\ \mathcal{N}\models\alpha\in\gamma\right\}.$
Clearly $O^{\ast}$ is nonempty since $\lambda\in O$ . We now consider the following two cases. As we shall see, Case I leads to the conclusion of the theorem, and Case II is impossible.
Case I. $O^{\ast}$ has a least ordinal (under $\in^{\mathcal{K}}).$
Case II. $O^{\ast}$ has no least ordinal.
Suppose Case I holds and let $\gamma_{0}=\min(O^{\ast})$. Clearly $\gamma_{0}$ is a limit ordinal of $\mathcal{N}$. By the choice of $\gamma_{0},$ $\mathrm{Ord}^{\mathcal{M}}$ is cofinal in $\mathrm{Ord}^{\mathcal{K}_{\gamma_{0}}}.$ Since by (3) and part (d) of Remark 2.6, $\mathrm{V}_{\alpha}^{\mathcal{M}}=\mathrm{V}_{\alpha}^{\mathcal{K}}$ for each $\alpha\in\mathrm{Ord}^{\mathcal{M}}$, this makes it clear that:
[…]
Since $\mathcal{K}_{\gamma_{0}}$ is the convex hull of $\mathcal{M}$ in $\mathcal{K}$, by Gaifman Splitting Theorem, (4) shows:
[…]
On the other hand since by (3) $\mathcal{K}\preceq\mathcal{N}_{\beta}$, we have:
(6) $\mathcal{K}_{\gamma_{0}}=\mathrm{V}_{\gamma_{0}}^{\mathcal{K}}\preceq\mathrm{V}_{\gamma_{0}}^{\mathcal{N}_{\beta}}=\mathrm{V}_{\gamma_{0}}^{\mathcal{N}}=\mathcal{N}_{\gamma_{0}}.$
By (5) and (6), $\mathcal{M}\preceq\mathcal{N}_{\gamma_{0}}$. Thus the proof of the theorem will be complete once we show that Case II leads to a contradiction. Within $\mathcal{N}$ let $s_{\beta}$ be the elementary diagram of $\left(\mathrm{V}_{\beta},\in\right)$, i.e., $\mathcal{N}$ views $s_{\beta}$ is the Tarskian satisfaction class for $\left(\mathrm{V}_{\beta},\in\right)$, and let $\Phi:=\bigcup\limits_{m\in M}\Phi_{m}$, where:
$\Phi_{m}:=\{x\in M:\mathcal{N}\models$ $x$ is (the code of) a formula $\varphi(c,\dot{m})$ such that $\varphi(\dot{\lambda},\dot{m})\in s_{\beta}\}\mathrm{.}$
So intuitively speaking, $\Phi$ is the result of replacing $\dot{\lambda}$ by $c$ (where $c$ is a fresh constant) in the sentences in the elementary diagram of $\mathcal{N}_{\beta}$ (as computed in $\mathcal{N}$) whose constants are in $\{\dot{\lambda}\}\cup\{\dot{m}:m\in M\}.$ Since $\mathcal{N}$ need not be $\omega$-standard, the elements of $\Phi$ might be nonstandard formulae.
By the assumption that $\mathcal{N}$ is a faithful extension of $\mathcal{M}$, $\Phi$ is $\mathcal{M}$-amenable. Next let:
$\Gamma:=\left\{\begin{array}[]{c}t(c,\dot{m})\in M:t(c,\dot{m})\in\Phi,\ \mathrm{and}\ \forall\theta\in\mathrm{Ord\ }\left(t(c,\dot{m})>\dot{\theta}\right)\in\Phi,\ \mathrm{and}\\ t\ \mathrm{is\ a\ definable\ term\ in\ the\ language}\ \mathcal{L}_{\mathrm{set}}\cup\{c\}\cup\{\dot{m}:m\in M\}\end{array}\right\}.$
Officially speaking, $\Gamma$ consists of syntactic objects $\varphi(c,\dot{m},x)$ in $\mathcal{M}$ that satisfy the following three conditions in $(\mathcal{M},\Phi)$:
  1. $(i)$
    $\left[\exists!x\varphi(c,\dot{m},x)\right]\in\Phi.$
  2. $(ii)$
    $\left[\forall x\left(\varphi(c,\dot{m},x)\rightarrow x\in\mathrm{Ord}\right)\right]\in\Phi.$
  3. $(iii)$
    $\forall\theta\in\mathrm{Ord}\ \left[\forall x\left(\varphi(c,\dot{m},x)\rightarrow\dot{\theta}\in x\right)\right]\in\Phi.$
Note that $\Gamma$ is definable in $\left(\mathcal{M},\Phi\right)$. Since we are considering Case II, $\left(\mathcal{M},\Phi\right)\models\psi$, where:
$\psi:=\forall t\left[t\in\Gamma\longrightarrow\exists t^{\prime}\in\Gamma\ (t^{\prime}\in t)\in\Phi\right].$
The veracity of the dependent choice scheme in $\mathcal{M}$ (see part (f) of Definition 2.1), together with the facts that $\Phi$ is $\mathcal{M}$-amenable and $\Gamma$ is definable in $\left(\mathcal{M},\Phi\right)$ make it clear that there is a sequence $s=\left\langle t_{n}:n\in\omega^{\mathcal{M}}\right\rangle$ in $\mathcal{M}$ such that:
$\left(\mathcal{M},\Phi\right)\models\forall n\in\omega\ \left[t_{n}\in\Gamma\wedge\left(\left(t_{n+1}\in t_{n}\right)\in\Phi\right)\right].$
Since $s$ is a countable object in $\mathcal{M}$, and $\mathcal{N}$ fixes $\omega^{M}$ by assumption, $s$ is fixed in the passage from $\mathcal{M}$ to $\mathcal{N}$ by Remark 2.6(c). On the other hand, since $\mathcal{N}$ has a satisfaction predicate $s_{\beta}$ for $\mathcal{N}_{\beta}$, this leads to a contradiction because we have:
$\mathcal{N}\models\left\langle t_{n}^{(\mathrm{V}_{\beta},\in)}(\dot{\lambda}):n\in\omega\right\rangle$ is an infinite decreasing sequence of ordinals.
In the above $t_{n}(\dot{\lambda})$ is the term obtained by replacing $c$ with $\dot{\lambda}$ in $t_{n},$ and $t_{n}^{(\mathrm{V}_{\beta},\in)}(\dot{\lambda})$ is the interpretation of $t_{n}(\dot{\lambda})$ in $(\mathrm{V},\in)$ using the Tarskian satisfaction class for $(\mathrm{V}_{\beta},\in)$. $\square$
4.2. Corollary. If $\mathcal{M}\models\mathsf{ZFC}$, $\mathcal{M}\prec_{\Delta_{0}^{\mathcal{P}},\ \mathrm{cons}}\mathcal{N}$, and $\mathcal{M}$ fixes $\omega^{\mathcal{M}},$ then $\mathcal{M}$ is cofinal in $\mathcal{N}$. Thus a conservative elementary extension of $\mathcal{M}$ that fixes $\omega^{\mathcal{M}}$ is a cofinal extension.
Proof. If $\mathcal{M}\models\mathsf{ZFC}$, $\mathcal{M}\prec_{\mathrm{cons}}\mathcal{N}$, and $\mathcal{M}$ is not cofinal in $\mathcal{N}$, then $\mathcal{N}$ is taller than $\mathcal{M}$. But then by part (b) of Theorem 4.1, there is an $\mathcal{M}$-definable full satisfaction class over $\mathcal{M}$, which contradicts Tarski’s Undefinability of Truth Theorem. $\square$
In Theorem 4.4 below we show that the conclusion of Theorem 4.1 holds if $\mathsf{ZFC}$ is weakened to $\mathsf{ZF}\mathrm{,}$ but the assumption that $\mathcal{N}$ fixes $\omega^{\mathcal{M}}$ is strengthened to $\mathcal{M}\subsetneq_{\mathrm{end}}\mathcal{N}$. We first establish a lemma that will come handy in the proof of Theorem 4.4.
4.3. Lemma Suppose $\mathcal{M}$ and $\mathcal{N}$ are models of $\mathsf{ZF}$, and $\mathcal{M}\subsetneq_{\mathrm{end,faithful}}\mathcal{N}$. Then $\mathcal{M}\subsetneq_{\mathrm{rank}}\mathcal{N}$.
Proof. Suppose $\mathcal{M}$ and $\mathcal{N}$ are as in the assumptions of the lemma. Since $\mathcal{N}$ end extends $\mathcal{M}$, in order to show that $\mathcal{N}$ rank extends $\mathcal{M}$ it is sufficient that if $\alpha\in\mathrm{Ord}^{\mathcal{M}}$ and $m\in M$ with $\mathcal{M}\models m=\mathrm{V}_{\alpha}$, then $\mathcal{N}\models m=\mathrm{V}_{\alpha}.$ For this purpose, it is sufficient to verify that $\mathcal{N}$ is a powerset-preserving end extension of $\mathcal{M}$, since the formula expressing $x=\mathrm{V}_{\alpha}$ is $\Sigma_{1}^{\mathcal{P}}$ and is therefore absolute for powerset-preserving end extensions (as noted in Remark 2.6(e)). Towards this goal, let $a\in M$ and suppose $b\in N$ such that:
[…]
Let $X_{b}=\mathrm{Ext}_{\mathcal{N}}(b)\cap\mathrm{Ext}_{\mathcal{M}}(a)=\mathrm{Ext}_{\mathcal{N}}(b).$ Thanks to the assumption that $\mathcal{N}$ is a faithful extension of $\mathcal{M}$, the expansion $(\mathcal{M},X_{b})$ satisfies Separation in the extended language and therefore there is an element $b^{\prime}\in M$ that codes $X_{b}.$ Thus
[…]
Since $\mathcal{N}$ end extends $\mathcal{M}$, we have:
(3) $\mathrm{Ext}_{\mathcal{M}}(a)=\mathrm{Ext}_{\mathcal{N}}(a)$ and $\mathrm{Ext}_{\mathcal{M}}(b^{\prime})=\mathrm{Ext}_{\mathcal{N}}(b^{\prime}).$
So by (1) and (2),
(4) $\mathrm{Ext}_{\mathcal{N}}(b^{\prime})=\mathrm{Ext}_{\mathcal{N}}(b)\cap\mathrm{Ext}_{\mathcal{N}}(a).$
By (1), $\mathrm{Ext}_{\mathcal{N}}(b)\subseteq\mathrm{Ext}_{\mathcal{N}}(a),$ so combined with (4) this yields:
[…]
which by Extensionality makes it clear that $b^{\prime}=b.$ $\square$
  • The proof of Theorem 4.4 below is a variant of the proof of Theorem 4.1, and will strike the reader who has worked through the proof of Theorem 4.1 as repetitious, but the proof is a bit more involved here since we do not have access to the axiom of choice in $\mathcal{M}$ or in $\mathcal{N}$. Note that in Theorem 4.4 $\mathcal{N}$ is assumed to end extend $\mathcal{M}$, whereas in Theorem 4.1 $\mathcal{N}$ is only assumed to fix $\omega^{\mathcal{M}}.$ Of course Theorem 4.4 follows from Theorem 4.1 if $\mathcal{M}$ and $\mathcal{N}$ are models of $\mathsf{ZFC}$.
4.4. Theorem. Suppose $\mathcal{M}$ and $\mathcal{N}$ are models of $\mathsf{ZF}$, and $\mathcal{M}\subsetneq_{\mathrm{end,faithful}}\mathcal{N}$. Then:
  1. (a)
[…]
  2. (b)
    In particular, there is a full satisfaction class $S$ for $\mathcal{M}$ such that $S$ can be written as $D\cap M$, where $D$ is $\mathcal{N}$-definable.
[…]
$O_{m}:=\left\{\gamma\in\beta:\mathcal{N}\models\gamma\in\mathrm{Def}(\mathrm{V}_{\beta},\in,\lambda,m)\right\}.$
where $x\in\mathrm{Def}(\mathrm{V}_{\beta},\in,\lambda,m)$ is shorthand for the formula of set theory that expresses:
$x$ is definable in $(\mathrm{V}_{\beta},\in)$ using at most the parameters $\lambda$ and $m.$
Clearly $\lambda\in O_{m}$ and $O_{m}$ is coded in $\mathcal{N}$. Next let:
$O:=\bigcup\limits_{m\in M}O_{m}.$
[…]
  1. (I)
    $O^{\ast}$ has a least element (under $\in^{\mathcal{N}}).$
  2. (II)
    $O^{\ast}$ has no least element.
Case I. Let $\gamma_{0}$ be the least element of $O^{\ast}.$ It is easy to see that $\gamma_{0}$ is a limit ordinal in $\mathcal{N}$. We claim that:
$\mathcal{M}\preceq\mathcal{N}_{\gamma_{0}}.$
To verify the claim, by the Tarski criterion of elementarity, it suffices to show that if $\mathcal{N}_{\gamma_{0}}\models\exists y\varphi(y,m)$ for some $m\in M$, then there is some $m_{0}\in M$ such that $\mathcal{N}_{\gamma_{0}}\models\varphi(m_{0},m).$ Let $\gamma_{m}$ be the first ordinal below $\gamma_{0}$ such that:
$\mathcal{N}_{\gamma_{0}}\models\exists y\in\mathrm{V}_{\gamma_{m}}\ \varphi(y,m).$
So $\gamma_{m}$ is definable in $\mathcal{N}_{\beta}$ with parameters $\gamma_{0}$ and $m$, and since $\gamma_{0}\in O$ by assumption, this shows that $\gamma_{m}$ is definable in $\mathcal{N}_{\beta}$ with parameter $m.$ So $\gamma_{m}\in O.$ By the fact that $\gamma_{0}$ was chosen to be the least element of $O^{\ast}$, this shows that $\gamma_{m}\in\mathrm{Ord}^{\mathcal{M}},$ which makes it clear that there is some $m_{0}\in M$ such that:
$\mathcal{N}_{\gamma_{0}}\models\varphi(m_{0},m).$
This concludes the verification of $\mathcal{M}\preceq\mathcal{N}_{\gamma_{0}}$. To finish the proof of Theorem 4.4 it suffices to show that Case II is impossible. Suppose to the contrary that $O^{\ast}$ has no least element. Within $\mathcal{N}$ let $s_{\beta}$ be the Tarskian satisfaction class for $\left(\mathrm{V}_{\beta},\in\right)$. For each $m\in M$ we define:
$\Phi_{m}:=\{\varphi(c,\dot{m})\in M:\mathcal{N}\models$ $\varphi(\dot{\lambda},\dot{m})\in s_{\beta}\}.$
Note the constant $c$ is interpreted as $\lambda$ in the right-hand side of the above definition of $\Phi_{m}.$ Next we define:
$\Phi:=\bigcup\limits_{m\in M}\Phi_{m}$.
Observe that $\Phi\subseteq M$ and $\Phi$ is $\mathcal{M}$-amenable since $\mathcal{N}$ is a faithful rank extension of $\mathcal{M}$. So intuitively speaking, $\Phi$ is the subset of the elementary diagram of $\mathcal{N}_{\beta}$ (as computed in $\mathcal{N}$) that consists of sentences whose constants are in $\{c\}\cup\{\dot{m}:m\in M\}.$ Since $\mathcal{N}$ need not be $\omega$-standard, the elements of $\Phi$ need not be standard formulae. Also note that the constant $c$ is interpreted as $\lambda$ in the right-hand side of the above definition of $\Phi_{m}.$ Now let:
$\Gamma:=\left\{\begin{array}[]{c}t(c,\dot{m})\in M:t(c,\dot{m})\in\Phi,\ \mathrm{and}\ \forall\theta\in\mathrm{Ord\ }\left(t(c,\dot{m})>\dot{\theta}\right)\in\Phi,\ \mathrm{and}\\ t\ \mathrm{is\ a\ definable\ term\ in\ the\ language}\ \mathcal{L}_{\mathrm{set}}\cup\{c\}\cup\{\dot{m}:m\in M\}\end{array}\right\},$
So, officially speaking, $\Gamma$ consists of syntactic objects $\varphi(c,\dot{m},x)$ in $\mathcal{M}$ that satisfy the following three conditions in $(\mathcal{M},\Phi)$:
  1. $(i)$
    $\left[\exists!x\varphi(c,\dot{m},x)\right]\in\Phi.$
  2. $(ii)$
    $\left[\forall x\left(\varphi(c,\dot{m},x)\rightarrow x\in\mathrm{Ord}\right)\right]\in\Phi.$
  3. $(iii)$
    $\forall\theta\in\mathrm{Ord}\ \left[\forall x\left(\varphi(c,\dot{m},x)\rightarrow\dot{\theta}\in x\right)\right]\in\Phi.$
Clearly $\Gamma$ is definable in $\left(\mathcal{M},\Phi\right)$. Since we are in Case II, $\left(\mathcal{M},\Phi\right)\models\psi$, where $\psi$ is the sentence that expresses:
$\Gamma\neq\varnothing\wedge\forall t\left[t\in\Gamma\longrightarrow\exists t^{\prime}\in\Gamma\ (t^{\prime}\in t)\in\Phi\right].$
In contrast to the proof of Theorem 4.1 we cannot at this point conclude that there is a countable descending chain $\left\langle t_{n}:n\in\omega\right\rangle$ in $\mathcal{M}$ since $\mathsf{DC}$ need not hold in $\mathcal{M}$. Instead, we will use the following argument that takes advantage of the Reflection Theorem. Choose $k\in\omega$ such that $\psi$ is a $\Sigma_{k}(\Phi)$-statement, and use the Reflection Theorem in $\left(\mathcal{M},\Phi\right)$ to pick $\mu\in\mathrm{Ord}^{\mathcal{M}}$ such that:
$\left(\mathcal{M}_{\mu},\Phi\cap M_{\mu}\right)\prec_{\Sigma_{k}(\mathrm{X})}\left(\mathcal{M},\Phi\right).$ [^28] Here $\mathrm{X}$ is interpreted by the left-hand side structure as $\Phi\cap M_{\mu}$ and by the right-hand side structure as $\Phi$.
Then $\psi$ holds in $\left(\mathcal{M}_{\mu},\Phi\cap M_{\mu}\right)$. Within $\mathcal{M}$, let
$w=\{v\in\mathrm{V}_{\mu}:v\in\Gamma\}.$
Thus $\left(\mathcal{M},\Phi\right)\models$ $w\neq\varnothing\wedge\exists t^{\prime}\in\Gamma\ (t^{\prime}\in t)\in\Phi.$ Observe that since $\mathcal{N}$ has access to the Tarskian satisfaction class for $\mathcal{N}_{\beta}$, $\mathcal{N}$ can evaluate each term $t(c)$ in $w$ as an ordinal $\delta<\beta$, where $\mathcal{N}_{\beta}\models t(\dot{\lambda})=\delta$, where $t(\dot{\lambda})$ is the term obtained by replacing all occurrences of the constant $c$ with $\dot{\lambda}$ in $t(c).$ So we can consider $s\in N$, where:
$\mathcal{N}\models s=\left\{\delta:\exists t(c)\in w\ (\mathrm{V}_{\beta},\in)\models\delta=t(\dot{\lambda})\right\}.$
Within $\mathcal{N}$, $s$ is a nonempty set of ordinals that has no least element, which of course is a contradiction. This shows that Case II is impossible, thus concluding the proof. $\square$
4.5. Remark. Arguing in $\mathsf{ZFC}$, suppose $\kappa$ is the first strongly inaccessible cardinal, $\lambda$ is the second strongly inaccessible cardinal, $\mathcal{M}=(\mathrm{V}_{\kappa},\in)$ and $\mathcal{N}=(\mathrm{V}_{\lambda},\in)$. Then $\mathcal{N}$ is a proper faithful rank extension of $\mathcal{M}$ in which the first clause of the dichotomy of the conclusion of part (a) of Theorem 4.1 (and 4.4) holds , but not the second one. On the other hand, if $\kappa$ is a weakly compact cardinal, then as noted by Kaufmann [Kau, Proposition 2.3], $(\mathrm{V}_{\kappa},\in)$ has a faithful topless elementary end extension $\mathcal{N}$; in this scenario the second clause of the dichotomy of the conclusion of part (a) of Theorem 4.1 (and 4.4) holds, but not the first one.
4.6. Remark. If the assumptions of Theorem 4.4 are strengthened by adding the assumption that $\mathcal{M}\prec\mathcal{N}$, then as in the proof of Theorem 3.3 of [En-3] we can use ‘Kaufmann’s trick’ to conclude that there is some $\mathcal{K}$ such that $\mathcal{M}\prec\mathcal{K}\preceq\mathcal{N}$ and $\mathcal{K}$ is a topped rank extension of $\mathcal{M}$. This yields a strengthening of Theorem 3.3 of [En-3] by eliminating the assumption that the axiom of choice holds in $\mathcal{M}$ and $\mathcal{N}$.
4.7. Remark. An inspection of the proof of Theorem 4.1 makes it clear that the assumption that $\mathcal{M}\models\mathsf{ZFC}$ can be reduced to $\mathcal{M}\models\mathsf{ZF}+\mathsf{DC}$, and the assumption that $\mathcal{N}\models\mathsf{ZFC}$ can be reduced to $\mathcal{N}\models\mathsf{KPR}$ + “every set can be well-ordered”. Here $\mathsf{KPR}$ is as in part (d) of Definition 2.1. Similarly, an inspection of the proof of Theorem 4.4 shows that the assumption that $\mathcal{N}\models\mathsf{ZF}$ can be reduced to $\mathcal{N}\models\mathsf{KPR}$.
4.8. Question. Can Theorem 4.1 and Corollary 4.2 be strengthened by assuming that $\mathcal{M}$ and $\mathcal{N}$ are models of $\mathsf{ZF}$?
5DEAD-END MODELS
In this section we will establish Theorem C of the abstract, the proof uses many ingredients, including the following one that refines a result obtained independently by Kaufmann and the author who demonstrated Theorem 5.1 for models of $\mathsf{ZF}$ in which the Axiom of Choice holds (see “Added in Proof” of [Kau] and Remark 1.6 of [En-2]).
5.1. Theorem. No model of $\mathsf{ZF}$ has a conservative proper end extension satisfying $\mathsf{ZF}$.
Proof. This follows from putting part (b) of Theorem 4.4 together with Tarski’s Undefinability of Truth Theorem (as in Theorem 2.8). $\square.$
The rest of the section is devoted to presenting results that in conjunction with Theorem 5.1 will allow us to establish Theorem C of the abstract (as Theorem 5.18).
5.2. Definition. Suppose $\mathcal{M}$ is an $\mathcal{L}_{\mathrm{set}}$-structure.
  1. (a)
    $X\subseteq M$ is a class [^29] Classes of $\mathcal{M}$ are sometimes referred to as piecewise coded subsets of $\mathcal{M}$. in $\mathcal{M}$ if $\forall a\in M\ \exists b\in M\ X\cap\mathrm{Ext}_{\mathcal{M}}(a)=\mathrm{Ext}_{\mathcal{M}}(b).$
  2. (b)
    $\mathcal{M}$ is rather classless if every class of $M$ is $\mathcal{M}$-definable.
  3. (c)
    $\mathcal{M}$ is $\aleph_{1}$-like if $\left|M\right|=\aleph_{1}$ but $\left|\mathrm{Ext}_{\mathcal{M}}(a)\right|\leq\aleph_{0}$ for each $a\in M.$
5.3. Theorem. (Keisler-Kunen [Ke-1], Shelah [Sh]). Every countable model of $\mathsf{ZF}$ has an elementary end extension to an $\aleph_{1}$-like rather classless model.
5.4. Proposition. No rather classless model of $\mathsf{ZF}$ has a proper rank extension to a model of $\mathsf{ZF}$.
Proof. This follows from putting Theorem 4.4 together with the observation that a rank extension of a rather classless model of $\mathsf{ZF}$ is a conservative extension, and therefore a faithful extension. $\square$
5.5. Remark. As pointed out in Remark 1.6 of [En-3] it is possible for a rather classless model to have a proper end extension satisfying $\mathsf{ZF}$, since $\aleph_{1}$-like rather classless models exist by Theorem 5.3, and one can use the Boolean-valued approach to forcing to construct set generic extensions of such models.
5.6. Definition. A ranked tree $\tau$ is a two-sorted structure $\tau=(\mathbb{T},\ \leq_{\mathbb{T}},\ \mathbb{L},\ \leq_{\mathbb{L}},\ \rho)$ satisfying the following three properties:
  1. (1)
    $(\mathbb{T},\ \leq_{\mathbb{T}})$ is a tree, i.e., a partial order such that any two elements below a given element are comparable.
  2. (2)
    $(\mathbb{L},\ \leq_{\mathbb{L}})$ is a linear order with no last element.
  3. (3)
    $\rho$ is an order preserving map from $(\mathbb{T},\ \leq_{\mathbb{T}})$ onto $(\mathbb{L},\ \leq_{\mathbb{L}})$ with the property that for each $t\in\mathbb{T},$ $\rho$ maps the set of predecessors of $t$ onto the initial segment of $(\mathbb{L},\ \leq_{\mathbb{L}})$ consisting of elements of $L$ less than $\rho(t).$
5.7. Definition. Suppose $\tau=(\mathbb{T},\ \leq_{\mathbb{T}},\ \mathbb{L},\ \leq_{\mathbb{L}},\ \rho)$ is a ranked tree. A linearly ordered subset $B$ of $\mathbb{T}$ is said to be a branch of $\tau$ if the image of $B$ under $\rho$ is $\mathbb{L}$. The cofinality of $\tau$ is the cofinality of $(\mathbb{L},\ \leq_{\mathbb{L}}).$
5.8. Definition. Given a structure $\mathcal{M}$ in a language $\mathcal{L}$, we say that a ranked tree $\tau$ is $\mathcal{M}$-definable if $\tau=\mathbf{t}^{\mathcal{M}}$, where $\mathbf{t}$ is an appropriate sequence of $\mathcal{L}$-formulae whose components define the corresponding components of $\tau$ in $\mathcal{M}.$ $\mathcal{M}$ is rather branchless if for each $\mathcal{M}$-definable ranked tree $\tau$, all branches of $\tau$ (if any) are $\mathcal{M}$-definable.
5.9. Theorem. Suppose $\mathcal{M}$ is a countable structure in a countable language.
  1. (a)
    (Keisler-Kunen [Ke-1], essentially). It is a theorem of $\mathsf{ZFC}+\Diamond_{\omega_{1}}$ that $\mathcal{M}$ can be elementarily extended to a rather branchless model.
  2. (b)
    (Shelah [Sh]). It is a theorem of $\mathsf{ZFC}$ that $\mathcal{M}$ can be elementarily extended to a rather branchless model.
5.10. Definition. Suppose $(\mathbb{P},\leq_{\mathbb{P}})$ is a poset (partially ordered set).
  1. (a)
    $(\mathbb{P},\leq_{\mathbb{P}})$ is directed if any given pair of elements of $\mathbb{P}$ has an upper bound.
  2. (b)
    A subset $F$ of $\mathbb{P}$ is a filter over $(\mathbb{P},\leq_{\mathbb{P}})$ if the sub-poset $(F,\leq_{\mathbb{P}})$ is directed.
  3. (c)
    A filter over $(\mathbb{P},\leq_{\mathbb{P}})$ is maximal if it cannot be properly extended to a filter over $(\mathbb{P},\leq_{\mathbb{P}})$. [^30] Thus a maximal filter can be described as a maximally compatible subset of $(\mathbb{P},\leq_{\mathbb{P}}).$
  4. (d)
    A subset $C$ of $\mathbb{P}$ is cofinal in $(\mathbb{P},\leq_{\mathbb{P}})$ if $\forall x\in\mathbb{P}\ \exists y\in C\ x\leq_{\mathbb{P}}y.$
5.11. Definition. Suppose $s$ is an infinite set.
  1. (a)
    $s]^{<\omega}$ is the directed poset of finite subsets of $s,$ ordered by containment. Note that $s]^{<\omega}$ is a directed set with no maximum element.
  2. (b)
    $\mathrm{Fin}(s,2)$ is the poset of finite functions from $s$ into $\{0,1\}$, ordered by containment (where a function is viewed as a set of ordered pairs).
5.12. Example. Given an infinite set $a$, and $s\subseteq a,$ let $\chi_{s}:s\rightarrow 2$ be the characteristic function of $s,$ i.e., $\chi_{s}(x)=1$ iff $x\in s.$ Clearly $\chi_{s}]^{<\omega}$ is a maximal filter of $\mathrm{Fin}(a,2).$ More generally, if $\mathcal{N}\models\mathsf{ZF}$, and $\mathcal{N}\modelssas\subseteq$ a and $a$ is infinite”, consider $(\mathbb{P},\leq_{\mathbb{P}})$, where:
$\mathbb{P}:=\mathrm{Ext}_{\mathcal{N}}(\mathrm{Fin}^{\mathcal{N}}(a,2))$,
and $\leq_{\mathbb{P}}$ is set-inclusion (among members of $\mathbb{P}$) in the sense of $\mathcal{N}$. Then $F_{s}$ is a maximal filter over $(\mathbb{P},\leq_{\mathbb{P}})$, where:
$F_{s}:=$ $\mathrm{Ext}_{\mathcal{N}}\left([\chi_{s}]^{<\omega}\right)^{\mathcal{N}}$.
5.13. Definition. A structure $\mathcal{M}$ is a Rubin model if it has the following two properties:
  1. (a)
    Every $\mathcal{M}$-definable directed poset $(\mathbb{D},\leq_{\mathbb{D}})$ with no maximum element has a cofinal chain of length $\omega_{1}.$
  2. (b)
    Given any $\mathcal{M}$-definable poset $\mathbb{P}$, and any maximal filter $F$ of $\mathbb{P}$, if $F$ has a cofinal chain of length $\omega_{1}$, then $F$ is coded in $\mathcal{M}$.
5.14. Remark. If $\tau=(\mathbb{T},\ \leq_{\mathbb{T}},\ \mathbb{L},\ \leq_{\mathbb{L}},\ \rho)$ is a ranked tree, then each branch of $\tau$ is a maximal filter over $(\mathbb{T},\ \leq_{\mathbb{T}}).$ This makes it clear that every Rubin model is rather branchless. Also, a rather branchless model of $\mathsf{ZF}$ is rather classless. To see this consider the ranked tree $\tau_{\mathrm{class}}$ defined within a model $\mathcal{M}$ of $\mathsf{ZF}$ as follows: The nodes of $\tau_{\mathrm{class}}$ are ordered pairs $(s,\alpha)$, where $s\subseteq\mathrm{V}_{\alpha}$, the rank of $(s,\alpha)$ is $\alpha$ and $(s,\alpha)<(t,\beta)$ if $\alpha\in\beta$ and $s=t\cap\mathrm{V}_{\alpha}$. It is easy to see that $\mathcal{M}$ is rather classless iff every branch of $\tau_{\mathrm{class}}^{\mathcal{M}}$ is $\mathcal{M}$-definable. [^31] The ranked tree $\tau_{\mathrm{class}}$ was first introduced by Keisler [Ke-1, Example 2.1], who noted that $\mathcal{M}$ is rather classless iff every branch of $\tau_{\mathrm{class}}^{\mathcal{M}}$ is $\mathcal{M}$-definable, which is the key property that we need here. As noted by the referee, this key property is also satisfied by the subtree $\sigma_{\mathrm{class}}$ of $\tau_{\mathrm{class}}$ whose tree-nodes are of the form $\left(\alpha,s\right)$ where the ordinal rank of $s$ is $\alpha.$ Note that the branches of $\tau_{\mathrm{class}}^{\mathcal{M}}$ correspond to all classes of $\mathcal{M}$ (including improper ones, i.e., those that form a set), whereas the branches of $\sigma_{\mathrm{class}}^{\mathcal{M}}$ correspond to all proper classes of $\mathcal{M}$. Hence we have the following chain of implications:
Rubin $\Rightarrow$ rather branchless $\Rightarrow$ rather classless.
5.15. Theorem. (Rubin [^32] The attribution to Rubin is informed by fact that the credit for [RS, Lemma 2.3] is explicitly given to Rubin (at the end of the introduction of the paper). [RS, Corollary 2.4]). It is a theorem of $\mathsf{ZFC}+\Diamond_{\omega_{1}}$ that if $\mathcal{M}$ is a countable structure in a countable language, then $\mathcal{M}$ has an elementary extension of cardinality $\aleph_{1}$ that is a Rubin model. [^33] The proof of Theorem 5.15 in [RS, Lemma 2.3] has a small gap (it is assumed that $T_{0}\vdash\exists\vec{x}\psi(\vec{x})$ instead of assuming the consistency of $T_{0}+\exists\vec{x}\psi(\vec{x})$). The improved proof was presented in [En-1, Theorem 2.1.3] and is reproduced here in the Appendix.
5.16. Definition. Suppose $\mathcal{M}$ is a model of $\mathsf{ZF}$. $\mathcal{M}$ is weakly Rubin if (a) and (b) below hold:
  1. (a)
    $\mathcal{M}$ is rather classless (this is the asymptotic case of $(ii)$ below, if one could use $a=\mathrm{V}$).
  2. (b)
    For every element $a$ of $\mathcal{M}$ that is infinite in the sense of $\mathcal{M}$ the following two statements hold:
    1. $(1)$
      $\left([a]^{<\omega}\right)^{\mathcal{M}}$ has a cofinal chain of length $\omega_{1}.$
    2. $(2)$
      If $F$ is a maximal filter of $\mathrm{Fin}^{\mathcal{M}}(a,2)$ and $F$ has a cofinal chain of length $\omega_{1}$, then $F$ is coded in $\mathcal{M}$.
As demonstrated in the Appendix, Schmerl’s strategy of $\Diamond_{\omega_{1}}$-elimination in [Schm-2] (which is based on an absoluteness argument first presented by Shelah [Sh]) can be employed to build weakly Rubin models within $\mathsf{ZFC}$. Thus we have:
5.17. Theorem (Rubin-Shelah-Schmerl) It is a theorem of $\mathsf{ZFC}$ that every countable model of $\mathsf{ZF}$ has an elementary extension to a weakly Rubin model of cardinality $\aleph_{1}$.
We are finally ready to establish the main result of this section on the existence of ‘dead-end’ models. Note that since every generic extension is an end extension (even for ill-founded models), the model $\mathcal{M}$ in Theorem 5.18 has no proper generic extension.
5.18. Theorem. Every countable model $\mathcal{M}_{0}\models\mathsf{ZF}$ has an elementary extension $\mathcal{M}$ of power $\aleph_{1}$ that has no proper end extension to a model $\mathcal{N}\models\mathsf{ZF}$. Thus every consistent extension of $\mathsf{ZF}$ has a model of power $\aleph_{1}$ that has no proper end extension to a model of $\mathsf{ZF}$.
Proof [^34] The proof of Fact $\left(\nabla\right)$ presented here is a reformulation of a proof that was suggested by the referee. Our original proof admittedly used too much machinery.. We begin with a basic fact that will be called upon in the proof.
Fact $\left(\nabla\right).$ Suppose $\mathcal{M}\models\mathsf{ZF}$ and $\mathcal{N}\models\mathsf{ZF}$ with $\mathcal{M}\subseteq_{\mathrm{end}}\mathcal{N}$, and $a\in M$. If $\in N, ss$ is finite as viewed in $\mathcal{N}$, and $\mathcal{N}\models s\subseteq a$, then $s\in M.$ Thus for all $a\in M$, we have:
$\left(\left[a\right]^{<\omega}\right)^{\mathcal{M}}=\left(\left[a\right]^{<\omega}\right)^{\mathcal{N}}.$
Proof. The assumptions on $\mathcal{M}$ and $\mathcal{N}$ readily imply:
(1) $\omega^{\mathcal{M}}=\omega^{\mathcal{N}}$, and
(2) $\mathcal{M}\preceq_{\Delta_{1}}\mathcal{N}$.
Let $x=\mathrm{Fn}(y,a)$ be the usual formula expressing “$x$ is the set of all functions $f:y\rightarrow a$. It is easy to see that:
(3) The predicate $\left(x=\mathrm{Fn}(y,a)\wedge y\in\omega\right)$ is $\Delta_{1}$ within $\mathsf{ZF.}$
Observe that $x=\mathrm{Fn}(y,a)$ is clearly expressible by a $\Pi_{1}$-formula within $\mathsf{ZF}$ (with no restriction on $y$). With the added condition that $y\in\omega$, we can take advantage of recursion to express $x=\mathrm{Fn}(y,a)$ by the following $\Sigma_{1}$-formula:
$\exists\left\langle s_{0},\cdot\cdot\cdot,s_{y}\right\rangle$ $\left[s_{y}=x\wedge s_{0}=\{\varnothing\}\wedge\forall z<y\ (s_{z+1}=\left\{f\cup\{(z,v)\}:f\in s_{z}\wedge v\in a\right\})\right].$
Now assume $s\in N$, $\mathcal{N}\models s\subseteq a$, and for some $k\in N,$ $\mathcal{N\models}$ $k\in\omega\wedge\left|s\right|=k$. Thanks to (1) and the assumption $\mathcal{M}\subseteq_{\mathrm{end}}\mathcal{N}$, $k\in M.$ Therefore by (2) and (3), if $b\in M$ such that $\mathcal{M}\models b=\mathrm{Fn}(k,a)$, then $\mathcal{N}\models b=\mathrm{Fn}(k,a)$. Thus if $f\in N$ such that $f:k\rightarrow a$ is an injective function such that $\mathrm{range}(f)=s$, then $f\in b$, and since $\mathcal{N}$ is an end extension of $\mathcal{M}$, $f\in M.$ Hence $\mathrm{range}(f)=s\in M.$ $\square$ Fact $\left(\nabla\right)$)
Given a countable model $\mathcal{M}_{0}$ of $\mathsf{ZFC}$, by Theorem 5.17 there is a weakly Rubin model $\mathcal{M}$ that elementarily extends $\mathcal{M}_{0}$. By Theorem 5.1, to prove Theorem 5.18 it is sufficient to verify that every end extension $\mathcal{N}$ of $\mathcal{M}$ that satisfies $\mathsf{ZF}$ is a conservative extension. Towards this goal, suppose $\mathcal{M}\subsetneq_{\mathrm{end}}\mathcal{N}\models\mathsf{ZF}$. In light of Remark 2.6(e), it suffices to show that $\mathcal{M}\subsetneq_{\mathrm{end}}^{\mathcal{P}}\mathcal{N}$.
To show $\mathcal{M}\subsetneq_{\mathrm{end}}^{\mathcal{P}}\mathcal{N}$, suppose that for $a\in M$ and $s\in N$, $\mathcal{N}\models s\subseteq a.$ We will show that $s\in M.$ By Fact $(\nabla)$ we may assume that $a$ is infinite as viewed from $\mathcal{M}$. Note that this implies that $\mathcal{M}$ views $a]^{<\omega}$ as a directed set with no maximum element. Fact $(\nabla)$ assures us that:
[…]
Since $\mathrm{Fin}(m,2)\subseteq[m\times\{0,1\}]^{<\omega}$, $(\ast)$ implies:
[…]
Let $\chi_{s}\in N$ be the characteristic function of $s$ in $\mathcal{N}$, i.e., as viewed from $\mathcal{N}$, $\chi_{s}:a\rightarrow\{0,1\}$ and $\forall x\in a(x\in s\leftrightarrow\chi_{s}(x)=1)$. Let
$F_{s}:=$ $\mathrm{Ext}_{\mathcal{N}}\left([\chi_{s}]^{<\omega}\right)^{\mathcal{N}}.$
As noted in Example 5.12, $F_{s}$ is a maximal filter over $\mathrm{Ext}_{\mathcal{N}}(\mathrm{Fin}^{\mathcal{N}}(a,2))$, so by $(\ast\ast)$ and the assumption that $\mathcal{N}$ end extends $\mathcal{M}$, $F_{s}$ is a maximal filter over $\mathrm{Ext}_{\mathcal{M}}(\mathrm{Fin}^{\mathcal{M}}(a,2))$. Note that directed set $\mathrm{Ext}_{\mathcal{M}}\left([a]^{<\omega}\right)^{\mathcal{M}}$ has a cofinal chain $\left\langle p_{\alpha}:\alpha\in\omega_{1}\right\rangle$ thanks to the assumption that $\mathcal{M}$ is weakly Rubin. Together with $(\ast)$, this readily implies that $F_{s}$ has a cofinal chain $\left\langle q_{\alpha}:\alpha\in\omega_{1}\right\rangle,$ where:
$q_{\alpha}:=(\chi_{s}\upharpoonright p_{\alpha})^{\mathcal{N}}=(\chi_{s}\upharpoonright p_{\alpha})^{\mathcal{M}}.$
Therefore, by the assumption that $\mathcal{M}$ is weakly Rubin, $F_{s}$ is $\mathcal{M}$-definable and thus coded in $\mathcal{M}$ by some $m\in M$. This makes it clear that $\mathcal{M}\models\chi_{s}=\cup m$, and thus $s\in$ $M$ since $s$ is $\Delta_{0}$-definable from $\chi_{s}$. This concludes the verification that $\mathcal{M}\subsetneq_{\mathrm{end}}^{\mathcal{P}}\mathcal{N}$, which as explained earlier, is sufficient to establish Theorem 5.18. $\square$
An examination of the proof of Theorem 5.1, together with Remark 4.7, makes it clear that Theorem 5.1 can be refined as follows:
5.19. TheoremNo model of $\mathsf{KPR+\Pi}_{\infty}^{1}$-$\mathsf{DC}$ has a conservative proper end extension to a model of $\mathsf{KPR}$.
The proof of Theorem 5.18, together with Theorem 5.19 allows us to refine Theorem 5.18 as follows:
5.20. TheoremEvery countable model $\mathcal{M}_{0}\models\mathsf{KPR}+\Pi_{\infty}^{1}$-$\mathsf{DC}$ has an elementary extension $\mathcal{M}$ of power $\aleph_{1}$ that has no proper end extension to a model $\mathcal{N}\models\mathsf{KPR}\mathrm{.}$ Thus every consistent extension of $\mathsf{KPR}+\Pi_{\infty}^{1}$-$\mathsf{DC}$ has a model of power $\aleph_{1}$ that has no proper end extension to model of $\mathsf{KPR}\mathrm{.}$
5.21. QuestionIs there an $\omega$-standard model of $\mathsf{ZF(C)}$ that has no proper end extension to a model of $\mathsf{ZF(C)}$?
5.22. Remark. Weakly Rubin models are never $\omega$-standard (even though $\omega$-standard rather classless models exist in abundance). However, a slight variation of the existence proof of Rubin models shows that a Rubin model $\mathcal{M}$ can be arranged to have an $\aleph_{1}$-like $\omega^{\mathcal{M}}.$ Note, however, that the answer to Question 5.21 is known to be in the negative if “end extension” is strengthened to “rank extension”; this follows from of [En-2, Theorem 1.5(b)], (which asserts that no $\aleph_{1}$-like model of $\mathsf{ZFC}$ that elementarily end extends a model all of whose ordinals are pointwise definable has a proper rank extension to a model of $\mathsf{ZFC}$. The proof of [En-2, Theorem 1.5(b)] together with part (b) of Theorem 4.4 makes it clear that the same goes for $\mathsf{ZF}$, i.e., there are $\omega$-standard models of $\mathsf{ZF}$ that have no proper rank extension to a model of $\mathsf{ZF}$.
6. ADDENDUM AND CORRIGENDUM TO [En-3]
Theorem 2.2 of [En-3] incorrectly stated (in the notation of the same paper, where Gothic letters are used for denoting structures) that if $\mathfrak{A}\models\mathsf{ZF}$, $\mathfrak{A}\prec_{\mathrm{cons}}\mathfrak{B}$ and $\mathfrak{B}$ fixes each $a\in\omega^{\mathfrak{A}},$ then $\mathfrak{A}$ is cofinal in $\mathfrak{B}$. As shown in Theorem 3.1 of this paper, it is possible to arrange taller conservative elementary extensions of models of $\mathsf{ZFC}$, thus Theorem 2.2 of [En-3] is false as stated. However, by assuming that ‘slightly’ strengthening the conditions:
$\mathfrak{A}\models\mathsf{ZF}$, and $\mathfrak{B}$ fixes each $a\in\omega^{\mathfrak{A}},$
to:
$\mathfrak{A}\models\mathsf{ZF}$, and $\mathfrak{B}$ fixes $\omega^{\mathfrak{A}},$
one obtains a true statement, as indicated in Corollary 4.2 of this paper. [^35] To my knowledge, there is only one place in the published literature in which Theorem 2.2 of [En-3] has been used, namely in a recent paper of Goldberg and Steel [GS]. In the discussion after Lemma 3.5 of the aforementioned paper, it is noted that Theorem 2.2 of [En-3] implies that an elementary embedding $\pi:\mathcal{M}\rightarrow\mathcal{N}$ (where $\mathcal{M}$ and $\mathcal{N}$ are inner models of $\mathsf{ZFC}$) has the property that $\pi^{-1}[S]\in M$ for each $S\in N$ (paraphrased in [GS] as $\pi$ is a close embedding of $\mathcal{M}$ into $\mathcal{N}$’) iff for each parametrically definable subset $D$ of $\mathcal{N}$, $\pi^{-1}[D]$ is parametrically definable in $\mathcal{M}$. In the context of [GS], the relevant models $\mathcal{M}$ and $\mathcal{N}$ are both $\omega$-models (they are inner models, thus well-founded), so Corollary 4.2 of this paper applies. As pointed out in Remark 3.9, $\mathfrak{B}$ fixes each $a\in\omega^{\mathfrak{A}}$ whenever $\mathfrak{A}\prec_{\mathrm{cons}}\mathfrak{B}\models\mathsf{ZF}$.
Several corollaries were drawn from Theorem 2.2 of [En-3]; what follows are the modifications in them, necessitated by the above correction. We assume that the reader has [En-3] handy for ready reference.
  • In Corollary 2.3(a) of [En-3], the assumption that $\mathfrak{B}$ fixes every integer of $\mathfrak{A}$ should be strengthened to: $\mathfrak{B}$ fixes $\omega^{\mathfrak{A}}$.
  • Part (b) of Corollary 2.3 of [En-3] is correct as stated in [En-3]: if $\mathfrak{A}$ is a Rubin model of $\mathsf{ZFC}$ and $\mathfrak{A}\prec\mathfrak{B}$ that fixes every $a\in\omega^{\mathfrak{A}}$, then $\mathfrak{B}$ is a conservative extension of $\mathfrak{A}$. This is really what the proof of part (a) of Corollary 2.3 of [En-3] shows. However, the following correction is needed: towards the end of the proof of part (a) of Corollary 2.3 of [En-3], after defining the set $X$, it must simply be stated as follows: But $X=Y\cap A,$ where $Y$ is the collection of members $a\in A$ satisfying $\varphi(b,a)\wedge a\in R_{\alpha}]$. Since $\mathfrak{B}$ is an end extension of $\mathfrak{C}$, $Y\in C,$ so by claim $(\ast)$, $X$ is definable in $\mathfrak{A}$.
  • Corollary 2.4 of [En-3] is correct as stated since it does not rely on the full force of Theorem 2.2 of [En-3] and only relies on the fact that no elementary conservative extension of a model of $\mathsf{ZF}$ is an end extension. The latter follows from Corollary 4.2 of this paper.
  • The condition stipulating that $\mathfrak{B}$ fixes $\omega^{\mathfrak{A}}$ must be added to condition $(i)$ of Corollary 2.5 of [En-3].
APPENDIX: PROOFS OF THEOREMS 5.15 & 5.17
A.1. Theorem (Rubin-Shelah-Schmerl) It is a theorem of $\mathsf{ZFC}$ that every countable model of $\mathsf{ZF}$ has an elementary extension to a weakly Rubin model of cardinality $\aleph_{1}$.
The proof of Theorem A.1 has two distinct stages: in the first stage we prove within $\mathsf{ZFC}+\Diamond_{\omega_{1}}$ that every countably infinite structure in a countable language has an elementary extension to a Rubin model of cardinality $\aleph_{1}$ (Theorem 5.15), and then in the second stage we use the result of the first stage together with a forcing-and-absoluteness argument to establish the theorem. [^36] It appears to be unknown whether Theorem 5.15 (existence of Rubin elementary extensions) can be proved in $\mathsf{ZFC}$ alone. However, as shown here, $\mathsf{ZFC}$ can prove that every countable model of $\mathsf{ZF}$ has an elementary extension to a weakly Rubin model. See also Remark A.7.
Stage 1 of the Proof of Theorem A.1
Fix a $\Diamond_{\omega_{1}}$ sequence $\left\langle S_{\alpha}:\alpha<\omega_{1}\right\rangle$. Given a countable language $\mathcal{L}$, and countably infinite $\mathcal{L}$-structure model $\mathcal{M}_{0}$, assume without loss of generality that $M_{0}=\omega.$ We plan to inductively build two sequences $\left\langle\mathcal{M}_{\alpha}:\alpha<\omega_{1}\right\rangle$, and $\left\langle\mathcal{O}_{\alpha}:\alpha<\omega_{1}\right\rangle.$ The first is a sequence of approximations to our final model $\mathcal{M}.$ The second sequence, on the other hand, keeps track of the increasing list of ‘obligations’ we need to abide by throughout the construction of the first sequence. More specifically, each $\mathcal{O}_{\alpha}$ will be of the form $\{\{V_{n}^{\alpha},\ W_{n}^{\alpha}\}:n\in\mathbb{\omega}\}$, where $\{V_{n}^{\alpha},\ W_{n}^{\alpha}\}$ is pair of disjoint subsets of $M_{\alpha}$ that are inseparable in $\mathcal{M}_{\alpha}$ and should be kept inseparable in each $\mathcal{M}_{\beta}$, for all $\beta>\alpha.$ Here we say that two disjoint subsets $V$ and $W$ of a model $\mathcal{N}$ are inseparable in $\mathcal{N}$ if there is no $\mathcal{N}$-definable $X\subseteq N$ such that $V\subseteq X,$ and $W\cap X=\varnothing.\mathfrak{\vskip 6.0pt plus 2.0pt minus 2.0pt}$
We only need to describe the construction of these two sequences for stages $\alpha+1$ for limit ordinals $\alpha$ since:
  • $\mathcal{O}_{0}:=\varnothing$.
  • For limit $\alpha$, $\mathcal{M}_{\alpha}:=\bigcup\limits_{\beta<\alpha}\mathcal{M}_{\beta}$ and $\mathcal{O}_{\alpha}:=\bigcup\limits_{\beta<\alpha}\mathcal{O}_{\beta}$.
  • For nonlimit $\alpha$ , $\mathcal{M}_{\alpha+1}:=\mathcal{M}_{\alpha}$ and $\mathcal{O}_{\alpha+1}:=\mathcal{O}_{\alpha}$.
At stage $\alpha+1,$ where $\alpha$ is a limit ordinal, we have access to a model $\mathcal{M}_{\alpha}$ (where $M_{\alpha}\in\omega_{1}$), and a collection $\mathcal{O}_{\alpha}$ of inseparable pairs of subsets of $M_{\alpha}$. We now look at $S_{\alpha}$, and consider two cases: either $S_{\alpha}$ is parametrically undefinable in $\mathcal{M}_{\alpha},$ or not. In the latter case we ‘do nothing’, and define $\mathcal{M}_{\alpha+1}:=\mathcal{M}_{\alpha}$ and $\mathcal{O}_{\alpha+1}:=\mathcal{O}_{\alpha}.$ But if the former is true, we augment our list of obligations via:
$\mathcal{O}_{\alpha+1}:=\mathcal{O}_{\alpha}\cup\{\left\{S_{\alpha},\ M_{\alpha}\backslash S_{\alpha}\right\}\}.$
Notice that if $S_{\alpha}$ is parametrically undefinable in $\mathcal{M}_{\alpha},$ then $S_{\alpha},\ M_{\alpha}\backslash S_{\alpha}\}$ is inseparable in $\mathcal{M}_{\alpha}.$ Then we use Lemma A.2 below to build an elementary extension $\mathcal{M}_{\alpha+1}$ of $\mathcal{M}_{\alpha}$ that satisfies the following three conditions:
(1) For each $\{V,\ W\}\in\mathcal{O}_{\alpha+1}$, $V$ and $W$ are inseparable in $\mathcal{M}_{\alpha+1}$.
(2) For every $\mathcal{M}_{\alpha}$-definable directed set $\mathbb{D}$ with no last element, there is some $d^{\ast}\in\mathbb{D}^{\mathcal{M}_{\alpha+1}}$ such that for each $d\in\mathbb{D}^{\mathcal{M}_{\alpha}},$ $\mathcal{M}_{\alpha+1}\models\left(d<_{\mathbb{D}}d^{\ast}\right).$
(3) $M_{\alpha+1}=M_{\alpha}+\omega$ (ordinal addition).
This concludes the description of the sequences $\left\langle\mathcal{M}_{\alpha+1}:\alpha<\omega_{1}\right\rangle$, and $\left\langle\mathcal{O}_{\alpha}:\alpha<\omega_{1}\right\rangle.$
Let $\mathcal{M}:=\bigcup\limits_{\alpha<\omega_{1}}\mathcal{M}_{\alpha},$ and $\mathcal{O}:=\bigcup\limits_{\alpha<\omega_{1}}\mathcal{O}_{\alpha}$; note that:
$(4)$ For each $\{V,\ W\}\in\mathcal{O}$, $V$ and $W$ are inseparable in $\mathcal{M}$.
We now verify that $\mathcal{M}$ is a Rubin model. Suppose, on the contrary, that for some $\mathcal{M}$-definable partial order $\mathbb{P}$, there is a maximal filter $F\subseteq\mathbb{P}$ that is not $\mathcal{M}$-definable, and $F$ contains a cofinal $\omega_{1}$-chain $E$. By a standard Löwenheim-Skolem argument there is some limit $\alpha<\omega_{1}$ such that:
[…]
In particular,
(6) $S_{\alpha}$ is an undefinable subset of $\mathcal{M}_{\alpha}.$
[…]
A.2. Lemma. Suppose $\mathcal{M}$ is a countable $\mathcal{L}$-structure, where $\mathcal{L}$ is countable. Let $\{\{V_{n},W_{n}\}:n\in\mathbb{\omega}\}$ is a countable list of $\mathcal{M}$-inseparable pairs of subsets of $M$, and $\left\{\left(\mathbb{D}_{n},\leq_{\mathbb{D}_{n}}\right):n\in\mathbb{\omega}\right\}$ be an enumeration of $\mathcal{M}$-definable directed sets with no last element. Then there exists an elementary extension $\mathcal{N}$ of $\mathcal{M}$ that satisfies the following two properties:
  1. (a)
    $V_{n}$ and $W_{n}$ remain inseparable in $\mathcal{N}$ for all $n\in\mathbb{\omega}$.
  2. (b)
    For each $n\in\mathbb{\omega}$ there is some $d_{n}\in N$ such that $\mathcal{N}\models d_{n}\in\mathbb{D}_{n}$ (i.e., $d_{n}$ satisfies the formula that defines $\mathbb{D}_{n}$ in $\mathcal{M}$), and for each $m\in\mathbb{D}_{n},$ $\mathcal{N}\models\left(m<_{\mathbb{D}_{n}}d_{n}\right).$
Let $\mathcal{L}^{+}$ be the language $\mathcal{L}$ augmented with constants from $C_{1}=\{\dot{m}:m\in M\}$ and $C_{2}=\left\{d_{n}:n\in\mathbb{\omega}\right\}$ (where $C_{1}$ and $C_{2}$ are disjoint), and consider the $\mathcal{L}^{+}$-theory $T$ below:
$T:=\mathrm{Th}(\mathcal{M},m)_{m\in M}+\left\{d_{n}\in\mathbb{D}_{n}:n\in\mathbb{\omega}\right\}\cup\left\{\dot{m}<_{\mathbb{D}_{n}}d_{n}:\mathcal{N}\models\left(m\in\mathbb{D}_{n}\right),n\in\mathbb{\omega}\right\}.$
$T$ is readily seen to be consistent since it is finitely satisfiable in $\mathcal{M}\mathfrak{.}$ Moreover, it is clear that if $\mathcal{N}\models T$, then $\mathcal{M}\prec\mathcal{N}$ and $\mathcal{N}$ satisfies condition (b) of the theorem. To arrange a model of $T$ in which conditions (a) also holds requires a delicate omitting types argument. First, we need a pair of preliminary lemmas. The proofs of Lemma A.3 and A.4 are routine and left to the reader. Note that they are each other’s contrapositive, so it is sufficient to only verify one of them.
A.3. Lemma. The following two conditions are equivalent for a sentence $\varphi(d_{k_{1}},\cdot\cdot\cdot,d_{k_{p}})$ of $\mathcal{L}^{+}$.
  1. $(i)$
    $T\vdash\varphi(d_{k_{1}},\cdot\cdot\cdot,d_{k_{p}})$.
  2. $(ii)$
    $\mathcal{M}\models\exists r_{1}\in\mathbb{D}_{k_{1}}\forall s_{1}\in\mathbb{D}_{k_{1}}\cdot\cdot\cdot\exists r_{p}\in\mathbb{D}_{k_{p}}\forall s_{p}\in\mathbb{D}_{k_{p}}\left[\left(\bigwedge\limits_{1\leq i\leq p}r_{i}<_{\mathbb{D}_{k_{i}}}s_{i}\right)\rightarrow\varphi(s_{1},\cdot\cdot\cdot,s_{p})\right].$
A.4. Lemma. The following two conditions are equivalent for a sentence $\varphi(d_{k_{1}},\cdot\cdot\cdot,d_{k_{p}})$ of $\mathcal{L}^{+}$.
  1. $(i)$
    $T+\varphi(d_{k_{1}},\cdot\cdot\cdot,d_{k_{p}})$ is consistent.
  2. $(ii)$
    $\mathcal{M}\models\forall r_{1}\in\mathbb{D}_{k_{1}}\exists s_{1}\in\mathbb{D}_{k_{1}}\cdot\cdot\cdot\forall r_{p}\in\mathbb{D}_{k_{p}}\exists s_{p}\in\mathbb{D}_{k_{p}}\left[\left(\bigwedge\limits_{1\leq i\leq p}r_{i}<_{\mathbb{D}_{k_{i}}}s_{i}\right)\wedge\varphi(s_{1},\cdot\cdot\cdot,s_{p})\right].$
We are now ready to carry out an omitting types argument to complete the proof of Lemma A.2. For each formula $\psi(y,\vec{x})$ of $\mathcal{L}^{+}$, and each $n\in\omega$, consider the following $q$-type formulated in the language $\mathcal{L}^{+}$, where $\vec{x}=\left\langle x_{1},\cdot\cdot\cdot,x_{q}\right\rangle$ is a $q$-tuple of variables:
$\Sigma_{n}^{\psi}(\vec{x}):=\{\psi(\dot{a},\vec{x}):a\in V_{n}\}\cup\{\lnot\psi(\dot{b},\vec{x}):b\in W_{n}\}.$
Note that $\Sigma_{n}^{\psi}$ expresses that $\psi(y,\vec{x})$ separates $V_{n}$ and $W_{n}.$
A.5. Lemma. $\Sigma_{n}^{\psi}$ is locally omitted by $T$ for each formula $\psi(y,\vec{x})$, and each $n\in\mathbb{\omega}$.
Proof. Suppose to the contrary that there is a formula $\theta(\vec{x},d_{k_{1}},\cdot\cdot\cdot,d_{k_{p}})$ of $\mathcal{L}^{+}$ and some $n\in\omega$ such that (1) through (3) below hold:
  1. $(1)$
    $T+\exists\vec{x}\ \theta(\vec{x},d_{k_{1}},\cdot\cdot\cdot,d_{k_{p}})$ is consistent.
  2. $(2)$
    For all $a\in V_{n},$ $T\vdash\forall\vec{x}\left[\theta(\vec{x},d_{k_{1}},\cdot\cdot\cdot,d_{k_{p}})\rightarrow\psi(\dot{a},\vec{x})\right].$
  3. $(3)$
    For all $b\in W_{n},$ $T\vdash\forall\vec{x}\left[\theta(\vec{x},d_{k_{1}},\cdot\cdot\cdot,d_{k_{p}})\rightarrow\lnot\psi(\dot{b},\vec{x})\right].$
Invoking Lemmas A.3 and A.4, (1) through (3) translate to $(1^{\prime})$ through $(3^{\prime})$ below:
  1. $(1^{\prime})$
    $\mathcal{M}\models\forall r_{1}\in\mathbb{D}_{k_{1}}\exists s_{1}\in\mathbb{D}_{k_{1}}\cdot\cdot\cdot\forall r_{p}\in\mathbb{D}_{k_{p}}\exists s_{p}\in\mathbb{D}_{k_{p}}\left[\left(\bigwedge\limits_{1\leq i\leq p}r_{i}<_{\mathbb{D}_{k_{i}}}s_{i}\right)\wedge\exists\vec{x}\ \theta(\vec{x},\vec{s})\right],$ where $\vec{s}=\left\langle s_{1},\cdot\cdot\cdot,s_{p}\right\rangle.$
  2. $(2^{\prime})$
    For all $a\in V_{n}$, $\mathcal{M}\models\lambda(a),$ where:
    $\lambda(\mathbf{v}):=\forall\vec{x}\ \exists r_{1}\in\mathbb{D}_{k_{1}}\forall s_{1}\in\mathbb{D}_{k_{1}}\cdot\cdot\cdot\exists r_{p}\in\mathbb{D}_{k_{p}}\forall s_{p}\in\mathbb{D}_{k_{p}}\left[\left(\bigwedge\limits_{1\leq i\leq p}r_{i}<_{\mathbb{D}_{k_{i}}}s_{i}\right)\rightarrow\left(\theta(\vec{x},\vec{s})\rightarrow\psi(\mathbf{v},\vec{x}\ )\right)\right].$
  3. $(3^{\prime})$
    For all $b\in W_{n},\ \mathcal{M}\models\gamma(b)$, where:
    $\gamma(\mathbf{w}):=\forall\vec{x}\ \exists r_{1}\in\mathbb{D}_{k_{1}}\forall s_{1}\in\mathbb{D}_{k_{1}}\cdot\cdot\cdot\exists r_{n}\in\mathbb{D}_{k_{p}}\forall s_{p}\in\mathbb{D}_{k_{p}}\left[\left(\bigwedge\limits_{1\leq i\leq p}r_{i}<_{\mathbb{D}_{k_{i}}}s_{i}\right)\rightarrow\left(\theta(\vec{x},\vec{s})\rightarrow\lnot\psi(\mathbf{w},\vec{x})\right)\right].$
Let $\Lambda:=\{m\in M:\mathcal{M}\models\lambda(m)\}$, and $\Gamma:=\{m\in M:\mathcal{M}\models\gamma(m)\}$, and observe that $V_{n}\subseteq\Lambda$ by $(2^{\prime})$ and $W_{n}\subseteq\Gamma$ by $(3^{\prime})$. To arrive at a contradiction, we will show that $\Lambda\cap\Gamma=\emptyset$, which implies that $V_{n}$ and $W_{n}$ are separable in $\mathcal{M}$. To this end, suppose to the contrary that for some $m\in M,$
  1. $(4)$
    $\mathcal{M}\models\lambda(m)\wedge\gamma(m).$
Since each $\left(\mathbb{D}_{n},\leq_{\mathbb{D}_{n}}\right)$ is a directed set, (4) implies:
  1. $(5)$
    $\mathcal{M}\models\forall\vec{x}\ \exists r_{1}\in\mathbb{D}_{k_{1}}\forall s_{1}\in\mathbb{D}_{k_{1}}\cdot\cdot\cdot\exists r_{p}\in\mathbb{D}_{k_{p}}\forall s_{p}\in\mathbb{D}_{k_{p}}\left[\left(\bigwedge\limits_{1\leq i\leq p}r_{i}<_{\mathbb{D}_{k_{i}}}s_{i}\right)\rightarrow\left(\psi(m,\vec{x})\wedge\lnot\psi(m,\vec{x})\right)\right].$
Recall that no $\left(\mathbb{D}_{n},\leq_{\mathbb{D}_{n}}\right)$ has a last element. This shows that (5) yields a contradiction, so the proof is complete.
$\square$ (Lemma A.5)
Proof of Lemma A.2. Put Lemma A.5 together with the Henkin-Orey omitting types theorem [CK, Theorem 2.2.9] to conclude that there exists a model $\mathcal{N}$ of $T$ that satisfies properties (a) and (b). $\square$
Stage 2 of the Proof of Theorem A.1
In this stage we employ the result of the first stage together with some set-theoretical considerations to establish the Rubin-Shelah-Schmerl Theorem. The main idea here is a variant of the one used by Schmerl, [Schm-2], which itself based on a method of $\diamondsuit_{\omega_{1}}$-elimination introduced by Shelah [Sh]. This stage has three steps.
Step 1 of Stage 2. It is well-known that there is an $\omega_{1}$-closed notion $\mathbb{Q}_{1}$ in the universe $\mathrm{V}$ of $\mathsf{ZFC}$ such that the $\mathbb{Q}_{1}$-generic extension $\mathrm{V}^{\mathbb{Q}_{1}}$ of $\mathrm{V}$ satisfies $\mathsf{ZFC}+\diamondsuit_{\mathbb{\omega}_{1}}$ [Ku, Theorem 8.3] (note that $\aleph_{1}$ is absolute between $\mathrm{V}$ and $\mathrm{V}^{\mathbb{Q}_{1}}$ since $\mathbb{Q}_{1}$ is $\omega_{1}$-closed). Thus by Step 1, given a model $\mathcal{M}$ in the universe $\mathrm{V}$ of $\mathsf{ZF}$, the forcing extension $\mathrm{V}^{\mathbb{Q}_{1}}$ satisfies “there is an elementary extension $\mathcal{N}$ of $\mathcal{M}$ such that $\mathcal{N}$ is a Rubin model”. In $\mathrm{V}^{\mathbb{Q}_{1}}$, since $\mathcal{N}$ is a Rubin model, for each $s\in N$ that is infinite in the sense of $\mathcal{N}$, there is a subset $C_{s}=\left\{c_{\alpha}^{s}:\alpha\in\omega_{1}\right\}$ of $\mathcal{N}$ such that $\left\langle c_{\alpha}^{s}:\alpha\in\omega_{1}\right\rangle$ is an $\omega_{1}$-chain that is cofinal in the directed set $\left([s]^{<\omega}\right)^{\mathcal{M}}.$
  • Let $\mathbb{P}_{s}$ be the suborder of $\left(\mathrm{Fin}(s,2)\right)^{\mathcal{M}}$ consisting of partial $\mathcal{M}$-finite functions $f$ from $s$ into $2$ such that $\mathrm{dom}(f)\in C_{s}$. Note that $\mathbb{P}_{s}$ is a tree-order (the predecessors of any given element are comparable).
  • $\mathbb{P}_{s}$ can be turned into a ranked-tree $\tau_{s}$ by defining $\rho(f)=\alpha$ if $\mathrm{dom}(f)=c_{\alpha}$. The ranked tree $\tau_{s}$ has the key property that each maximal filter of $\left(\mathrm{Fin}(s,2)\right)^{\mathcal{M}}$ is uniquely determined by a branch of $\tau_{s}.$
  • Let $C$ be the binary predicate on $\mathcal{N}$ defined by $C(s,x)$ iff $x\in C_{s}$, and let
$\mathcal{N}^{+}=\left(\mathcal{N},C\right)$.
Note that the $\mathcal{N}^{+}$-definable ranked trees include trees of the form $\tau_{s}$, as well as the ranked tree $\tau_{\mathrm{class}}^{\mathcal{N}}$ (whose branches uniquely determine classes of $\mathcal{N}$, as noted in Remark 5.14).
  • This gives us $\aleph_{1}$-many ranked trees ‘of interest’ in $\mathrm{V}^{\mathbb{Q}_{1}},$ namely $\tau_{\mathrm{class}}^{\mathcal{N}}$ together with $\tau_{s}$ for each infinite set $s$ of $\mathcal{N}$.
Step 2 of Stage 2. The concept of a weakly specializing function plays a key role in this step. Given a ranked tree $(\mathbb{T},\ \leq_{\mathbb{T}},\ \mathbb{L},\ \leq_{\mathbb{L}},\ \rho)$, we say that $f:\mathbb{T\rightarrow\omega}$ weakly specializes $(\mathbb{T},\ \leq_{\mathbb{T}})$ if $f$ has the following property:
If $x\leq_{\mathbb{T}}y$ and $x\leq_{\mathbb{T}}z$ and $f(x)=f(y)=f(z),$ then $y$ and $z$ are $\leq_{\mathbb{T}}$-comparable.
The following lemma captures the essence of a weakly specializing function for our purposes.
A.6. Lemma. Suppose $\tau=(\mathbb{T},\ \leq_{\mathbb{T}},\ \mathbb{L},\ \leq_{\mathbb{L}},\ \rho)$ is a ranked tree, where $(\mathbb{L},\ \leq_{\mathbb{L}})$ has cofinality $\omega_{1}$, and $f$ weakly specializes $(\mathbb{T},\ \leq_{\mathbb{T}})$. Then every branch of $\tau$ is parametrically definable in the expansion $\left(\tau,f\right)$ of $\tau.$
Proof. Suppose $B$ is a branch of $\tau$. Then since ($\mathbb{L},\ \leq_{\mathbb{L}})$ has cofinality $\omega_{1},$ there is a subset $S$ of $B$ of order type $\omega_{1}$ that is unbounded in $B$. Therefore there is some subset $B_{0}\$of $B$ that is unbounded in $B$ such that $f$ is constant on $B_{0}.$ Fix some element $b_{0}\in B_{0}$ and consider subset $X$ of $\mathbb{T}$ that is defined in $\left(\tau,f\right)$ by the formula:
$\varphi(x,b_{0}):=[x\leq_{\mathbb{T}}b_{0}\vee(b_{0}\leq_{\mathbb{T}}x\wedge f(b_{0})=f(x))].$
By the defining property of weakly specializing functions, $X$ is linearly ordered and is unbounded in $B$. This makes it clear that $B$ is definable in $\left(\tau,f\right)$ as the downward closure of $X$, i.e., by the formula $\psi(x,b_{0}):=\exists y\left[\varphi(y,b_{0})\wedge x\leq_{\mathbb{T}}y\right].$ $\square$
By a remarkable theorem, due independently to Baumgartner [Bau] and Shelah [Sh], given any ranked tree $\tau$ of cardinality $\aleph_{1}$ and of cofinality $\omega_{1}$, there is a c.c.c. notion of forcing $\mathbb{Q}_{\tau}$ such that $\mathrm{V}^{\mathbb{Q}_{1}\ast\mathbb{Q}_{\tau}}:=\left(\mathrm{V}^{\mathbb{Q}_{1}}\right)^{\mathbb{Q}_{\tau}}$ contains a function $f_{\tau}$ such that weakly specializes $\tau.$ [^37] This result generalizes the fact, established by Baumgartner, Malitz, and Reinhardt [BMR] that every $\omega_{1}$-Aronszajn tree can be turned into a special Aronszajn tree in a c.c.c. forcing extension of the universe. Here $\mathbb{Q}_{\tau}$ consists of all finite attempts for building a function that weakly specializes $\tau$ (ordered by inclusion). Moreover, given any collection $\{\tau_{\alpha}:\alpha\in\kappa\}$ of ranked trees of cofinality $\omega_{1}$, the finite support product $\mathbb{Q}$ of $\left\{\mathbb{Q}_{\tau_{\alpha}}:\alpha\in\kappa\right\}$ is also known to have the c.c.c. property [^38] It is well-known that the finite support product of a given collection of forcing notions $\left\{\mathbb{P}_{\alpha}:\alpha\in\kappa\right\}$ has the c.c.c. property iff every finite sub-product of $\left\{\mathbb{P}_{\alpha}:\alpha\in\kappa\right\}$ has the c.c.c. property., as indicated in [CZ, Corollary 3.3]393939[CZ, Corollary 3.3] is overtly about specializing $\omega_{1}$-Aronszajn trees, but as shown by Baumgartner [Bau, Section 7], given a tree $\tau$ of height $\omega_{1}$ with at most $\aleph_{1}$ many branches, there is an Aronszajn subtree $\tau_{0}$ of $\tau$ of height $\omega_{1}$, and moreover, any function $f_{0}$ that specializes $\tau_{0}$ can be canonically extended to a function $f$ that specializes $\tau.$ For more detail, see the proof of [Sw, Lemma 3.5].. Note that for each $\alpha\in\kappa$, $\mathbb{Q}$ forces the existence of a function $f_{\tau_{\alpha}}$ that weakly specializes $\tau_{\alpha}.$
  • Thus by choosing $\{\tau_{\alpha}:\alpha\in\omega_{1}\}$ to be the collection of ranked trees that are parametrically definable in $\mathcal{N}^{+}$, there is a function $g(s,x)$ in $\mathrm{V}^{\mathbb{Q}_{1}\ast\mathbb{Q}}:=\left(\mathrm{V}^{\mathbb{Q}_{1}}\right)^{\mathbb{Q}}$ such that for each infinite set $s$ of $\mathcal{N}$, $g(s,x)$ weakly specializes $\tau_{s}$ ($\tau_{s}$ was defined in Step 1 of Stage 2), and a function $f$ that weakly specializes $\tau_{\mathrm{class}}^{\mathcal{N}}.$
The next lemma shows that forcing with $\mathbb{Q}$ does not add new branches to ranked trees in the ground model that have cofinality $\omega_{1}$. Note that when the lemma below is applied to the case when V is replaced by $\mathrm{V}^{\mathbb{Q}_{1}}$
, it shows that the ample supply of rather branchless models in $\mathrm{V}^{\mathbb{Q}_{1}}$ remain rather branchless after forcing with $\mathbb{Q}$, i.e., in $\mathrm{V}^{\mathbb{Q}_{1}\ast\mathbb{Q}}.$
A.7. Lemma. Suppose $\tau=(\mathbb{T},\ \leq_{\mathbb{T}},\ \mathbb{L},\ \leq_{\mathbb{L}},\ \rho)$ is a ranked tree, where $(\mathbb{L},\ \leq_{\mathbb{L}})$ has cofinality $\omega_{1}.$ If $B$ is a branch of $\tau$ in $\mathrm{V}^{\mathbb{Q}}$, then $B\in\mathrm{V}$.
Proof. Suppose $B$ is a branch of $\tau$ in $\mathrm{V}^{\mathbb{Q}}$, and let $f$ be the $\mathbb{Q}$-generic function in $\mathrm{V}^{\mathbb{Q}}$ that specializes $\tau.$ Since $\mathbb{Q}$ does not collapse $\omega_{1}$ and the rank order of $\tau$ is assumed to have cofinality $\omega_{1}$ in $\mathrm{V}$, there is for some $k\in\omega$ such that:
[…]
In $\mathrm{V}^{\mathbb{Q}},$ let $x_{0}\in B$ with $f(x_{0})=k.$ Coupled with (1) this shows that there is some $q_{0}\in\mathbb{Q}$ such that:
(2) $q_{0}\Vdash$ $\left[f(x_{0})=k\wedge\exists y>_{\mathbb{T}}x_{0}(y\in B\wedge f(y)=k)\right]$.
Let $B_{0}$ be the subset of $\mathbb{T}$ defined in $\mathrm{V}$ as:
$B_{0}=\left\{y\in\mathbb{T}:\exists q\in\mathbb{Q\ }(q\supseteq q_{0}\wedge q\Vdash\left[y>_{\mathbb{T}}x_{0}\wedge y\in B\wedge f(y)=k\right]\right\}.$
We wish to show:
$(\ast)\ \ \mathrm{V}^{\mathbb{Q}}\models$ $B_{0}$ is a cofinal subset of $B$.
To verify $(\ast)$, suppose $b_{0}\in B$. By (1), for some $y\in\mathbb{T}$, $\mathrm{V}^{\mathbb{Q}}$ satisfies $y>_{\mathbb{T}}b_{0}$, $y\in B$, and $f(y)=k$. Therefore there is some $\mathbb{Q}$-condition $q\supseteq q_{0}$ such that:
$q\Vdash\left[y>_{\mathbb{T}}x_{0}\wedge y\in B\wedge f(y)=k\right].\vskip 6.0pt plus 2.0pt minus 2.0pt$
Since $B_{0}\in\mathrm{V}$, $(\ast)$ makes it clear that $B\in\mathrm{V}$ as well, since $B$ can be defined in $\mathrm{V}$ as the downward closure of $B_{0}.$ $\square$
  • Let $\Omega$ be the collection consisting of the ranked tree $\tau_{\mathrm{class}}^{\mathcal{N}}$ together with the family of ranked trees $\left\{\tau_{s}^{\mathcal{N}}:\mathcal{N}\models\left|s\right|\geq\aleph_{0}\right\}.$ Since $\mathcal{N}$ is a Rubin model in $\mathrm{V}$, each of the trees in $\Omega$ are rather branchless in $\mathrm{V}$, and Lemma A.7 assures us the ranked trees in $\Omega$ remain rather branchless in $\mathrm{V}^{\mathbb{Q}}$. Lemma A.6, on the other hand, assures us that each branch of the ranked trees in $\Omega$ has a simple definition in $\mathrm{V}^{\mathbb{Q}}.$
  • Let $\mathcal{L}_{\mathrm{set}}^{+}$ be the finite language obtained by augmenting the usual language $\mathcal{L}_{\mathrm{set}}$ of set theory with extra symbols for denoting the binary relation $C$, the binary function $g,$ and the unary function $f$. Then let $\mathcal{L}$ be the countable language that is the result of augmenting $\mathcal{L}_{\mathrm{set}}^{+}$ with constants for each element of $\mathcal{M}$. The content of the above bullet item allows us to write a single sentence $\psi$ in $\mathcal{L}_{\mathsf{\omega}_{\mathsf{1}}\mathsf{,\omega}}(\mathsf{Q})$ (where $\mathsf{Q}$ is the quantifier “there exist uncountably many”) that captures the following salient features of the structure $\left(\mathcal{N}^{+},C,g,f\right)$ that hold in $\mathrm{V}^{\mathbb{Q}}$:
  1. $(1)$
    The $\mathcal{L}_{\mathrm{set}}$-reduct of $\left(\mathcal{N}^{+},C,g,f\right)$ elementarily extends $\mathcal{M}$.
  2. $(2)$
    For each infinite $s$, $\tau_{s}^{\mathcal{N}}$ has uncountable cofinality, is rather branchless, and $g(s,x)$ weakly specializes $\tau_{s}^{\mathcal{N}}$.
  3. $(3)$
    $\tau_{\mathrm{class}}^{\mathcal{N}}$ has uncountable cofinality, is rather branchless, and $f$ weakly specializes $\tau_{\mathrm{class}}^{\mathcal{N}}$.
More specifically, $\psi$ is the conjunction of the following sentences:
  • $\psi_{1}=$ the conjunction of the countably many sentences in the elementary diagram of $\mathcal{M}.$
  • $\psi_{2}=$ the sentence expressing (2) above.
  • $\psi_{3}=$ the sentence expressing (3) above.
We elaborate on how to write $\psi_{2}$, a similar idea is used to write $\psi_{3}.$ The quantifier Q allows us to express that $\tau_{s}$ has uncountable cofinality, and the range of the specializing function $g(s,x)$ is countable. To expresses the rather branchless feature of $\tau_{s}$, let $\psi(x,y)$ be as in the proof of Lemma A.6. Note that in the notation of Lemma A.6, for any $b\in\mathbb{T}$ $\{x\in\mathbb{T}_{s}:\psi(x,b)\}$ is linearly ordered by $\leq_{\mathbb{T}}.$ By Lemma A.6, in $\mathrm{V}^{\mathbb{Q}},$ if $s$ is infinite and $B$ is a branch of $\tau_{s}$, then there is some $b_{0}\in B$ such that $\psi(x,b_{0})$ defines $B$. Together with the fact that $\tau_{s}$ is rather branchless in $\mathrm{V}^{\mathbb{Q}}$ (thanks to Lemma A.7) the $\mathcal{L}_{\mathsf{\omega}_{\mathsf{1}}\mathsf{,\omega}}$-formula below (no need for the quantifier $\mathsf{Q}$ here) expresses the rather branchless feature of $\tau_{s}$. In the formula below $\mathbb{T}_{s}$ is the set of the nodes of the ranked tree $\tau_{s}$, and “unbounded rank” is short for “unbounded $\tau_{s}$-rank”.
$\forall b\in\mathbb{T}_{s}\left[\{x\in\mathbb{T}_{s}:\psi(x,b)\}\ \mathrm{has\ unbounded\ rank\ }\rightarrow\bigvee\limits_{\theta(x,y)\in\mathcal{L}_{\mathrm{set}}^{+}}\exists y\ \forall x\ \left(\psi(x,b)\leftrightarrow\theta(x,y)\right)\right].$
Step 3 of Stage 2. Let $\psi$ be the sentence described above. Since $\psi$ has a model in $\mathrm{V}^{\mathbb{Q}_{1}\ast\mathbb{Q}}$, and by Keisler’s completeness theorem for $\mathcal{L}_{\omega_{1},\omega}(\mathsf{Q})$ [Ke-2], the consistency of sentences in $\mathcal{L}_{\omega_{1},\omega}(\mathsf{Q})$ is absolute for extensions of $\mathrm{V}$ that do not collapse $\aleph_{1},$ we can conclude that $\psi$ has a model in $\mathrm{V}$ since $\aleph_{1}$ is absolute between $\mathrm{V}$ and $\mathrm{V}^{\mathbb{Q}_{1}\ast\mathbb{Q}}$ (recall that $\mathbb{Q}_{1}$ is $\omega_{1}$-closed and $\mathbb{Q}$ is c.c.c.).). Thanks to Lemmas A.6 and A.7, the $\mathcal{L}_{\mathrm{set}}$-reduct of any model of $\psi$ is a weakly Rubin model that elementarily extends $\mathcal{M}$ (and the cardinality of the model can be arranged to be $\aleph_{1}$ by a Löwenheim-Skolem argument). This concludes the proof of Theorem A.1. $\square$
A.7. Remark. Schmerl [Schm-2] used a similar argument as the one used in the proof of Theorem A.1 to show that every countable model $\mathcal{M}$ of $\mathsf{ZFC}$ has an elementary extension $\mathcal{N}$ of cardinality $\aleph_{1}$ such that every $\omega^{\mathcal{N}}$-complete ultrafilter over $\mathcal{N}$ is $\mathcal{N}$-definable (and thus coded in $\mathcal{N}$). The proof of Theorem A.1 can be dovetailed with Schmerl’s proof so as to show that every countable model $\mathcal{M}$ of $\mathsf{ZFC}$ has an elementary extension $\mathcal{N}$ of cardinality $\aleph_{1}$ that is weakly Rubin and also has the additional property that every $\omega^{\mathcal{N}}$-complete ultrafilter over $\mathcal{N}$ is $\mathcal{N}$-definable.